### v1

# JEL: C81, O32

Across fields, the quantity of research has increased substantially, without an attendant increase in output. We argue that, in medicine, this indicator of declining productivity reflects a compositional shift toward low-capital, low-productivity research. Using a fine-tuned, open-source large language model, we construct a novel census of capital-intensive, high-productivity medical investment—clinical trials. Since 2010, the annual quantity of clinical trials has been constant. By contrast, the quantity of other forms of clinical research has increased substantially. Within clinical trials, there is substantial heterogeneity in productivity. The distribution of this heterogeneity is stable over time.
Keywords: Medical Innovation, Productivity, Large Language Models
JEL: C81, O32
1. Introduction
Technological progress in medicine is widely credited with enabling large improvements in human health (Easterlin,, 1999; Mokyr et al.,, 2002; Acemoglu and Johnson,, 2007). Breakthroughs at the turn of the century—the completion of the Human Genome Project, the advent of DNA sequencing, the adoption of combinatorial chemistry—promised transformation of diagnosis and treatment or, at least, an expansion of the set of available scientific opportunities. [^1] See Scannell et al., (2012) for details on these factors, in a broader discussion of declining productivity in the pharmaceutical industry. Combinatorial chemistry, in the 1980s and 1990s, increased the number of molecules (potential drug candidates) that a chemist could synthesize “perhaps 800-fold.” Similarly, since the first genome sequence was determined in the 1970s, DNA sequencing has become “over a billion times faster.” Yet, an emerging literature finds that biomedical research productivity is in a state of decline; see Goldin et al., (2024) for a review. This literature is structured around an incongruity identified by Bloom et al., (2020). The quantity of medical research—proxied by estimates of the number of clinical trials and scientific publications—has increased substantially, without an accompanying improvement in mortality. Many explanations center on characteristics of the “idea production function.” Perhaps low-hanging fruit have been depleted (the fishing out hypothesis, described in Jones,, 1995), [^2] See, also, Cowen, (2011) and Gordon, (2017). or else the educational costs to a young innovator of reaching the frontier of knowledge are increasingly high (the burden of knowledge hypothesis, due to Jones,, 2009). Though the specific parameter implicated differs, the mechanisms underlying these hypotheses suggest a structural change in research productivity.
This paper tests an alternative, compositional hypothesis. We document that increases in the quantity of medical research have been driven by increases in the quantity of unproductive, low-capital research with limited capacity to effect quantitatively significant changes to mortality. As a consequence, increased quantities of medical research need not indicate that the productivity of high-capital research investment is declining.
We construct a novel census of capital-intensive, high-productivity medical investment—clinical trials. We contrast this set with its complement in clinical research, proxied by publications that cite clinical trials. This comparison is natural and reflects the standard hierarchy of medical evidence (Jones and Podolsky,, 2015). Decision-making in medicine—by regulators, firms, insurers, physicians, and patients—hinges on the causal, mortality-relevant data produced by costly, time-intensive randomized experiments (Chavez-MacGregor and Giordano,, 2016; Alsan et al.,, 2024). Although research engaging with trial evidence undoubtedly produces valuable information, it does so at comparatively low cost and with lesser impact (Kones et al.,, 2014). Since 2010, we find that the quantity, quality, and composition of published clinical trials has been stable. By contrast, the quantity of other clinical research has doubled.
No existing dataset contains sufficient information to assess the composition of clinical research. Administrative databases—such as ClinicalTrials.gov—aim to provide an index of trials, but suffer from widespread noncompliance with reporting requirements (DeVito et al.,, 2020). Proprietary databases take as an input the contents of ClinicalTrials.gov and other registries and, thus, capture changes in reporting patterns over time (e.g., Cortellis,, 2024). Best practices for identifying trials in publication data do so imprecisely, with both high true and false positive rates (Thomas et al.,, 2021).
We construct a new census of clinical trials. We collect all records of scientific publications indexed in the National Library of Medicine’s PubMed / MEDLINE database, from 2010 forward, that disclose the results of a clinical trial studying a medicine in human subjects. Our definition is restrictive and intended to capture only those research inputs that are most relevant for human mortality. Our goal is to identify clinical trials in our sample of interest—from the set of roughly 35 million PubMed records—with an accuracy and precision close to that of a human labeller. Standard machine learning methods perform poorly and hand-labelling at scale is infeasible. Instead, we train a bespoke large language model, optimized for our task.
We proceed in stages. We devise an interface to iteratively revise model prompts. With the highest-performing prompts, we use proprietary language models, OpenAI’s GPT-3.5 and GPT-4, to classify a large number of records (Achiam et al.,, 2023; Bubeck et al.,, 2023). These noisy labels are used as training data to fine-tune a set of open-source models. The resulting model matches the performance of the best proprietary model and produces a sample of approximately 150,000 clinical trials.
We document two sets of facts. First, there is quantitatively significant heterogeneity in both the composition of medical research and the composition of clinical trials. While clinical trial counts are essentially constant over time, the quantity of papers citing trials has increased by nearly a factor of two. This trend reflects a large increase in the quantity of research originating from China, in addition to other countries who, individually, do not account for a large share of the total quantity of medical research. Within clinical trials, proxies for quality and informativeness suggest considerable heterogeneity. Roughly half of trials in our sample are never cited by a leading medical journal and nearly 15 percent are never cited by another scientific publication. Second, these forms of heterogeneity within trials are stable over time. We find that—across measures of quality, importance, and geography—the distribution of clinical trial heterogeneity is unchanging. Taken together, we interpret these facts as evidence that the productivity of high-capital medical investment is stable over our time period.
Our choice of setting reflects two substantive considerations. First, the determinants of technological progress in medicine are interesting in their own right. [^3] Cutler and McClellan, (2001) observe that technological change has accounted for the bulk of increases in health spending over time, but that health has improved alongside these increases. In a series of case studies, they argue that the benefits of technological progress exceed their costs. See, also, Almond et al., (2010), who estimate that the returns to medical spending are positive on the margin. Second, and more centrally, nearly all micro-level evidence on rising research investment, and declining productivity, comes from studies of clinical trials and scientific publications. [^4] See Goldin et al., (2024) for a summary. Notable exceptions are studies of agriculture and semiconductors. Measured increases in the quantity of clinical trials (Bloom et al.,, 2020) and average “declining disruptiveness” of scientific papers (Park et al.,, 2023) are at the center of active debates about whether research in science and medicine has hit a point of diminishing returns (e.g., Dieppe and Kose,, 2020). [^5] See discussions in the popular press, e.g. Collison and Nielsen, (2018); Broad, (2023); Thompson, (2021); Piper, (2023). Concerns about declining productivity have influenced a wave of policy interest in new mechanisms to promote research and development, as well as the revival of “industrial policy” (see e.g., Council of Economic Advisers,, 2021, 2024; Lindsey and Hammond,, 2020; Hausmann,, 2023). A precise characterization of compositional changes in this industry, then, is a directly policy-relevant object.
We propose and validate the existence of quantitatively significant heterogeneity in medical research. [^6] Park et al., (2023) examine “disruptiveness” in scientific research, using data on the universe of publications across six decades. They document that the absolute counts of highly disruptive papers remain constant, consistent with our finding that high-capital, high-productivity research remains stable. Our setting allows for a clear delineation between types of research investments and, importantly, allows us to take a stand on the value of information produced by these distinct types for welfare-relevant outcomes. This is the primary contribution of this paper. We measure research inputs with a precision that is new to this literature, enabled by the existence of novel technology for data construction at scale (Achiam et al.,, 2023; Bubeck et al.,, 2023). Our decomposition suggests that measures of research investment that neglect these differences—and, instead, infer uniform productivity or capital for scientific investments—can yield misleading conclusions about industry-wide productivity.
This paper contributes to a growing literature on the causes of declining productivity, in general and in medicine. Additional explanations include market imperfections and failures—including barriers to innovation (see e.g., Aghion et al.,, 2019), inadequate public sector support for research (Gruber and Johnson,, 2019), and productivity imbalances across sectors (Acemoglu et al.,, 2024). Our work is related to a literature—which Syverson, (2017) summarizes as putting forth a “mismeasurement hypothesis”—that views the measured decline in aggregate productivity as illusory (see e.g., Mokyr,, 2014). In contrast, our hypothesis is not that productivity is mismeasured, per se, but that failure to account for heterogeneity can generate a misattribution of trends. More directly, Myers and Pauly, (2019), Scannell et al., (2012), and Cockburn, (2006) investigate the determinants of declining productivity in the pharmaceutical sector, with a focus on large increases in factor prices for clinical trials rather than quantities.
Finally, we draw on, and contribute to, a long literature on the measurement of innovation. Since Pakes, (1986), economists have recognized that patents—the most commonly used measure of research output—differ greatly in economic and technological significance. Griliches, (1990) provides a classic discussion of the difficulties associated with using patents as economic indicators, given this heterogeneity. [^7] Bryan and Williams, (2021) provide an updated survey and discuss the challenges of devising credible measures of innovative effort. Our work builds on these insights to investigate the implications of heterogeneity in inputs for the use of medical research as a productivity indicator.
2. Productivity, Capital, and Medical Research
Studies of research productivity have, at their core, some notion of an “idea production function”—a mapping from research inputs (effort) to outputs (ideas). The canonical Romer, (1990) specification is given by
$\frac{\text{d}A_{t}}{\text{d}t}\frac{1}{A_{t}}=\alpha_{t}S_{t}\leavevmode\nobreak\ .$ (2.1)
Here, the variable $A_{t}$ denotes total factor productivity. The left-hand side of (2.1) captures the rate at which “ideas” are produced. On the right-hand side, the factor $S_{t}$ measures capital devoted to research. The structural parameter $\alpha_{t}$ is interpreted as research productivity—the efficiency with which inputs are transformed into outputs. [^8] See e.g., Romer, (1990), Jones, (1995), and Aghion and Howitt, (1992), among many others.
The revival of interest in tools, such as industrial policy, to better allocate public resources to research and development is a reaction, in part, to two stylized facts (Council of Economic Advisers,, 2021, 2024; Lindsey and Hammond,, 2020; Hausmann,, 2023). First, the growth rate of research output has been constant (i.e., the left-hand side of (2.1) is constant). Second, the quantity of resources devoted to research (i.e., the factor $S_{t}$) has increased substantially. The linear specification (2.1) then implies that research productivity $\alpha_{t}$ has declined. That is, ideas are getting harder to find. For a review of the relevant literature, see Goldin et al., (2024). Detailed characterization of these facts is due to Bloom et al., (2020).
2.1. Heterogeneity
We propose an alternative reconciliation of these facts. Our hypothesis hinges on heterogeneity. If measured increases in the quantity of research are accompanied by a compositional shift toward unproductive or low-capital research, then constant growth does not necessarily imply declining productivity for capital-intensive research. That is, large increases in the quantity of low-capital-intensity or low-productivity research need not have quantitatively important impacts on growth.
To fix ideas, let the quantity $\mathcal{I}_{t}$ index the set of “research units” available at time $t$. There are two types of research units, high, $\mathsf{H}$, and low, $\mathsf{L}$. Each research unit of type $\mathsf{H}$ has a stock of research capital $S^{\mathsf{H}}_{t}$ and a productivity coefficient $\alpha^{\mathsf{H}}_{t}$. Similarly, research units of type $\mathsf{L}$ have research capital and productivity $S^{\mathsf{L}}_{t}$ and $\alpha^{\mathsf{L}}_{t}$, respectively. Consider the generalized production function
$\frac{\text{d}A_{t}}{\text{d}t}\frac{1}{A_{t}}=|\mathcal{I}^{\mathsf{H}}_{t}|\text{ }\alpha^{\mathsf{H}}_{t}S^{\mathsf{H}}_{t}+|\mathcal{I}^{\mathsf{L}}_{t}|\text{ }\alpha^{\mathsf{L}}_{t}S^{\mathsf{L}}_{t}\leavevmode\nobreak\ ,$ (2.2)
where $\mathcal{I}^{\mathsf{H}}_{t}$ and $\mathcal{I}^{\mathsf{L}}_{t}$ index the set of $\mathsf{H}$-type and $\mathsf{L}$-type research units and ||$\cdot$|| ⋅ | denotes the cardinality operator. That is, relative to the canonical specification (2.1), we entertain heterogeneity across research units in both productivity and capital-intensity.
Suppose that the quantity, productivity, and capital of $\mathsf{H}$-type units is constant over time. Thus, in each period $t$, the first term in (2.2) is given by $|\mathcal{I}^{\mathsf{H}}|\text{ }\alpha^{\mathsf{H}}S^{\mathsf{H}}$, where we have removed the time subscript. Suppose that at time $t=0$, there are no $\mathsf{L}$-type units, and that at time $t=1$, the quantity of $\mathsf{L}$-type units $\mathcal{I}^{\mathsf{L}}_{1}$ becomes non-zero. The impact of this change only affects growth through the factors $S_{1}^{\mathsf{L}}$ and $\alpha_{1}^{\mathsf{L}}$. If either quantity is negligible, relative to the contribution of the $\mathsf{H}$-type units, then there is a negligible impact on growth.
This story is closely related to a strand of the economic growth literature that argues that measured increases in the quantity of research reflect increases in the number of product varieties (Dinopoulos and Thompson,, 1998; Peretto,, 1998; Young,, 1998; Howitt,, 1999). By contrast, we examine the idea that shifts in the relative shares of established varieties can rationalize the same trends.
2.2. Medical Research
The objective of this paper is to test the compositional hypothesis described in Section 2.1. We do so in the context of medical research. We argue that in this setting we are able to decompose research units into coarse types. Moreover, for each type, institutional context allows us to infer absolute and relative productivity and capital-intensity. Measures of quantity can be recovered from existing data. An additional advantage of this setting is that medical research has an intuitive notion of output: improvement in mortality. To keep our focus on the composition of inputs—the right-hand side of (2.2)—we note that it is a well-established empirical regularity that life expectancy has risen at a steady rate since 1840 (see e.g., Oeppen and Vaupel,, 2002; Bloom et al.,, 2020). We proceed assuming that the left-hand side of (2.2) is constant, as is standard in this literature. [^9] We make the assumption that life expectancy is improving at a steady rate, given the stability of these trends for more than 150 years. It is of course unlikely that medical innovations developed in the early 2010s are clearly discernible in recent mortality data, especially given the impact of the COVID-19 pandemic.
Medical research takes many forms. We focus on two types: clinical trials, and other forms of clinical research that aim to inform the provision of medical care. Clinical trials are expensive, time-consuming, highly-regulated experiments, which have as their goal the production of causal and mortality-relevant evidence (Alsan et al.,, 2024; Wouters et al.,, 2020). The contours of the second set, other forms of clinical research, are less precise. This set includes observational studies, meta-analyses and systematic reviews, cost-effectiveness studies, etc. (Food and Drug Administration,, 2024). In contrast to clinical trials, these other study types are less capital-intensive and less productive, where productivity is defined as their efficiency in turning research inputs into changes in mortality (Jones and Podolsky,, 2015; Chavez-MacGregor and Giordano,, 2016; Kones et al.,, 2014). [^10] Of course, one can identify specific studies that deviate from our categorization. See Kones et al., (2014) for discussion of a setting in which, the authors argue, observational data are as informative as randomized controlled trial evidence. However, clinical trials remain the “gold standard” for medical research (Jones and Podolsky,, 2015). Non-randomized evidence is consistently excluded from prescribing guidelines, on the grounds that it provides evidence of lesser quality. Chavez-MacGregor and Giordano, (2016) and Kones et al., (2014) discuss hierarchies of medical evidence in more detail.
3. Identifying Clinical Trials in Publication Data
We construct a census of clinical trials. No existing dataset, used off-the-shelf, is sufficient. Although dollar-denominated expenditures are an obvious metric, available records of investment in clinical trials are self-reported, in some form, by pharmaceutical firms and are essentially impossible to validate independently. [^11] There are two frequently used sources of data on dollar-denominated expenditures. DiMasi et al., (2016) provide one set of estimates of drug research and development costs, based on confidential data obtained from 10 firms. Myers and Pauly, (2019) and Scannell et al., (2012) use research and development expenditure figures published by the Pharmaceutical Research and Manufacturers of America, a trade organization representing the industry. Administrative databases—such as ClinicalTrials.gov—aim to provide an index of trials, but suffer from widespread non-compliance with reporting requirements, with substantial variation in enforcement over time (DeVito et al.,, 2020). Proprietary databases pose two challenges: first, it is, again, difficult to validate their contents and, second, leading databases take as inputs data in public registries, thus adopting the same shifts in reporting over time (see e.g., the list of registry inputs for one frequently-used database, Cortellis,, 2024). For each of these options, there is no symmetric way to identify the quantity of other clinical research.
3.1. Publication Data
Publication data provide a natural alternative. The database PubMed / MEDLINE indexes roughly $34$ million publications, a near universe of published biomedical research. [^12] See Section A.1 for further details on on our treatment of the PubMed data. Scientific publications provide an intuitive proxy for the quantity of research inputs. [^13] The existence of publication bias means that scientific publications do not perfectly capture research inputs. See Andrews and Kasy, (2019) for a discussion. In a setting such as ours—where objects of interest are trends in research over time—we need only assume that changes in publication bias year-on-year do not perfectly offset changes in research investment. Of these $34$ million records, which are clinical trials?
The existing literature takes several approaches. In health sciences—where data on the universe of clinical trials are relevant for the construction of meta-analyses and the development of prescribing guidelines—researchers typically employ over-inclusive methods with high true positive rates. To use one prominent example, Cochrane, a British organization that synthesizes the findings of medical research, developed the “Cochrane RCT Classifier,” a machine learning-based classifier for retrieving randomized controlled trials. The Cochrane approach successfully identifies trials with a 99 percent true positive rate. For every eight true positives, however, it returns 92 false positives. See Thomas et al., (2021) for details.
Approaches used by researchers in social science, instead, largely draw on metadata contained in PubMed. [^14] Oostrom, (2024) and Kao et al., (2023) are two exceptions. Oostrom, (2024) uses data from a meta-analysis of published clinical trials. Kao et al., (2023) collect trial records from the proceedings of scientific conferences. Lichtenberg, (2018) and Bloom et al., (2020) collect all records assigned a “publication type”—in the PubMed indexing process—of “clinical trial.” Feldman et al., (2019) collect all records with a publication type similar to clinical trial (e.g., also including “randomized controlled trial”).
On inspection, these tags are imprecise. Both approaches retrieve large numbers of records that are not, in fact, clinical trials, while excluding potentially relevant records. Figure 1 contrasts counts of published trials constructed in various ways. The light green line (‘Clinical Trial’ NLM Tag) plots the number of publications, in each calendar year, indexed with the publication type “clinical trial.” The teal line (Any NLM Tag) plots counts of records indexed with any of 18 types that are likely to include clinical trials or related medical research. [^15] On inspection, it appears as though the NLM shifted from using the tag “Clinical Trial” to “Randomized Controlled Trial”, as well as more specific tags (e.g., “Clinical Trial, Phase 1”) . Alternative approaches to identifying trials are similarly imprecise. Medical journals increasingly require trials to be registered in open registries, e.g., ClinicalTrials.gov, as a condition of publication. The blue line (Any Registry Tag) counts records that report, in their abstract, an identifier associated with one of the four largest international trial registries. Language in a record’s abstract can indicate that it might be a clinical trial if, for example, it references a “treatment group” and “control group.” The dark blue line (Any Keyword) plots the count of records with such keywords over time. We define the “potential universe” of clinical trials as those publications with any of the following: a publication type variable similar to clinical trial, a registry identifier reported in its abstract, or a keyword reported in its abstract. The purple line (Potential Universe) plots this trend over time. Figure 1 highlights that alternative approaches to data construction, in this setting, yield meaningfully different conclusions about levels, trends, and composition. See Section A.2 for further details on the construction of these series.
Figure 1. Universe of Potential Clinical Trials
Notes: Figure 1 displays counts of the number of clinical trials indexed in PubMed / MEDLINE over time, constructed using alternative search strategies. The National Library of Medicine (NLM) categorizes each publication into a “pubtype.” The light green line (‘Clinical Trial’ NLM Tag) displays the number of publications in the “Clinical Trial” pubtype. The teal line (Any NLM Tag) displays the number of publication whose pubtype is an element of a set of 18 categories likely to include clinical trials. The blue line (Any Registry Tag) gives the number of publications that report, in their abstract, an identifier associated with one of the four largest international trial registries. The dark blue line (Any Keyword) indicates the number of publications whose abstract contains a keyword indicative of a clinical trial. The purple line (Potential Universe) displays the number of clinical trials in the union of the sets of publications identified with the other lines. See Section A.2 for further details.
3.2. Objective
What exactly are we trying to recover when we search for clinical trials? We argue that the object of interest is a census of clinical trials that study the effects of a medicine in human subjects.
Definition 3.1.
The sample of interest is composed of all publications that report the results of a prospective, interventional clinical trial that evaluates the effects of investigational or approved drugs in a setting with exclusively human subjects.
3.1 embeds several restrictions, intended to yield a measure of new investment in information relevant to human-facing medical care. We view this set of studies as the set most likely to generate information relevant to human mortality. Studies involving animals, literature summaries, and re-analyses of existing data, for example, are excluded. We take the “potential universe” plotted in Figure 1 as our baseline sample of potential clinical trials. We further restrict attention to the 1,821,429 records published in or after 2010.161616There is a trade-off between data coverage and quality. Around 2010, a set of policy changes—including trial registration requirements and harmonization in publication formatting—affected trial disclosure in scientific publications. See Laine et al., (2007) and Schulz et al., (2010) for examples of two such changes.
To quantify our ability to identify the clinical trials in this set of publications, we hand-label approximately 3,000 randomly selected publications based on the content of their abstracts. [^17] We develop a custom labelling interface that reduces the time required to label 100 records, for the authors, by a factor of seven. Section B.1 describes this tool and its application to this context in detail. Of the hand-labelled publications, 11.2% meet the criteria for inclusion in our sample. The hand-labelled data are split into three subsets—validation, training, and testing—based on their eventual use.
As a baseline, we assess the performance of several standard machine learning algorithms. The results are displayed in Figure 2. Further details on the construction of this figure are given in Section B.2. Each algorithm is estimated on the hand-labels assigned to the publications in the training and validation sample splits. For feature vectors, we use either TF-IDF embeddings computed in the corpus of abstracts in the hand-labeled sample or the embeddings of each abstract obtained from the SentenceTransformer language model (Reimers and Gurevych,, 2019). We find that the performance of the standard machine learning algorithms is again unsuitable for our application. At a 90% true positive rate, the best performing model identifies 50 true positives for every 50 false positives.
Figure 2. Performance of Standard Machine Learning Methods
Notes: Figure 2 displays the receiver operating characteristic of several machine learning models, trained on two types of embeddings, for classifying whether a publication satisfies the restrictions enumerated in 3.1. The hyper-parameters of each model are determined with 10-fold cross-validation. The models are trained in the training and validation data set. Error rates are estimated in the testing data set. In addition, we display the performance of our, ensemble, fine-tuned large language model in purple. This is the same curve labelled “Ensemble” in Figure 3.
3.3. Prompt Design and Fine-Tuning
We construct a large language model optimized for our task. This is accomplished in three steps. First, we iteratively construct a set of prompts that exhibit good performance when posed to proprietary models—OpenAI’s GPT-3.5 and GPT-4. Second, we extract noisy labels for a large number of publications in our sample by querying the proprietary models. Third, the noisy labels are used to train an off-the-shelf large language model. The resulting model is then used to identify clinical trials in our baseline sample. This process of constructing specialized large language models is referred to as “fine-tuning” (see e.g., Taori et al.,, 2023, for a prominent application of this idea).
3.3.1. Prompt Design
We devise three types of prompts and measure their performance on samples of abstracts taken from our validation dataset. We measure model performance with the hand-labeled data, conduct error analyses, and iteratively revise prompts. When modifications to the prompt cease to yield improvements in performance, we finalize prompts best suited for GPT-3.5 and GPT-4, respectively. Section B.3 provides more detail on this process. Figure 3 displays estimates of the true and false positive rates computed in the test data, for both proprietary models, with black dots.
Figure 3. Receiver Operating Characteristic
Notes: Figure 3 displays estimates of the receiver operating characteristic of several models used to classify the publications indexed by PubMed / MEDLINE according to whether they satisfy the restrictions enumerated in 3.1. All metrics are computed in the testing split of the hand-labelled publications. Estimates of the true positive rate and false positive rate of OpenAI’s proprietary models GPT-3 and GPT-4 are indicated with black dots. The curves give the performance of four fine-tuned, open-source large language models, in addition to an ensemble model. The blue dots indicate the true positive rate and false positive rate of the models used to construct the “Conservative,” “Moderate,” and “Liberal” samples of clinical trials.
GPT-4 performs remarkably well. For every 71 true positives, GPT-4 identifies 29 false positives, while achieving a true positive rate of approximately 0.974. This represents a substantial improvement over existing methods. However, practical and substantive considerations mitigate the applicability of proprietary models for classification of the complete baseline sample. Practically, it is—and likely will remain—prohibitively expensive to deploy GPT-4 at this scale. [^18] We incur a cost of roughly \$4,500 to extract noisy labels for 64,000 abstracts. Extrapolating this figure to the full sample of 1.8 million abstracts gives a price of approximately \$130,000. Substantively, proprietary models are black boxes. Their details and substance are not public and are known to change at regular intervals.
3.3.2. Fine-Tuning
We compute noisy labels for 64,000 randomly selected publications in our baseline sample using the best performing prompts for both GPT-3 and GPT-4. These noisy labels are used as data to train off-the-shelf language models from two architecture classes: (1) RoBERTa (Liu et al.,, 2019) and (2) BioMedBERT (Gu et al.,, 2021). Both models are available in either “baseline” or “large” varieties. [^19] There are many open-source large language models (e.g., LLaMa, Mistral, Pythia). Many are several orders of magnitude (e.g., 7-70 billion parameters) larger than RoBERTA ($\sim$ 350 million parameters). We selected RoBERTA and BioMedBERT via trial-and-error that included testing the performance of these larger models. In our setting, RoBERTA outperforms LLaMA (both 7 and 70 billion parameter versions, trained with QLoRA). Both models exhibited comparable performance, but RoBERTa was substantially faster and simpler to train. The BioMedBERT model is an instance of a BERT model (Devlin et al.,, 2019), that has itself been fine-tuned on the abstracts of the publications indexed in PubMed. We systematically vary model architecture, size, and pre-training regimen. We observe no large differences in performance stemming from these hyper-parameters. Sections B.4 and B.5 give further details the fine-tuning process and the results of these experiments, respectively.
3.4. Performance
Given the text of an abstract, our fine-tuned language models output a probability that the publication satisfies the restrictions enumerated in 3.1. Publications whose probabilities fall above a chosen threshold are classified as belonging in our sample. Figure 3 displays estimates of the true and false positive rates, computed with the test data, as we vary this threshold for both sizes of the RoBERTa and BioMedBERT models. [^20] Each of the models whose results are displayed in Figure 3 are trained with noisy labels obtained with GPT-4. We document in Section B.5 that models trained with noisy labels obtained from GPT-3.5 or with the 1000 hand-labeled records in the training data perform significantly worse. The figure additionally displays the performance of an ensemble model estimated in the training data. This model is obtained by estimating a logistic regression of the hand-labels on the probabilities output by all four models displayed Figure 3. The ensemble model is used to produce our final sample.
The fine-tuned models are able to match the performance of GPT-4 in the test data. We choose three thresholds according to the stringency with which they enforce the sample restrictions. The estimated true and false positive rates associated with these points are displayed in Figure 3 and labeled as “Conservative,” “Moderate,” and “Liberal.” Our preferred sample is associated with the conservative threshold. [^21] The liberal model may be particularly useful for conducting literature reviews, where a near-perfect true positive rate is needed. For every 82 true positives, the conservative model identifies 18 false positives. [^22] In the test data, the conservative model assigns incorrect labels to 27 of 993 papers. We conduct an error analysis. See Section B.6 for further details. In 13 cases, there is a clear error. In 14 cases, however, errors are associated with records that were difficult to categorize for a human labeller. For example, PubMed record 32737793 is flagged as satisfying 3.1 with all three thresholds, but is a literature review. By contrast, PubMed record 27880726 was categorized as satisfying 3.1 twice by a human labeller. It is excluded from both the conservative and moderate model-generated samples. On inspection, the abstract does not explicitly state that the study enrolled human subjects, but hints that it may have been conducted in an animal model. Review of the associated full text confirms that this study did, in fact, enroll only rats. We report results associated with the moderate and liberal thresholds in the appendix as tests of robustness. Our final, conservative, sample consists of 152,027 publications classified as satisfying the restrictions enumerated in 3.1.
Appendix A compares the contents of our sample to the counts of potential clinical trials plotted in Figure 1. Although certain elements of PubMed metadata—including the union of all NLM tags—capture many of the records in our final sample, we confirm that they miss many records that we flag as trials and include many records that do not satisfy 3.1. No combination of existing search strategies, then, classifies records with the same accuracy or precision as our final sample.
4. Medical Research as a Productivity Indicator
Figure 1 suggests that, across metrics, there has been a substantial increase in the quantity of medical research over time, including since 2010. This is consistent with documented increases in the quantity of scientific papers in Bloom et al., (2020) and Park et al., (2023), as well as economy-wide increases in research effort reported in Goldin et al., (2024).
4.1. Decomposing Medical Research
We decompose this pattern, into changes attributable to high-productivity, high-capital investments—clinical trials—and changes attributable to other forms of clinical research. Our census of clinical trials provides measures for the first category. Although there are many reasonable ways to construct samples of other clinical research, our preferred approach defines this category as those scientific publications that cite clinical trials. [^23] Alternative approaches to construction of this sample, as in Figure 1, also suggest an increase in the quantity of non-clinical trial research. We prefer to use this citation-based definition because it allows for consistency in sample definition. Other approaches that rely on keyword searches, NLM tags, etc., as Section 3 and its appendices make clear, capture different sets of records over time. These are records of research investment that necessarily engage with findings from trials (such as meta-analyses and re-analyses of existing data), but, as discussed in Section 2, are essentially always lower productivity and less capital-intensive.
Figure 4 documents that the quantity of clinical trials—the black line—has remained constant over the past ten years. In contrast, other clinical research has increased substantially. Observe that we display the percent growth in the quantity of clinical trials in our sample published in each calendar year, alongside the growth in the set of papers that cite clinical trials. Each series is normalized to begin at zero. To account for concerns about citation delays and truncation, we construct various measures of papers that cite clinical trials, using 2–6 year citation counts. Two-year citation counts, thus, are available starting in the year 2012 (two years into our sample) and count publications that cite clinical trials published in the preceding two years. [^24] Figure C.3 and Figure C.4, displayed in the supplemental appendix, document that this trend is robust to alternative sample definitions, using either our liberal or moderate samples, and to citation-weighting count data. This presentation masks substantial differences in the levels of these series. There are 10,903 clinical trials in our census in 2010, and 73,253 publications that cite clinical trials (using two-year citations) in 2012.
Figure 4. Growth in Clinical Research, Stability of Clinical Trials
Notes: Figure 4 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year. Each series is reported in terms of the percent change relative to its initial value. The sample of published clinical trials is constructed with the conservative model. To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
4.2. Stable Composition of Clinical Trials
To contextualize the trends displayed in Figure 4, it is helpful to return to the simple model from Section 2. We document that the quantity of capital-intensive, high-productivity research inputs remains unchanged over our period of interest, while low-capital, low-productivity inputs increase. As changes in outputs in this sector (mortality) are assumed to remain constant (see Section 2), there are thus three ways of reconciling these facts in our stylized framework: (i) other clinical research is sufficiently low capital or low productivity, across the board, as to have essentially no impact on growth, (ii) increases in other clinical research have, heterogeneously, low capital or low productivity, or (iii) clinical trial research productivity is declining over this time period, fully offsetting the growth in other types of research. If either Case (i) or Case (ii) were an accurate characterization, rising observed research effort—increases in the number of scientific publications or the number of records in PubMed that the previous methods have identified as clinical trials—need not imply diminishing returns to high-capital investment.
Our data allow us to reject the hypothesis that we are in Case (iii). Figure 5 documents that—on four dimensions—measures of trial productivity, across clinical trials, have remained stable over the past decade. Panel A of Figure 5 considers the share of publications in our census with the following characteristics: any funding from a government agency (purple), any (three-year) citations from a leading journal in medicine (teal), or any citations from a leading journal conditional on having public funding (green). [^25] We designate a publication as having public funding if it appears in the National Institutes of Health RePORTER data, as linked to a funded research grant, or if PubMed reports a source of research funding, including funding from non-U.S. agencies. To designate a set of the “top” journals in medicine, we follow Angrist et al., (2020) We collect all citations originating from the “trunk journals” in medicine, the Journal of the American Medical Association and the New England Journal of Medicine, to records in PubMed between 2010 and 2022. We designate a journal as being “leading” of it received at least 100 citations from a trunk journal over this time period. This yields a list of 84 journals. See Section A.3 for details. Appendix C reproduces this figure using five-year citations. The trend is unchanged. The trends in Panel A communicate two findings. First, all three proxies for research productivity are stable, essentially unchanging, over our time period. Second, nearly 55 percent of clinical trials are never cited by a leading journal, suggesting substantial, but time-invariant, heterogeneity across trials.
Figure 5. Stability of Heterogeneity
Panel A: Public Funding and Low-Quality Research
Panel B: The Composition of Quality over Time
Notes: Figure 5 illustrates the stability of the heterogeneity in the composition and quality of published clinical trials between 2010 and 2019. Panel A displays three time series: the proportion of clinical trials that are publicly funded, the proportion of clinical trials that are ever cited by a leading journal, and the proportion of publicly funded clinical trials that are ever cited by a leading journal. Panel B displays two heat maps measuring the distribution of citations received by clinical trials in each calendar year from leading journals and from all publications, respectively. The $y$-axis has been stretched to elongate the right-tail of the citation distribution. Colors are displayed in a log scale. For both panels, the sample of published clinical trials is constructed with the conservative model.
These proxy variables capture slightly different conceptual objects. Interpreting each, separately, provides suggestive evidence on different aspects of the research production function. Stability in public funding suggests minimal changes, at least in the cross section, in allocation of resources to clinical trials over time. [^26] NIH RePORTER data allow us to tab total expenditures associated with grants acknowledged in these publications. These grant-based expenditure measures, too, are stable. Note, however, that linking research grant dollars to specific scientific papers is an imprecise exercise. See Li, (2017) for an extended discussion of this disconnect. Stability in citation patterns, in contrast, suggests that the “usefulness” of information produced by trials also remains stable.
Panel B of Figure 5 provides a more granular view of this heterogeneity. For each publication in our census, we collect citations from publications in leading journals and from all of PubMed. We plot the total (three-year) citation counts to publications at each citation quantile over time. [^27] Appendix C presents these results using alternative cuts of our data and five-year citations. The patterns remain unchanged. Grey areas at the bottom of each plot represent publications that receive zero citations. This distribution is stable over our time period. [^28] Reassuringly, Panel B indicates that this exercise does capture large shifts in the research environment: citation counts appear especially high in the later years of data, when three-year citation counts (beginning for year 2017) capture research written during the COVID-19 pandemic. Specifically, we document that the tails of the citation distribution are unchanging. Observe that we elongate the y-axis, to highlight the right-tail. Citations received at the top of this distribution are constant across periods. The left-tail is similarly stable: the share of records receiving zero citations is essentially constant, year-on-year.
This result is consistent with the idea that, for clinical trials, the right-hand side of the production function, (2.1), in our simple framework—i.e., $\alpha_{t}S_{t}$—is both heterogeneous and stable over time. For this pattern to be consistent with a systematic decrease in the research productivity of clinical trials, it must be the case that there is an exactly offsetting increase in the distribution of research capital. Although we cannot reject that there are shifts in upstream, basic inputs into clinical trials or large changes in factor prices during this period, there is little credible evidence consistent with either during this time period. A small number of papers examine frictions in markets for basic research. See, for example, Hill and Stein, (2021) and Myers, (2020). To our knowledge, this literature has not provided evidence of substantial shifts in these frictions over time, which in turn would increase the cost or decrease the productivity of basic science. [^29] Anecdotal evidence suggests that there may have been substantial decreases in the cost of basic science during this time period. See Scannell et al., (2012). On factor prices, as discussed in Section 3, nearly all estimates of research and development costs are self-reported by the pharmaceutical industry. Any measured increases in these expenditures that cannot be independently verified should be viewed with caution, especially in light of our findings of stability.
More generally, any type of cancellation—an argument against the conclusion that research productivity across clinical trials is stable—would appear to be a knife-edge case. In our view, it is far more likely that the distributions of both research productivity and research capital within clinical trials have been constant over this time period.
5. Discussion
All research investments are not alike. Certain investments are more likely to yield innovations that have welfare-relevant impacts. Others, while potentially valuable contributions to a research ecosystem, may have insignificant impacts on growth. We argue that this type of heterogeneity is quantitatively significant in medical research. We use an open-source, fine-tuned large language model to construct a novel census of clinical trials from 2010 to the present. We contrast stable trends in clinical trials—in quantity, quality, and composition—with substantial growth in the quantity of other forms of medical research. Our findings rationalize, in this field, two stylized facts about research investments and innovative output: the quantity of research inputs has grown substantially, while improvements in mortality continue at the same stable rate.
A second dimension of heterogeneity is helpful in contextualizing our findings: geography. Figure 6 displays trends in the annual quantity of published clinical trials, and papers citing clinical trials, disaggregated by the location of the first-listed author. [^30] In most domains of scientific and medical publication, last-listed authors are senior investigators. First-listed authors are typically junior investigators. Agha and Molitor, (2018) observe that, for large-scale clinical trials, first-authors are more likely to be the principal investigator. In our context, the first- and last-author have the same listed country for 85 percent of clinical trial records and 91 percent of non-clinical trial records. We collect details on author location from Clarivate Analytics’ Web of Science. See Section A.4 for further details. In our sample, the top four producers of published clinical trials, and papers citing clinical trials, are, in order, the United States, China, Germany, and Japan. Clinical trial investment—across geographies—is stable. [^31] Note that Panel A of Figure 6 displays measurements of the location of clinical trial publication authors, not the location of clinical trial sites. We cannot rule out that, for example, trials authored by researchers with United States mailing addresses were conducted elsewhere. Thus, we do not interpret these as facts on clinical trial “offshoring.” See Petryna, (2007) and Durvasula, (2023) for longer discussions of the geography of clinical trial sites. Roughly 30 percent of published clinical trials originate in the United States. This share is unchanging across periods. By contrast, Panel B indicates a large shift in the medical research ecosystem. Since 2013, there has been a small increase in the quantity of other medical research published by authors in the United States. This small change is dwarfed by increases—on the order of 30-50 percent—in the quantity of other medical research from China and from countries outside of the top four producers of medical research.
Figure 6. Composition of Medical Research Across Countries
Panel A: Clinical Trials Panel B: Papers Citing Clinical Trials
Refer to caption Refer to caption
Notes: Figure 6 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, by country. The sample of published clinical trials is constructed with the conservative model. To address truncation, we report the number of publications that cite clinical trials published in the preceding three years.
These compositional facts are, in many ways, unsurprising. In recent years, editorials in prominent journals in medicine have observed that changes in the incentive structures for physicians, scientists, and even medical students, across countries, have led to “an extreme of quantity at the expense of quality” in medical publishing (Siegel et al.,, 2018). [^32] See, as one example, program certification requirements, which describe publication expectations for medical students and clinical faculty, from the Accreditation Council for Graduate Medical Education in Accreditation Council for Graduate Medical Education, (2017). Recent work documents a substantial increase in the quantity of meta-analyses and systematic reviews—summaries of existing clinical trials—driven by scientists from outside of the United States (Ioannidis, 2016a, ). In parallel, there has been a rapid increase in the number, and scale, of “mega-journals”—scientific journals published frequently, with a large and growing number of pages (Ioannidis et al.,, 2023).
Nonetheless, substantial constraints—financial, ethical, and practical—keep firms, physicians, and scientists from increasing the frequency of clinical trials. Clinical trials are infrastructure-intensive. Trials are run at specific sites, which must identify patients who satisfy all relevant eligibility criteria and monitor them for periods ranging from six months to twenty years (Piantadosi,, 2024). Best available estimates suggest that trials can cost on the order of 500 million to one billion dollars to complete (Qiao et al.,, 2019). Perhaps most importantly, sites must be capable of producing data sufficient to persuade regulators. Certain regulators require that sponsors of new drug applications submit trial evidence collected from a domestic population. [^33] Until December 2023, Japan required domestic phase I clinical trials of drugs developed overseas before Japanese individuals could participate in international phase III trials for pharmaceutical regulatory approval. This policy reflected, in part, differences in disease burden in Japan, relative to countries where drugs are routinely tested. This may explain why Japan is a top producer of clinical trials over our sample period. See Namba et al., (2024) for one discussion. In recent years, however, the U.S. Food and Drug Administration (FDA)—responding to concerns from patient groups—has indicated a limited willingness to approve drugs on the basis of evidence collected from exclusively foreign sites (for a longer discussion, see Alsan et al.,, 2024). In one especially high-profile example, a cancer drug was rejected by the FDA after being tested exclusively in China (Kolata,, 2022). As long as the U.S. market continues to have outsized value for pharmaceutical firms—a consequence of the especially high prices paid by American consumers—it may be unsurprising to find that evidence disproportionately originates in the United States. Observational studies, meta-analyses, and case reports are, by contrast, cheap and not subject to the same financial and regulatory pressures. Digital tools, including search engines and language models, will likely further depress these costs.
Existing work documents patterns that are not inconsistent with those in this paper. However, we offer an alternative interpretation. Like Park et al., (2023) and Bloom et al., (2020), we find a substantial increase in the size of the medical research ecosystem. Park et al., (2023) document that this increase in scale brings with it a decline in the average importance of published scientific articles. Technically, we find the same. As increases in non-clinical trials outpace (non-existent) changes in clinical trials, the average productivity of studies in Figure 3 mechanically decreases. In debates on productivity, however, we argue that this measure of central tendency is not the policy-relevant object. To be sure, these findings raise a host of other important questions—about misaligned incentives, wasteful spending, and the productive efficiency of science. But canonical frameworks, such as Romer, (1990), center on the simple idea that it is high-intensity investments that generate meaningful shifts in outcomes. An explosion of low-capital research that has a negligible impact on outcomes, then, is not cause for concern from a growth perspective. That is, growth-oriented studies may prefer to focus—like much of the research on science and innovation—on the right-tail of the productivity distribution (e.g., Kelly et al.,, 2021).
Our findings do not speak to shifts in upstream, basic science inputs into clinical trials, or to shifts in factor prices. In our view, there is little credible evidence indicating changes on either dimension during this time period. This setting allows us, only, to note that any shift in inputs or factor prices must be perfectly offset by a countervailing force, to account for the complete stability of clinical trial counts, citations, journal placements, and reliance on public funding in our data.
Our findings, also, apply to roughly one decade: 2010 to 2022. We impose this restriction because policy changes in the early 2000s substantially shifted incentives for both publication and standardized formatting of clinical trial results. See Laine et al., (2007) and Schulz et al., (2010) for details. Given our methods, these changes mean that data construction over a longer period risks decreasing the quality of our data. Thus, we cannot speak to trends that precede this period. That clinical trial investment is remarkably stable in the past decade, however, is likely the relevant consideration for contemporary public policy.
What our findings, and the simple framework borrowed from the growth literature, do suggest is that there is likely value in efforts to increase the productivity and decrease the costs of clinical trials. A long-standing discussion in medicine and statistics, spurred by Altman, (1994), centers on the argument that “we need less research, better research, and research done for the right reasons.” Our data indicate that nearly half of all clinical trials published in our sample are never cited by leading medical journals, and roughly 15 percent are never cited at all. This heterogeneity is consistent with findings in Ioannidis, 2016b and van Zwet et al., (2023), who imply that there may exist trade-offs between quantity and data quality in this setting. If increasing the scale of research infrastructure is a policy objective, such efforts should invest in tools, like patient registries, that enable the production of high-quality, representative information (Alsan et al.,, 2024; Durvasula,, 2023).
Supplemental Appendix to: Medical Research as a Productivity Indicator [^1] Date: May 12, 2024
$\star$ Stanford Department of Economics and Stanford Law School, maya.durvasula@stanford.edu
$\dagger$ Stanford Department of Computer Science, eyuboglu@stanford.edu
$\ddagger$ Stanford Graduate School of Business, ritzwoll@stanford.edu
Maya M. Durvasula Sabri Eyuboglu David M. Ritzwoller
Contents  appendix.Asubsection.A.1subsection.A.2subsubsection.A.2.1subsubsection.A.2.2subsubsection.A.2.3subsubsection.A.2.4subsubsection.A.2.5subsection.A.3subsection.A.4appendix.Bsubsection.B.1subsection.B.2subsection.B.3subsection.B.4subsection.B.5subsection.B.6appendix.Cappendix.D
Appendix A Publication Data
In this Appendix, we detail our treatment of the PubMed / MEDLINE database (hereafter, “PubMed”). We give a general overview of the PubMed data in Section A.1. In Section A.2, we describe the process of identifying our initial sample of abstracts, i.e., the “Potential Universe” displayed in Figure 4. We discuss the process of determining the set of leading medical journals in Section A.3.
A.1. The PubMed / MEDLINE Database
Our version of the PubMed / MEDLINE database contains roughly $34$ million records and is current through December 2022. These records were constructed by parsing bulk MEDLINE XML files. [^34] We are grateful to Heidi Williams for sharing this processed data. Technically, PubMed and MEDLINE are different products. MEDLINE, a subset of PubMed, is the U.S. National Library of Medicine’s (NLM) bibliographic database, which contains references to journal articles in life sciences, with a primary focus on biomedicine. A committee at the NLM determines the set of indexed journals, meaning that only journals that meet certain quality and content standards are indexed. In practice, this means that so-called predatory journals are excluded, as are pre-prints and non-peer reviewed articles. Informal conversations with staff at the NLM suggest that MEDLINE should contain the universe of peer-reviewed publications in legitimate journals. PubMed includes a broader set of records, including pre-prints and publications deposited through alternative processes. Official documentation for MEDLINE is available at the link: https://www.nlm.nih.gov/databases/download/pubmed_medline_documentation.html. Throughout the text, and from this point forward in the Appendix, we refer to this database as “PubMed.”
A natural question for researchers who rely on these data as a census of research investments is whether PubMed records are complete and accurate. To our knowledge, there is no paper that reports such validation exercises for each field of PubMed. We focus on those data elements most relevant to our work. There are 2,335,653 records in PubMed that have no associated year of publication, which we drop from our baseline sample. There are 16 records, across all years of data, that are not indexed with tags describing their contents (NML tags). From 2010 forward, roughly 92 percent of records have associated abstracts. We randomly inspect roughly 150 records with no abstracts. None correspond to publications that satisfy our definition of a clinical trial.
We assume, throughout this paper, that records with missing citation data have zero associated citations. We confirm that this is generally accurate in two ways. First, we search for Google Scholar records corresponding to roughly 100 randomly selected records with zero citations. In each case, Google Scholar indicates that the paper has no more than two citations. Second, we link records in PubMed to Clarivate Analytics’ Web of Science database. We compare citation counts constructed using information in the two databases. We find that the two measures have a correlation of 0.9.
Researchers interested in different cuts of the data may need to conduct additional validation exercises. Figure 1 highlights that the frequencies of certain NLM tags have changed over time, often sharply. Researchers should account for such changes in any use of these flags. For other assessments of the completeness of certain PubMed fields, see Durvasula et al., (2021) and Ouellette and Sampat, (2024).
A.2. The Universe of Potential Clinical Trials
We construct our sample using records drawn from the universe of publications indexed in PubMed. The initial sample consists of 34,957,127 unique publications indexed in PubMed, as of 15 April 2023. We drop 2,335,653 records that are missing information on publication year, to yield a base sample of 32,621,474 records. From 2010 forward, at least 92 percent of scientific publications published in each year have associated abstracts. We also drop 14,179 publications with publication year 2023, as we have incomplete data for 2023.
A.2.1. NLM Tags
The National Library of Medicine (NLM) assigns each publication a ‘pubtype.’ In the entirely of PubMed, there are 16 records missing pubtype tags. To the best of our knowledge, there have been no efforts to validate the PubMed indexing process used to generate these flags. We flag records with each of the following pubtype (or associated unique identifiers, reported in the field ‘pubtypeUI’) tags:
Adaptive trial; Clinical conference; Clinical study; Clinical trial; Clinical trial protocol; Clinical trial, Phase 1; Clinical trial, Phase 2; Clinical trial, Phase 3; Clinical trial, Phase 4; Comparative study; Controlled clinical trial; Equivalence trial; Evaluation study; Observational study; Pragmatic clinical trial; Randomized controlled trial; Twin study; Validation study.
These categories are chosen to include all pubtypes with the potential to contain a publication satisfying the restriction enumerated in 3.1. In particular, we follow Feldman et al., (2019), who use multiple pubtype fields to retrieve a sample of records. Here, two authors reviewed the list of potential pubtypes to identify categories likely to include records of interest. Table A.1 reports the frequency of each NLM tag across all records in PubMed and subset to those published after 1 January 2010.
Table A.1. Composition of PubMed by NLM Tag
A. All records B. 2010-2022 C. Conserv. Sample
Frequency % Frequency % Frequency %
adaptive trial 36 0.00 36 0.00 11 0.01
clinical conference 7,045 0.02 1,724 0.01 12 0.01
clinical study 5,053 0.02 5,050 0.04 641 0.42
clinical trial 498,722 1.53 79,029 0.55 13,969 9.19
clinical trial protocol 9,542 0.03 9,542 0.07 124 0.08
clinical trial, phase 1 23,704 0.07 14,139 0.10 10,951 7.20
clinical trial, phase 2 37,623 0.12 22,749 0.16 17,166 11.29
clinical trial, phase 3 20,597 0.06 15,647 0.11 11,112 7.31
clinical trial, phase 4 2,251 0.01 1,790 0.01 1,250 0.82
comparative study 1,752,380 5.37 424,681 2.97 22,521 14.82
controlled clinical trial 88,132 0.27 13,839 0.10 2,181 1.43
equivalence trial 1,047 0.00 1,047 0.01 505 0.33
evaluation study 243,290 0.75 123,897 0.87 998 0.66
observational study 127,461 0.39 127,370 0.89 5,996 3.94
pragmatic clinical trial 2,112 0.01 2,112 0.01 253 0.17
randomized controlled trial 549,711 1.69 284,774 1.99 79,907 53.57
twin study 9,030 0.03 4,964 0.03 14 0.01
validation study 101,817 0.31 62,261 0.44 495 0.33
$N$=32,621,474 $N$=14,316,494 $N$=151,997
Notes: Table A.1 reports the frequency and percentage of records indexed in PubMed that have been categorized by the NLM as falling into each of 18 categories. These categories are selected for their potential to contain a publication satisfying the restriction enumerated in 3.1.
A.2.2. Clinical Trial Registry Identifiers
In 2004, the International Committee of Medical Journal Editors recommended that research journals decline to publish outcomes associated with trials not pre-registered in some repository. Laine et al., (2007) summarizes these policies, based on a 2007 revision, in more detail. The U.S. Food and Drug Administration Amendments Act of 2007, Section 801, mandates registration of all clinical trials regulated by the FDA in ClinicalTrials.gov. Many countries have adopted similar guidance. Several countries and international organizations now maintain registries of trials.
Registry identifiers are distinctive strings of letters and numbers. We flag records that include clinical trial registry identifiers in their abstract text. In particular, we search for records containing acronyms associated with the following registries:
ClinicalTrials.gov (NCT); European Union Drug Regulating Authorities Clinical Trials Database (EUDRACT); International Traditional Medicine Clinical Trial Registry (ISRCTN); Australian New Zealand Clinical Trials Registry (ACTRN).
We note the potential to overcount records using these searches. We focus on instances where each trial identifier prefix is followed by numbers, letters, or punctuation (e.g., NCT12345, ISRCTN: 12345). However, we may collect records that include these characters in other settings (e.g., the world distinct). Table A.2 reports the frequency of each registry identifier flagged across all records in PubMed and subset to those published after 1 January 2010. Figure 1 plots the number of trials with any registry identifier over time. Reassuringly, we observe registry identifiers in PubMed data beginning around 2010, as registration mandates were implemented, and find a steady, smooth increase over time.
Table A.2. Composition of PubMed by Trial Registry Identifiers in Abstract Text
A. All records B. 2010-present C. Conserv. Sample
Frequency % Frequency % Frequency %
NCT 97,923 0.30 94,279 0.66 30,656 20.17
EUDRACT 2,835 0.0 2,816 0.00 1,564 1.03
ISRCTN 12,261 0.03 11,325 0.07 1,275 0.84
ACTRN 5,818 0.01 5,656 0.03 553 0.36
$N$=32,621,474 $N$=14,316,494 $N$=151,997
Notes: Table A.2 reports the frequency and percentage of records indexed in PubMed that include a string associated with a clinical trial registry identifier in their abstract text. That is, the first row counts the number and percentage of publications that contain the string “nct.”
A.2.3. Keywords
We flag records that include any keyword likely to indicate that the record reports the results of a clinical trial in their abstract text. We selected these keywords after several reviewing roughly 200 abstracts flagged through the pubtype process, described above.
These keywords are:
Randomized; Controlled trial; Control trial; Clinical trial; Treatment group; Control group; Intervention; Clinical study.
Table A.3 reports the frequency of each keyword flagged across all records in across all records in PubMed and subset to those published after 1 January 2010.
Table A.3. Composition of PubMed by Keywords in Abstract Text
A. All records B. 2010-present C. Conserv. Sample
Frequency % Frequency % Frequency %
randomized 550,165 1.69 366,378 2.56 65,637 43.18
controlled trial 107,935 0.33 80,789 0.56 11,535 7.59
control trial 6,049 0.02 4,846 0.03 371 0.24
clinical trial 129,510 0.40 92,035 0.64 17,644 11.61
treatment group 44,627 0.14 27,826 0.19 4,296 2.63
control group 447,901 1.37 286,250 2.00 14,245 9.37
intervention 629,237 1.93 463,373 3.24 10,460 6.88
clinical study 32,652 0.10 18,589 0.13 2,411 1.59
$N$=32,621,474 $N$=14,316,494 $N$=151,997
Notes: Table A.3 reports the frequency and percentage of records indexed in PubMed that include each of a collection of keywords in their abstract text. That is, the first row counts the number and percentage of publications that contain the string “randomized.”
A.2.4. Intersection
There are 3,925,958 records with at least one of the following attributes: a clinical-trial indicative NLM tag; a clinical trial registry identifier; a clinical-trial indicative keyword in the abstract text. The three categories overlap. Table A.4 records the size of the overlap of each category. In the main text, we restrict attention to this sample for the years 2010-2022. This sample includes 1,821,429 publications.
A.2.5. Novel Census
We add columns to Tables A.2, A.3, A.1, and A.4 that document the frequency of each flag in the data constructed for this paper (we use the conservative sample). Several facts are worth noting. First, Table A.2 suggests that—between 2010 and 2022—less than 25 percent of records identified as reporting the results of a clinical trial reported a registry identifier in their abstract. Although, in principle, registry identifiers may be reported in publication full-texts, rather than abstracts, the low frequency of identifiers in our sample suggests non-compliance with requirements, consistent with patterns in ClinicalTrials.gov documented in DeVito et al., (2020). Second, Table A.1 sheds light on the differences in trial composition that we observe in Figure 1: NLM tags that, ostensibly, capture clinical trials have little overlap with one another and do not fully capture the records in our sample. Of the roughly 150,000 records in our census, approximately 100,000 could be identified using NLM tags for “clinical trial” (and its variants) and “randomized controlled trial.” Observe, however, that use of such flags captures a much larger number of records than one would intend to include. Finally, Table A.4 suggests that although various text-based flags that may indicate a record is a clinical trial, in the sense of 3.1, are correlated, there is less overlap than one might expect.
Table A.4. Overlap of Publication Attributes
Records with any: that have any: A. All records B. 2010-present C. Conserv. Sample
Registry ID 116,564 111,837 33,040
NLM tag 81,894 78,366 26,885
Keyword 73,164 70,414 19,103
NLM tag 2,858,924 1,804,081 124,260
Registry ID 81,894 3,528 26,885
Keyword 515,755 220,681 67,153
Keyword 1,564,659 1,044,461 91,349
Registry ID 73,164 70,414 19,103
NLM tag 515,755 295,074 67,153
Notes: Table A.4 reports the size of the intersection between three sets of records indexed in PubMed. These categories are defined by the properties: Possession of a clinical-trial indicative NLM tag, inclusion of a clinical trial registry identifier in abstract text, and inclusion of a clinical-trial indicative keyword in abstract text. These properties are defined in Sections A.2.1, A.2.2, and A.2.3, respectively.
A.3. Determining the Set of Leading Medical Journals
Angrist et al., (2020) propose a strategy to identify leading journals in a field, which we adapt to this context. In medicine, the “trunk journals,” per Angrist et al., (2020), are the Journal of the American Medical Association and the New England Journal of Medicine. We collect all citations originating from these journals to records in PubMed between 2010 and 2022. Angrist et al., (2020) collect initial journal lists for their citation exercise of the journal is one of the top fifty most cited by a trunk journal. As medicine includes a large number of subfields, we instead collect a list of journals that receive at least 100 citations from a trunk journal over this time period, which yields a list of 84 journals.
A.4. The Web of Science
We collect supplementary information about papers in our sample from Clarivate Analytics’ Web of Science. [^35] We use a copy of the Web of Science licensed to Stanford University. Web of Science data allow us to examine more detailed information about each author in our sample. While PubMed includes some details on authors, including limited institutional affiliation and address information, the Web of Science provides standardized author addresses for roughly 75 percent of observations in our sample.
For each paper in our sample, we extract countries from available mailing addresses for first- and last-listed authors. First- and last-authors have the same listed country for 85 percent of clinical trial records and 91 percent of non-clinical trial records. Thus, we impute to each record the country associated with the first-listed author. In Figure 6, we plot geographic trends over time.
Appendix B Prompt Design, Fine-Tuning, and Performance Assessment
In this appendix, we detail the construction of the sample of clinical trials studied in Section 4. We begin in Section B.1 by describing the process we used to hand-label a sample of publications according to whether they satisfy the restrictions enumerated in 3.1. In Section B.2, we document poor performance of several approaches for classifying publications according to whether they belong in our sample based on standard machine learning methods. Section B.3 overviews the process of constructing prompts that perform well when used to query proprietary language models. Details associated with fine-tuning an open-source language model are given in Section B.4. Additional measurements of the performance of our model are given in Section B.5.
B.1. Hand-Labeling
A standard concern when using large language models is the quality of the output. We hand-label a set of 3,000 abstracts, selected randomly from our candidate set of records, which allow us to quantify the performance of our classification procedure.
We devise an interface for abstract labeling, which allows us to review records quickly. Figure B.1 provides an example of this interface. A labeller reads in a set of publication records and is shown one PubMed identifier (PMID) and one abstract at a time. The labeller can select that the abstract satisfies our sample criteria, or else can indicate a reason that the abstract should be excluded. The buttons for exclusion mirror the inclusion and exclusion criteria in 3.1, using short-hand convenient for labellers. We exclude records that report, only, the protocol for a clinical trial without results (“protocol”), that report the results of an observational (“observational”) or retrospective (“retrospective”) study, that report the results of a study involving animals (“animal”), that do not include human subjects (“no human subjects”), [^36] In practice, “animal” and “no human subjects” catch different sets of records. We flag records that include any animal with “animal.” Many studies in our sample of candidate records occur in vitro—in test tubes or other laboratory-based settings. There, no living subjects of any kind are enrolled and the “no human subjects” button is used. that do not study a drug (“no drug”), that report the results of a specific clinical case or patient experience (“case report”), or that summarize the findings of existing work (“meta-analysis”). By our own estimates, the labelling interface increased the speed of labelling by a factor of seven. Each record was reviewed, and labelled, twice.
The hand-labeled data are split into three subsets—validation, training, and testing—based on their eventual use. Splits are assigned randomly. We assign 1000 records to a test set, 1000 to a validation set, and 1082 to a training set. Of these records, we assign labels only to records that have abstracts. Our final training data includes 1082 records, our validation dataset includes 1000 records, and our test set includes 993 records. The loss of seven records in our test set reflects a coding error. We add additional records, randomly selected, to the training and validation datasets when abstracts are missing. We do not add extra records to the testing dataset.
Figure B.1. Abstract Labeling Interface
Notes: Figure B.1 displays a screenshot of the application we developed to hand-label abstracts. Each abstract was labelled according to whether it satisfied the sample criteria enumerated in 3.1. If a publication did not satisfy our sample restrictions, we stipulated a reason. The user interface was developed with the “Meerkat” Python package available at the link: https://github.com/HazyResearch/meerkat.
B.2. Benchmark Comparison to Standard Machine Learning Methods
We measure the performance of a variety of standard machine learning algorithms for classifying publications according to whether they meet our sample restrictions. Models are trained in the training and validation samples and tested in the testing sample.
We compute two embeddings for each abstract in our hand-labeled sample. First, we compute Term Frequency, Inverse Document Frequency (TF-IDF) embeddings based on the corpus of abstracts in the hand-labeled sample. Second, we extract embeddings associated with the abstract of each publication in the hand-labeled sample with the SentenceTransformers language model (Reimers and Gurevych,, 2019). The SentenceTransformers language model is a standard source for embeddings, and is a modification of the BERT language model (Devlin et al.,, 2019).
We consider six classes of prediction algorithms: logistic ridge regression, logistic lasso regression, Support Vector Machines, Random Forest Regression, Boosted Regression, and Convolutional neural networks. Each model is obtained from the “sklearn” Python package. The hyper-parameters of each model are chosen with 10-fold cross-validation implemented in the training and validation samples. A balanced loss is used to measure the performance of each hyper-parameter.
Figure 2 displays receiver operating characteristics for the each class of models estimated with both types of embedding. Error rates are estimated in the testing data. The performance of the receiver operating characteristic for the “ensemble” model whose performance is displayed in Figure 3 is given in purple. The fine-tuned large language model developed in this paper substantially out-paces the performance of standard machine learning methods.
B.3. Prompt Design and Error Analysis
We extract weak labels for 64,000 randomly selected abstracts with two proprietary large language models: OpenAI’s GPT-3.5 and GPT-4 (Nori et al.,, 2023; Bubeck et al.,, 2023). A priori, we expect GPT-4 to output labels of slightly lower quality than hand-labeled output and GPT-3.5 to produce slightly noisier labels than GPT-4. The extraction of GPT-3.5 labels is substantially faster and cheaper.
To produce these labels at scale, each of the two models must be appropriately prompted. That is, each pre-trained model must be provided with a block of text as an input and asked to return the appropriate completion of this input. A priori, it is not clear what prompt structure will work well. The relative infancy of this area of research renders it difficult to identify a set of “best practices.” The proprietary nature of these models makes them difficult to study. See, for details, a guide to prompt engineering from OpenAI here: https://platform.openai.com/docs/guides/prompt-engineering. All of the prompts that we consider are displayed as figures in Appendix D.
We identify three general prompt formats, which differ both in the amount of detail provided about our classification task and in the structure of the requested model completion. The simplest prompt provides a version of our sample definition, 3.1, and the text of an abstract, and asks the model to return ‘TRUE’ of the abstract satisfies these criteria. Otherwise, it returns false. The text of this prompt is displayed in Figure D.8. We refer to this prompt as Prompt 1.0. A second, more complicated prompt provides the same definition, with a set of examples of publication characteristics that do and do not satisfy the definition. The text of this prompt is displayed in Figure D.14. We refer to this prompt as Prompt 2.0. A third, more complex prompt provides the same definitions and examples, but asks the model to return an explanation of why the record does or does not satisfy these criteria. The text of this prompt is displayed in Figure D.17. We refer to this prompt as Prompt 3.0. We devise initial language for each prompt iteratively, using small numbers of records from our hand-labelled validation data.
We test each of these three prompts in our 1000-record hand-labelled validation dataset, using both GPT-3.5 and GPT-4. We conduct a detailed error analysis, reported in LABEL:tab:Error-Analysis. For each instance in which a model returns a label (TRUE/FALSE) that differs from that in the hand-labelled dataset, we inspect the record. We categorize errors into types and subtypes. To make this concrete, suppose that GPT-3.5, given Prompt 1.0 and an abstract that reports the results of a meta-analysis, erroneously classifies the record as clinical trial, per our definition. In LABEL:tab:Error-Analysis, we flag this error under Prompt 1 for GPT-3.5, as an error of type “meta-analysis.” Within meta-analysis, we assign the error to a sub-type. If the abstract explicitly includes terms such as “meta-analysis,” “literature search,” or “literature review,” we categorize this as an explicit error. If the abstract references that the publication is summarizing existing studies, or searching a database for records and collecting their findings, we categorize this as an implicit error. We devise error types based on our inclusion/exclusion criteria, in 3.1, and select sub-types based on common categories of errors.
This exercise is instructive on three margins. First, it highlights the highest performing prompts. Second, it indicates particular types of errors in categorization—which suggest opportunities for more precise language in a prompt. Third, it draws attention to differences in the performance of the two models. Observe that, in the “Other” error type, we include a subtype called “overly literal interpretation of inclusion criteria.” We primary record errors of this type for Prompt 3, which asked the model to return a completion that described why, or why not, a record was classified as being in our sample. Here, GPT-3.5 makes 30 such errors—classifying records as FALSE by recapitulating the sample definition provided. GPT-4 makes two errors. Insight that emerges from this exercise, then, is that different prompt structures may be preferable depending on the model used.
We revise each class of prompt based on these findings. For Prompt 1.0, the simplest true/false prompt, we consider three variants. We refer to these revised prompts as Prompts 1.1, 1.2, and 1.3. The text of these prompts is displayed in Figures D.9, D.10, and D.12. For Prompts 2 and 3, we consider one variant each. We refer to these revised prompts as Prompts 2.1 and 3.1. The text of these prompts is displayed in Figures D.15 and D.18. These changes reflect the differences in performance catalogued in LABEL:tab:Error-Analysis.
Table B.1 reports estimates of the true positive rate and false positive rate for each prompt in both models computed in the validation sample. Observe—for example, with Prompts 1.1, 1.2, and 1.3 queried to GPT-4—that even small changes in the text of a prompt yield substantial differences in performance. We attempt a second round of prompt iteration (Prompt 1.3). We observe that performance deteriorates with even small modification. Thus, we select the two highest performing prompts—one for each proprietary model. We use Prompt 2.0 to extract weak labels using GPT-3.5, and Prompt 1.2 to extract weak labels using GPT-4. These prompts are used to extract 64,000 labels—TRUE/FALSE flags—using GPT-3.5 and GPT-4. We refer to these as our set of “noisy labels.”
Table B.1. Prompt Performance in Validation Data
Panel A: GPT-3.5
Prompt Type Prompt Sub-Type False Positive Rate True Positive Rate
1 0 0.202 0.971
1 0.065 0.949
2 0.049 0.934
3 0.056 0.912
2 0 0.081 0.964
1 0.072 0.964
3 0 0.167 0.971
1 0.068 0.956
Panel B: GPT-4
Prompt Type Prompt Sub-Type False Positive Rate True Positive Rate
1 0 0.247 0.788
1 0.172 0.898
2 0.037 0.584
3 0.037 0.489
2 0 0.162 0.876
1 0.176 0.905
3 0 0.248 0.722
1 0.171 0.883
Notes: Table B.1 records the performance of each of eight prompt variants in a sample of 1,000 validation dataset records. Panel A reports performance associated with the proprietary model GPT-3.5. Panel B reports analogous statistics for model GPT-4. Prompt 1 asks the model to return TRUE or FALSE. Prompt 2 asks the model to return TRUE or the name of a specific, excluded category. Prompt 3 asks the model to return TRUE or an explanation of why the record should be excluded. Prompt sub-types correspond to various iterations. Sub-type 0 is the initial version of the prompt. Sub-types 1-3, where applicable, are subsequent iterations. Appendix D records the text of each prompt. We use Prompt 2, Sub-Type 0, to extract weak labels using GPT-3.5, and Prompt 2, Sub-Type 1 to extract weak labels using GPT-4.
Table B.2. Baseline Prompt Error Analysis
Model GPT-3 GPT-4
Prompt Prompt
Error Type 1.0 2.0 3.0 1.0 2.0 3.0
No Drug
vitamin/supplement 12 12 9 15 4 15
surgical/medical procedure/diagnostic 40 27 29 35 21 29
abstract not specific about name of medicine 3 4 4 6 5 6
food/beverages 8 3 6 13 1 12
surgical/medical device 8 4 5 7 2 6
surgical/medical imaging 1 1 1 2 0 2
behavioral/physical therapy/exercise 40 22 36 25 8 13
misses the mention of a drug 25 10 5 2 1 0
surgical/medical material (e.g., resin/dental filling) 1 0 1 3 1 1
non-medical pollution/chemical/drug 1 1 1 1 2 1
other 1 0 1 2 1 2
Meta-Analysis
explicit mention of meta-analysis/literature search/literature review 13 5 13 8 0 6
mention of database searched, explicit mention of meta-analysis 3 3 3 3 0 2
summary existing studies, without reference to data search / meta-analysis 6 7 4 3 4 3
Retrospective
re-analysis of previously collected data without explicitly mentioning “retrospective” 8 2 6 6 5 4
explicitly described as “retrospective” or “retrospective analysis” 6 1 6 0 0 0
miscategorizes a study conducted in the past as “retrospective” 0 1 0 0 0 0
other
Observational
no active intervention described; does not explicitly say “observational” 10 9 11 6 0 4
no active intervention described; explicitly says “observational” 14 13 13 13 2 11
misses reference to active intervention 1 1 1 0 0 0
Protocol
no results reported about current study; explicitly says “protocol” 10 9 11 6 0 4
no results reported about current study; does not explicitly say “protocol” 14 13 13 13 2 11
study is based on a simulation, not real world data 1 1 1 0 0 0
No Human Subjects
only references to laboratory experiments, cells drawn from humans 0 1 0 0 0 0
Animal
study conducted on animals 5 10 6 2 3 2
Other
overly literal interpretation of inclusion criteria 0 2 30 0 0 2
studying outcomes/objects unrelated to a randomized trial 1 1 1 1 1 1
Notes: LABEL:tab:Error-Analysis categorizes errors documented in the first round of prompt iteration, using Prompts of sub-type 0. See Table B.1 for details on model performance. For each model error—an instance in which a model returned a classification that deviated from hand-labelled data—we examined the associated record. We categorized errors by Type and Sub-Type. Types are drawn from the exclusion restrictions in 3.1. Sub-types describe consistent characteristics of publications that resulted in such an error. We report counts of errors, by model and prompt, of each Type and Sub-type.
B.4. Fine-tuning
This appendix gives technical details related to fine-tuning open-source large language models with noisy labels extracted from GPT-3.5 and GPT-4. We fine-tune pre-trained BERT models from two different architecture classes: (1) BigBird and (2) BioMedBERT (Gu et al.,, 2021; Zaheer et al.,, 2021). Both models us medium-scale Transformer architectures, i.e., between 100 and 300 million parameters, pre-trained with the masked-language modeling (MLM) objective (Vaswani et al.,, 2023; Devlin et al.,, 2019). The models differ in their architectural details and pre-training corpora.
BigBird uses a sparse attention mechanism to reduce the computational cost of processing long text sequences. The computational complexity of regular attention is quadratic in sequence length, while BigBird reduces this to linear complexity. This enables efficient handling of long documents up to 4,096 tokens. In contrast, standard BERT can only process 512 tokens. BigBird is useful because abstracts of clinical trials regularly exceed 512 tokens. Prior to pre-training, the model was warm-started from the RoBERTa checkpoint (Liu et al.,, 2019) and then pre-trained on the standard BERT corpus consisting of the Books (Zhu et al.,, 2015), CC-News (Guu et al.,, 2020), Stories (Trinh and Le,, 2019), and Wikipedia (Wikimedia Foundation,, 2023) datasets. We evaluate two model sizes: BigBird Base (125 million parameters) and BigBird Large (355 million parameters).
BioMedBert uses a domain-specific pretraining corpus sourced from articles on PubMed Gu et al., (2021). It uses the same model architecture as RoBERTa, but a different, domain-specific token vocabulary optimized for PubMed articles. Like BigBird, we evaluate two model sizes: BioMedBert Base (125 million parameters) and BioMedBert Large (355 million parameters).
We fine-tune the pre-trained architectures to classify abstracts according to the inclusion and exclusion criteria enumerated in 3.1. We replace the language modeling classification head with a binary classification head consisting of a single linear layer and softmax activation. We fine-tune the full model (i.e. not just the classification head) with cross-entropy loss. We use the Adam optimizer with learning rate $1\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$ (determined via hyperparameter sweep). We use a maximum sequence length of $4{,}096$ tokens. If an abstract does not fit into the maximum sequence length provided by a model, the abstract is truncated. All models were trained on a single $V100$ GPU with 16GB of HBM.
B.5. Performance Assessment
In this appendix, we assess the performance of various fine-tuned language models for classifying whether a publication satisfies the restrictions enumerated in 3.1. Figure B.2 displays estimates of the receiver operating characteristic of several fine-tuned large language models. Here, we consider only the performance of the RoBERTa base model. The top and bottom panels are trained on noisy labels extracted from GPT-3.5 and GPT-4, respectively. We vary the quantity of noisy labels used to train the language model. Additionally, in the bottom panel, we display the performance of a fine-tuned model trained with the 1000 hand-labelled observations in the training dataset. The performance of each language model is measured in the testing dataset.
Figure B.2. Receiver Operating Characteristic by Data Size and Origin
Notes: Figure B.2 displays estimates of the the receiver operating characteristic of several fine-tuned language models. The models differ in terms of the source of their training data. Solid lines display the performance of models trained on noisy labels extracted from either GPT-3.5 or GPT-4. The dotted line displays the performance of a model trained on the hand-labeled data in the training dataset. Estimates of the true positive rate and false positive rate of GPT-3.5 and GPT-4 are indicated with black dots.
The fine-tuned models trained with noisy labels extracted from GPT-4 significantly outperform the models trained with noisy labels extracted from GPT-3.5 or with the hand-labels. There appears to be a threshold, where performance dramatically improves, at around 8,000 training labels. The models displayed in Figure 3 were each obtained by using all 64,000 training labels.
B.6. Final Model: Error Analysis
The Conservative model incorrectly labels 27 papers. We inspect each of these errors. We review each abstract and aim to understand what might have generated an error. In roughly half of cases–13 of 27—there is a clear error. These clear errors include publications that explicitly report the results of observational studies (one error), that re-analyze existing data (eight errors), and that report literature reviews (one error). This set also includes clinical trials that satisfy all criteria in 3.1, except the treatment being studied is not a drug (three errors). In 14 cases, however, errors are associated with records that are difficult for human labellers to categorize, either because the content of the publication does not fit neatly into the inclusion and exclusion criteria implied by 3.1 or because the publication is written in a way that makes it difficult to determine details of the study. PubMed record 27880726 is illustrative. This publication studies the effect of chrysophanic acid (CA) on benign prostatic hyperplasia. The abstract is unclear about whether this treatment is studied in human subjects or in an animal model–though the publication title and full-text make clear that this was conducted in a sample of rats. Based on the abstract alone, however, nothing suggests that this is an animal study. It is excluded from both the conservative and moderate model-generated samples, but a human-labeller flagged it as satisfying our criteria.
Appendix C Additional Figures and Further Analyses
This appendix collects additional figures and analyses. Figure C.3 displays series analogous to the series displayed in Figure 4, constructed with the Moderate and Liberal sample restrictions. The results are very similar.
One concern with the comparison made in Figure 4 is that researchers may cite a growing number of publications over time. That is, more papers may cite clinical trials over time because citations are becoming, in some sense, “cheaper.” The relevant clinical literature, then, may not be any larger. To alleviate such concerns, Figure C.4 reports counts, where we scale each citation assigned to a paper by the total number of citations in the originating paper. That is, a paper that receives two citations, each from a paper that cites two papers, will have a weighted citation count of 1.
Figure C.5 displays series analogous to Panel A of Figure 5 with the Moderate and Liberal sample restrictions. Figures C.6 and C.7 display heat maps analogous to Panel B of Figure 5, with the Moderate and Liberal sample restrictions and both 3-Year and 5-Year citation counts. In each case, the results are very similar across specifications.
Figure C.3. Growth in Clinical Research, by Sample Stringency
Notes: Figure C.3 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” Each series is reported in terms of the percent change relative to its initial value. To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
Figure C.4. Growth in Cost-of-Citation Weighted Clinical Research
Notes: Figure C.4 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” Here, we report the number of scientific papers using a citation-weighted metric. In particular, we scale each citation assigned to a paper by the total number of citations in the originating paper. Each series is reported in terms of the percent change relative to its initial value. To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
Figure C.5. Public Funding and Low-Quality Research, by Sample Stringency
Notes: Figure C.5 illustrates the stability of the heterogeneity in the composition and quality of published clinical trials between 2010 and 2019 in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” See Figure 3 for the true positive and false positive rates of these models. The figure displays three time series: the proportion of clinical trials that are publicly funded, the proportion of clinical trials that are ever cited by a leading journal, and the proportion of publicly funded clinical trials that are ever cited by a leading journal.
Figure C.6. Quality over Time, 3-Year Cites, by Sample Stringency
Panel A: Citations From Publications in Leading Journals
Panel B: Citations From All Publications
Notes: Figure C.6 displays heat maps measuring the distribution of three-year citations received by clinical trials in each calendar year from leading journals and from all publications, respectively. The $y$-axis has been stretched to elongate the right-tail of the citation distribution. Colors are displayed in a log scale. We consider the sample of clinical trials constructed with the conservative, moderate, and liberal model.
Figure C.7. Quality over Time, 5-Year Cites, by Sample Stringency
Panel A: Citations From Publications in Leading Journals
Panel B: Citations From All Publications
Notes: Figure C.7 displays heat maps measuring the distribution of five-year citations received by clinical trials in each calendar year from leading journals and from all publications, respectively. The $y$-axis has been stretched to elongate the right-tail of the citation distribution. Colors are displayed in a log scale. We consider the sample of clinical trials constructed with the conservative, moderate, and liberal model.
Appendix D Prompt Repository
This appendix serves as a repository for figures displaying each of the prompts considered in Section B.3. Throughout, the dummy text “{abstract}” is placed in the location where the abstract of a publication would be input. The baseline “True/False” Prompt 1.0 is displayed in Figure D.8. Refinements to this prompt are displayed in Figures D.9, D.10, and D.12. The prompt that asks the model to categorize excluded publications, i.e., Prompt 2.0, is displayed in Figure D.14. A refinement to this prompt is given in Figure D.15. The prompt that asks the model to explain why it has excluded a publication, i.e., Prompt 3.0, is displayed in Figure D.17. In turn, a refinement to this prompt is given in Figure D.18.
Figure D.8. Prompt 1.0: True/False, Base
Figure D.9. Prompt 1.1: True/False, Initial Examples
Figure D.10. Prompt 1.2: True/False, Extended Examples, Short
Figure D.11. Prompt 1.2: True/False, Extended Examples, Short (Cont.)
Figure D.12. Prompt 1.3: True/False, Extended Examples, Long
Figure D.13. Prompt 1.3: True/False, Extended Examples, Long (Cont.)
Figure D.14. Prompt 2.0: Categorize Exclusion Restriction
Figure D.15. Prompt 2.1: Categorize Exclusion Restriction, Examples
Figure D.16. Prompt 2.1: Categorize Exclusion Restriction, Examples (Cont.)
Figure D.17. Prompt 3.0: Provide Reason for Exclusion
Figure D.18. Prompt 3.1: Provide Reason for Exclusion, Examples
Figure D.19. Prompt 3.1: Provide Reason for Exclusion, Examples (Cont.)

### v2

@@
 JEL: C81, O32
-Across fields, the quantity of research has increased substantially, without an attendant increase in output.
-We argue that, in medicine, this indicator of declining productivity reflects a compositional shift toward low-capital, low-productivity research.
-Using a fine-tuned, open-source large language model, we construct a novel census of capital-intensive, high-productivity medical investment—clinical trials.
-Since 2010, the annual quantity of clinical trials has been constant.
-By contrast, the quantity of other forms of clinical research has increased substantially.
-Within clinical trials, there is substantial heterogeneity in productivity.
-The distribution of this heterogeneity is stable over time.
-Keywords: Medical Innovation, Productivity, Large Language Models
+A recent literature contends that the returns to research in medicine are declining.
+This conclusion is premised on two empirical findings: the quantity of medical research is rising, while improvements in health outcomes are stagnant.
+We argue that the first finding is an artifact of conceptual and empirical problems in the measurement of research.
+We propose a method for annotating unstructured text based on fine-tuning a large language model.
+We apply this method to construct a census of published clinical trials.
+We argue that clinical trials function as a natural proxy for productivity-relevant research in medicine.
+In our data, the quantity, quality, and composition of clinical trials have been stable since 2010.
+We find that measured increases in the total quantity of medical research are driven by shifts in the composition of scientific publications that do not correspond to real increases in the quantity of productivity-relevant research.
+We document large increases in the growth of non-trial publications, driven by publications from China and publications that review existing literature.
+This increase in quantity is coincident with a decline in various measures of average publication quality.
+Taken together, we find no support for declining medical research productivity over this period.
+Keywords: Medical Innovation, Productivity, Clinical Trials, Large Language Models, Generative AI
 JEL: C81, O32
 1. Introduction
-Technological progress in medicine is widely credited with enabling large improvements in human health (Easterlin,, 1999; Mokyr et al.,, 2002; Acemoglu and Johnson,, 2007).
-Breakthroughs at the turn of the century—the completion of the Human Genome Project, the advent of DNA sequencing, the adoption of combinatorial chemistry—promised transformation of diagnosis and treatment or, at least, an expansion of the set of available scientific opportunities.
-[^1] See Scannell et al., (2012) for details on these factors, in a broader discussion of declining productivity in the pharmaceutical industry.
-Combinatorial chemistry, in the 1980s and 1990s, increased the number of molecules (potential drug candidates) that a chemist could synthesize “perhaps 800-fold.” Similarly, since the first genome sequence was determined in the 1970s, DNA sequencing has become “over a billion times faster.” Yet, an emerging literature finds that biomedical research productivity is in a state of decline; see Goldin et al., (2024) for a review.
-This literature is structured around an incongruity identified by Bloom et al., (2020).
-The quantity of medical research—proxied by estimates of the number of clinical trials and scientific publications—has increased substantially, without an accompanying improvement in mortality.
-Many explanations center on characteristics of the “idea production function.” Perhaps low-hanging fruit have been depleted (the fishing out hypothesis, described in Jones,, 1995), [^2] See, also, Cowen, (2011) and Gordon, (2017).
-or else the educational costs to a young innovator of reaching the frontier of knowledge are increasingly high (the burden of knowledge hypothesis, due to Jones,, 2009).
-Though the specific parameter implicated differs, the mechanisms underlying these hypotheses suggest a structural change in research productivity.
-This paper tests an alternative, compositional hypothesis.
-We document that increases in the quantity of medical research have been driven by increases in the quantity of unproductive, low-capital research with limited capacity to effect quantitatively significant changes to mortality.
-As a consequence, increased quantities of medical research need not indicate that the productivity of high-capital research investment is declining.
-We construct a novel census of capital-intensive, high-productivity medical investment—clinical trials.
-We contrast this set with its complement in clinical research, proxied by publications that cite clinical trials.
-This comparison is natural and reflects the standard hierarchy of medical evidence (Jones and Podolsky,, 2015).
-Decision-making in medicine—by regulators, firms, insurers, physicians, and patients—hinges on the causal, mortality-relevant data produced by costly, time-intensive randomized experiments (Chavez-MacGregor and Giordano,, 2016; Alsan et al.,, 2024).
-Although research engaging with trial evidence undoubtedly produces valuable information, it does so at comparatively low cost and with lesser impact (Kones et al.,, 2014).
-Since 2010, we find that the quantity, quality, and composition of published clinical trials has been stable.
-By contrast, the quantity of other clinical research has doubled.
-No existing dataset contains sufficient information to assess the composition of clinical research.
-Administrative databases—such as ClinicalTrials.gov—aim to provide an index of trials, but suffer from widespread noncompliance with reporting requirements (DeVito et al.,, 2020).
-Proprietary databases take as an input the contents of ClinicalTrials.gov and other registries and, thus, capture changes in reporting patterns over time (e.g., Cortellis,, 2024).
-Best practices for identifying trials in publication data do so imprecisely, with both high true and false positive rates (Thomas et al.,, 2021).
-We construct a new census of clinical trials.
-We collect all records of scientific publications indexed in the National Library of Medicine’s PubMed / MEDLINE database, from 2010 forward, that disclose the results of a clinical trial studying a medicine in human subjects.
-Our definition is restrictive and intended to capture only those research inputs that are most relevant for human mortality.
-Our goal is to identify clinical trials in our sample of interest—from the set of roughly 35 million PubMed records—with an accuracy and precision close to that of a human labeller.
-Standard machine learning methods perform poorly and hand-labelling at scale is infeasible.
-Instead, we train a bespoke large language model, optimized for our task.
+Two stylized facts support the argument that the returns to medical research are diminishing.
+[^1] Goldin et al., (2024) survey evidence related to declining productivity across the economy, including evidence of diminishing returns to research investment in sectors beyond medicine.
+The quantity of resources directed toward medical research, measured in dollars and scientific publications, appears to be increasing (Scannell et al.,, 2012; DiMasi et al.,, 2016; Bloom et al.,, 2020; Park et al.,, 2023).
+[^2] See Park et al., (2023), Bloom et al., (2020), and Scannell et al., (2012) for case studies examining the quantity of investment in science, medicine, and drug development, respectively.
+Bloom et al., (2020) collect case studies from other research-intensive fields.
+See, also, Chu and Evans, (2021), who document large increases in the quantity of papers in certain scientific fields and examine associated patterns in citation networks.
+Meanwhile, improvements in human health (e.g., increases in life expectancy) continue at the same steady pace (Mehta et al.,, 2020; Ho and Hendi,, 2018).
+If more resources are necessary to yield the same rates of technological progress, simple models of “idea production” imply that research productivity—the efficiency with which research is converted into new and useful knowledge—must be falling (Jones,, 1995; Bloom et al.,, 2020).
+[^3] In some models, these patterns can be reconciled without diminishing returns to research.
+For instance, Young, (1998) investigates the idea that investment in new varieties of existing products can yield the same trends.
+These facts, and their forecast of technological stagnation in a sector closely linked to health and welfare (Cutler and McClellan,, 2001; Mokyr et al.,, 2002; Acemoglu and Johnson,, 2007), have influenced prominent policy debates.
+[^4] These facts were most prominently documented in Bloom et al., (2020).
+Concerns about technological stagnation in modern policy discourse predate Bloom et al., (2020) and respond in part to Cowen, (2011) and Gordon, (2017).
+As we discuss below, the prospect of diminishing returns to scientific research has long been a subject of concern.
+Proposals that would restructure public subsidies administered by the National Institutes for Health (NIH) (e.g., Executive Office of the President,, 2019; Moss et al.,, 2020; The Economist,, 2022), relax the regulatory standards for approval of new medicines by the U.S.
+Food and Drug Administration (FDA) (e.g., Roy,, 2021; Chertman,, 2023; Graboyes and Reimschisel,, 2017), and prevent the Centers for Medicare and Medicaid Services (CMS) from negotiating drug prices (e.g., York,, 2023; Long,, 2023; Inflation Reduction Act,, 2022) are premised, in no small part, on the concern that science and medicine are in a state of decline.
+[^5] This concern has been explored extensively in the popular press.
+Collison and Nielsen, (2018), Broad, (2023), Thompson, (2021), and Piper, (2023), among others, interpret these facts in the popular press as a “warning sign” for the future of technological progress (Piper,, 2023).
+In this paper, we find that one of these stylized facts—the increase in the quantity of resources directed toward medical research—is an artifact of conceptual and empirical problems in the measurement of medical research.
+An ideal, and infeasible, measure of the quantity of research captures the amount of “effort” directed toward increasing the stock of useful (medical) knowledge.
+Informative proxies for this measure have two characteristics.
+First, as researchers differ in skill and research designs differ in quality, the proxy must account for heterogeneity in the likelihood that research projects yield “productivity-relevant” information.
+Second, units of the proxy must be comparable over time.
+If the quantity of research embedded in, say, a patent or paper decreases with time, increases in the quantity of patents or papers provide very little insight into research productivity.
+We find that existing measures of medical research fall short on both accounts.
+The primary contribution of this paper is the construction of a measure of medical research—a census of published clinical trials—that addresses both issues and fills this gap.
+We show that standard data construction tools classify research imprecisely, and that this imprecision yields inaccurate conclusions about trends in productivity.
+Thus, we introduce a method for annotating unstructured text at scale.
+When applied to a corpus of scientific text, this method allows us to identify and classify heterogeneous forms of research, relaxing a long-standing data constraint in the economics of innovation.
+We fine-tune an open-source large language model and devise a workflow that allows us to identify clinical trial records from the unstructured text of scientific publications with an accuracy comparable to that of a human labeler.
+In our data, the quantity, quality, and composition of clinical trials have been stable since 2010.
+We find evidence that measured increases in the quantity of scientific publications capture shifts in scientific incentives that are not indicative of real increases in the quantity of (productivity-relevant) research.
+Taken together, we find no support for claims of declining medical research productivity over this period.
+We begin by arguing that clinical trials are a natural proxy for the research object of interest in this setting.
+Trials are discrete, regulated experiments that generate useful statistical information about new and existing technologies (Chavez-MacGregor and Giordano,, 2016; Alsan et al.,, 2024).
+These experiments are designed in consultation with regulators after sufficiently robust preliminary evidence has been produced but before health technologies can be marketed.
+Trials top the hierarchy of medical evidence employed by firms, regulators, physicians, and patients (Jones and Podolsky,, 2015).
+Since the early 2000s, clinical trials have been reported in a consistent, standardized format across scientific publications.
+[^6] In the 1990s and 2000s, medical experts created a checklist that provides authors of papers reporting the results of clinical trials with a standardized structure and list of elements for publication.
+The most recent Consolidated Standards of Reporting Trials (CONSORT) Statement was published in 2010 and is now the reporting standard for most scientific journals that publish the results of trials.
+See Laine et al., (2007); Schulz et al., (2010) for details.
+To our knowledge, this type of standardization is unique to this form of science.
+No existing dataset allows us to track quantities of clinical trials over time.
+Although dollar-denominated expenditures are an obvious metric, available records of investment in trials are self-reported, in some form, by pharmaceutical firms and are essentially impossible to validate independently.
+Administrative databases—such as ClinicalTrials.gov—aim to provide an index of trials, but suffer from widespread noncompliance with reporting requirements (e.g., DeVito et al.,, 2020).
+Proprietary databases (e.g., Cortellis,, 2024) take as an input the contents of publicly-maintained registries and, thus, are biased by these same noncompliance patterns.
+We construct a new census of clinical trials indexed in scientific publications.
+We collect all records in the National Library of Medicine’s PubMed / MEDLINE database, from 2010 forward, that disclose the results of a clinical trial that studied the effects of a medicine in human subjects.
+This definition is intentionally restrictive and aims to include only those records that clearly report trials.
+Our goal is to identify clinical trials in our sample of interest—from the set of roughly 35 million PubMed records—with an accuracy and precision close to that of a human labeler.
+Manually identifying clinical trial records, at this scale, is infeasible.
+[^7] A small number of papers, including Oostrom, (2024), examine medical investment using small samples of clinical trial records that are constructed by hand.
+Standard methods perform poorly.
+In particular, we show that documented increases in the quantity of clinical trials during this time period, as reported by Bloom et al., (2020), are driven by a categorization error caused by keyword-based searches.
+[^8] Specifically, Bloom et al., (2020) categorize many records as clinical trials that discuss trials (e.g., meta-analyses and literature reviews).
+We find that changes in the scientific publishing landscape, discussed in Section 4, increase the size of this categorization error over time.
+In our period of interest, this generates a spurious increase in the measured quantity of trials.
+Machine learning methods used in the health services literature similarly lack the precision necessary for our purposes (Thomas et al.,, 2021).
+Instead, we leverage recent advances in generative AI, which allow for data extraction from unstructured text at scale.
+We devise a workflow for extracting precisely defined objects from a large corpus of text, and train a bespoke large language model that is specifically optimized for this task.
 We proceed in stages.
@@
 These noisy labels are used as training data to fine-tune a set of open-source models.
-The resulting model matches the performance of the best proprietary model and produces a sample of approximately 150,000 clinical trials.
-We document two sets of facts.
-First, there is quantitatively significant heterogeneity in both the composition of medical research and the composition of clinical trials.
-While clinical trial counts are essentially constant over time, the quantity of papers citing trials has increased by nearly a factor of two.
-This trend reflects a large increase in the quantity of research originating from China, in addition to other countries who, individually, do not account for a large share of the total quantity of medical research.
-Within clinical trials, proxies for quality and informativeness suggest considerable heterogeneity.
-Roughly half of trials in our sample are never cited by a leading medical journal and nearly 15 percent are never cited by another scientific publication.
-Second, these forms of heterogeneity within trials are stable over time.
-We find that—across measures of quality, importance, and geography—the distribution of clinical trial heterogeneity is unchanging.
-Taken together, we interpret these facts as evidence that the productivity of high-capital medical investment is stable over our time period.
-Our choice of setting reflects two substantive considerations.
-First, the determinants of technological progress in medicine are interesting in their own right.
-[^3] Cutler and McClellan, (2001) observe that technological change has accounted for the bulk of increases in health spending over time, but that health has improved alongside these increases.
-In a series of case studies, they argue that the benefits of technological progress exceed their costs.
-See, also, Almond et al., (2010), who estimate that the returns to medical spending are positive on the margin.
-Second, and more centrally, nearly all micro-level evidence on rising research investment, and declining productivity, comes from studies of clinical trials and scientific publications.
+An ensemble of the fine-tuned models matches the performance of the best proprietary model, with false positive and false negative rates below 5 percent, and produces a sample of approximately 150,000 publications that report the results of clinical trials from 2010 to 2022.
+These data allow us to document two sets of facts about the quantity of medical research over time.
+First, we find that the quantity of clinical trials has been effectively constant, year-on-year, across our period of interest.
+We collect a set of ex-ante and ex-post quality measures, derived from scientific publication data.
+On these dimensions, we find considerable heterogeneity.
+Roughly half of clinical trial publications in our sample are never cited by a leading medical journal.
+Nearly 15 percent are never cited by any other scientific publication.
+However, we find that the distribution across clinical trials of measures of quality, importance, and geography is unchanging.
+Taken together, we interpret these patterns as evidence that “productivity-relevant” research in medicine is essentially constant in quantity and stable in composition over our time period.
+This contradicts existing findings, at least for our period of interest, and calls into question the conclusion that the productivity of medical research is falling.
+Second, our data allow us to examine the source of measured increases in the quantity of scientific publications (Bloom et al.,, 2020; Park et al.,, 2023; Chu and Evans,, 2021).
+We contrast clinical trial production with trends in the publication of other forms of medical research.
+While the quantity of clinical trials is constant across our period of interest, the quantity of other medical research—defined, in our preferred version, as the set of papers citing clinical trials—has increased by nearly a factor of two.
+[^9] In Section 4, we inspect the contents of this set and document that nearly all records fall into one of three categories: observational studies, preclinical studies, and meta-analyses/literature reviews.
+In doing so, we confirm that our focus on clinical trials is appropriate.
+These increases in quantity have coincided with shifts in the quality and composition of these other types of research.
+We highlight three patterns.
+First, the large increase in the total number of medical papers in our sample is largely driven by researchers in China.
+Second, this rise in quantity coincides with a decrease (on various measures) in the average quality of publications.
+Third, there is a substantial increase in the quantity of meta-analyses and literature reviews during our period of interest.
+[^10] These patterns are consistent with the geometric increase in the quantity of meta-analyses produced by Chinese researchers, documented in Ioannidis et al., (2013).
+We interpret these patterns as evidence that—for this category of other medical research—the quantity of productivity-relevant research “embedded” in a single document may be declining over time.
+Quantities of publications, then, provide essentially no insight on the question of research productivity.
+[^11] Park et al., (2023) examine the “disruptiveness” of scientific research, using data on the universe of publications across six decades.
+Their findings are similar to ours, in the sense that they document that the absolute counts of highly disruptive papers remain constant over time, while average “disruptiveness” is declining.
+They interpret this as evidence that science is becoming less disruptive over time.
+Our paper differs in two important ways.
+First, we distinguish between types of publications on the basis of the economic value they produce.
+Second, we document that measured increases in the quantity of publications are disproportionately concentrated in types of publication that focus on diffusion, rather than production, of new information.
+Pooled averages, then, may be difficult to interpret.
+Our findings speak directly to ongoing policy discussions and branches of the macroeconomic literature, which consider the hypothesis that the rate of technological progress may be declining.
+Nearly all micro-level evidence on rising levels of research investment comes from studies of scientific publications and, in particular, clinical trials.
 [^12] See Goldin et al., (2024) for a summary.
 Notable exceptions are studies of agriculture and semiconductors.
-Measured increases in the quantity of clinical trials (Bloom et al.,, 2020) and average “declining disruptiveness” of scientific papers (Park et al.,, 2023) are at the center of active debates about whether research in science and medicine has hit a point of diminishing returns (e.g., Dieppe and Kose,, 2020).
-[^5] See discussions in the popular press, e.g.
-Collison and Nielsen, (2018); Broad, (2023); Thompson, (2021); Piper, (2023).
-Concerns about declining productivity have influenced a wave of policy interest in new mechanisms to promote research and development, as well as the revival of “industrial policy” (see e.g., Council of Economic Advisers,, 2021, 2024; Lindsey and Hammond,, 2020; Hausmann,, 2023).
-A precise characterization of compositional changes in this industry, then, is a directly policy-relevant object.
-We propose and validate the existence of quantitatively significant heterogeneity in medical research.
-[^6] Park et al., (2023) examine “disruptiveness” in scientific research, using data on the universe of publications across six decades.
-They document that the absolute counts of highly disruptive papers remain constant, consistent with our finding that high-capital, high-productivity research remains stable.
-Our setting allows for a clear delineation between types of research investments and, importantly, allows us to take a stand on the value of information produced by these distinct types for welfare-relevant outcomes.
-This is the primary contribution of this paper.
-We measure research inputs with a precision that is new to this literature, enabled by the existence of novel technology for data construction at scale (Achiam et al.,, 2023; Bubeck et al.,, 2023).
-Our decomposition suggests that measures of research investment that neglect these differences—and, instead, infer uniform productivity or capital for scientific investments—can yield misleading conclusions about industry-wide productivity.
-This paper contributes to a growing literature on the causes of declining productivity, in general and in medicine.
-Additional explanations include market imperfections and failures—including barriers to innovation (see e.g., Aghion et al.,, 2019), inadequate public sector support for research (Gruber and Johnson,, 2019), and productivity imbalances across sectors (Acemoglu et al.,, 2024).
-Our work is related to a literature—which Syverson, (2017) summarizes as putting forth a “mismeasurement hypothesis”—that views the measured decline in aggregate productivity as illusory (see e.g., Mokyr,, 2014).
-In contrast, our hypothesis is not that productivity is mismeasured, per se, but that failure to account for heterogeneity can generate a misattribution of trends.
+We show that these findings result from conceptual and empirical challenges in the quantification of innovation.
+Thus, our work is most closely related to a series of papers—which Syverson, (2017) summarizes as putting forth a “mismeasurement hypothesis”—that view the measured decline in aggregate productivity as illusory (see e.g., Mokyr,, 2014).
+[^13] Additional explanations include market imperfections and failures—including barriers to innovation (see e.g., Aghion et al.,, 2019), inadequate public sector support for research (Gruber and Johnson,, 2019), and productivity imbalances across sectors (Acemoglu et al.,, 2024).
 More directly, Myers and Pauly, (2019), Scannell et al., (2012), and Cockburn, (2006) investigate the determinants of declining productivity in the pharmaceutical sector, with a focus on large increases in factor prices for clinical trials rather than quantities.
-Finally, we draw on, and contribute to, a long literature on the measurement of innovation.
-Since Pakes, (1986), economists have recognized that patents—the most commonly used measure of research output—differ greatly in economic and technological significance.
-Griliches, (1990) provides a classic discussion of the difficulties associated with using patents as economic indicators, given this heterogeneity.
-[^7] Bryan and Williams, (2021) provide an updated survey and discuss the challenges of devising credible measures of innovative effort.
-Our work builds on these insights to investigate the implications of heterogeneity in inputs for the use of medical research as a productivity indicator.
-2. Productivity, Capital, and Medical Research
-Studies of research productivity have, at their core, some notion of an “idea production function”—a mapping from research inputs (effort) to outputs (ideas).
-The canonical Romer, (1990) specification is given by
+In doing so, we draw on, and contribute to, a long literature on the measurement of invention.
+Pakes and Griliches, (1980) begin a discussion of inventive activity led by U.S.
+firms by remarking: “Perhaps the most serious task facing empirical work in the area of ‘technological change’ and ‘invention and innovation’ is the construction and interpretation of measures (indices) of advances in knowledge.” In a more recent survey, Bryan and Williams, (2021) note that, although much progress has been made over half a century of research, construction of such measures remains an open challenge.
+Decades of work have focused, in particular, on the development of proxies for the outputs of technological innovation.
+An especially prominent approach, due to Hall et al., (2000), uses citation-weighted counts of patents.
+Kogan et al., (2017) and Kelly et al., (2021) provide two recent, novel, alternatives.
+Our focus, by contrast, is on the difficulties associated with measuring the inputs of innovation, which can similarly bias estimates of macroeconomic trends.
+While scientific papers are a natural alternative to patents on the input side, existing methods for the use of patent data are not easily applied to publications (Popp,, 2015; Popp et al.,, 2010).
+[^14] In particular, citations across scientific papers are not analogous to citations across patents.
+Whereas patent citations are legal objects, which are defined in part by patent examiners and delineate the boundaries of a right enforceable by courts, citations across scientific papers reflect norms and networks.
+See Hall et al., (2000).
+The idea that patterns like those that motivate this paper—rising research quantities and stagnant outcomes—signal the “exhaustion of inventive opportunities” (a phrase due to Griliches, (1994)) is, similarly, an old one.
+Early appearances in the literature include Robert K.
+Merton’s 1935 discussion of the “possibility of a slackening in the rate of technologic advance” (Merton,, 1935), Alfred B.
+Stafford’s 1952 paper titled “Is the Rate of Invention Declining?” (Stafford,, 1952) and Jacob Schmookler’s various investigations (Schmookler,, 1952, 1954, 1966) into patterns suggestive of diminishing returns to research.
+Our reconciliation of these patterns hews closely to that of this earlier literature, which viewed the evidence base on productivity decline as inconclusive, given the shortcomings of available measures of innovation (Pakes and Griliches,, 1980; Griliches,, 1990, 1994).
+[^15] Griliches, (1990) summarizes his version of this conclusion more prosaically: “One can always worry that the world is coming to an end.
+Someday it undoubtedly will, but it does not look as if the end is already upon us, at least not yet.” This paper differs on one important dimension.
+We show that advances in data construction technology, if deployed carefully, provide researchers with sufficiently precise measures to draw conclusions.
+A nascent literature investigates the use of large language models in economics (Dell,, 2024; Korinek,, 2023).
+We join a small set of papers, e.g., Bartik et al., (2023) and Dell, (2024), that use generative AI as a tool to construct high-quality economic data.
+Our approach is distinct in this literature, in that we fine-tune an open-source language model using data extracted from a larger, proprietary model.
+In doing so, we show that it is feasible for researchers to achieve the quality of hand-constructed data at a practically interesting scale, with a reproducible and transparent methodology, at a feasible cost.
+Our use of these technologies is distinct from other applications considered in e.g., Korinek, (2023), Manning et al., (2024), Bybee, (2023), Horton, (2023), and Mullainathan and Rambachan, (2024), which focus on the use of generative AI for other aspects of the scientific process, including hypothesis identification, survey participation, idea generation, and scientific writing.
+More generally, we contribute to the large empirical literature that uses text as data (Gentzkow et al.,, 2019).
+We anticipate that our procedure for the construction of new, precise data on scientific research may be applicable to other contexts where unstructured text is available and manual construction of data is infeasible.
+We highlight a set of steps—in validation, prompt design, and fine-tuning of open source models—that yield data of sufficiently high quality at comparatively low cost.
+2. Measuring Research
+Studies of research productivity have, at their core, some notion of an “idea production function”—a mapping from research (e.g., scientists, dollars, research capital) to new, useful information.
+The canonical Romer, (1990) specification, which motivates a recent empirical investigation in Bloom et al., (2020), is given by
 $\frac{\text{d}A_{t}}{\text{d}t}\frac{1}{A_{t}}=\alpha_{t}S_{t}\leavevmode\nobreak\ .$ (2.1)
@@
 [^16] See e.g., Romer, (1990), Jones, (1995), and Aghion and Howitt, (1992), among many others.
-The revival of interest in tools, such as industrial policy, to better allocate public resources to research and development is a reaction, in part, to two stylized facts (Council of Economic Advisers,, 2021, 2024; Lindsey and Hammond,, 2020; Hausmann,, 2023).
-First, the growth rate of research output has been constant (i.e., the left-hand side of (2.1) is constant).
-Second, the quantity of resources devoted to research (i.e., the factor $S_{t}$) has increased substantially.
-The linear specification (2.1) then implies that research productivity $\alpha_{t}$ has declined.
-That is, ideas are getting harder to find.
-For a review of the relevant literature, see Goldin et al., (2024).
-Detailed characterization of these facts is due to Bloom et al., (2020).
-2.1. Heterogeneity
-We propose an alternative reconciliation of these facts.
-Our hypothesis hinges on heterogeneity.
-If measured increases in the quantity of research are accompanied by a compositional shift toward unproductive or low-capital research, then constant growth does not necessarily imply declining productivity for capital-intensive research.
-That is, large increases in the quantity of low-capital-intensity or low-productivity research need not have quantitatively important impacts on growth.
-To fix ideas, let the quantity $\mathcal{I}_{t}$ index the set of “research units” available at time $t$.
-There are two types of research units, high, $\mathsf{H}$, and low, $\mathsf{L}$.
-Each research unit of type $\mathsf{H}$ has a stock of research capital $S^{\mathsf{H}}_{t}$ and a productivity coefficient $\alpha^{\mathsf{H}}_{t}$.
-Similarly, research units of type $\mathsf{L}$ have research capital and productivity $S^{\mathsf{L}}_{t}$ and $\alpha^{\mathsf{L}}_{t}$, respectively.
-Consider the generalized production function
-$\frac{\text{d}A_{t}}{\text{d}t}\frac{1}{A_{t}}=|\mathcal{I}^{\mathsf{H}}_{t}|\text{ }\alpha^{\mathsf{H}}_{t}S^{\mathsf{H}}_{t}+|\mathcal{I}^{\mathsf{L}}_{t}|\text{ }\alpha^{\mathsf{L}}_{t}S^{\mathsf{L}}_{t}\leavevmode\nobreak\ ,$ (2.2)
-where $\mathcal{I}^{\mathsf{H}}_{t}$ and $\mathcal{I}^{\mathsf{L}}_{t}$ index the set of $\mathsf{H}$-type and $\mathsf{L}$-type research units and ||$\cdot$|| ⋅ | denotes the cardinality operator.
-That is, relative to the canonical specification (2.1), we entertain heterogeneity across research units in both productivity and capital-intensity.
-Suppose that the quantity, productivity, and capital of $\mathsf{H}$-type units is constant over time.
-Thus, in each period $t$, the first term in (2.2) is given by $|\mathcal{I}^{\mathsf{H}}|\text{ }\alpha^{\mathsf{H}}S^{\mathsf{H}}$, where we have removed the time subscript.
-Suppose that at time $t=0$, there are no $\mathsf{L}$-type units, and that at time $t=1$, the quantity of $\mathsf{L}$-type units $\mathcal{I}^{\mathsf{L}}_{1}$ becomes non-zero.
-The impact of this change only affects growth through the factors $S_{1}^{\mathsf{L}}$ and $\alpha_{1}^{\mathsf{L}}$.
-If either quantity is negligible, relative to the contribution of the $\mathsf{H}$-type units, then there is a negligible impact on growth.
-This story is closely related to a strand of the economic growth literature that argues that measured increases in the quantity of research reflect increases in the number of product varieties (Dinopoulos and Thompson,, 1998; Peretto,, 1998; Young,, 1998; Howitt,, 1999).
-By contrast, we examine the idea that shifts in the relative shares of established varieties can rationalize the same trends.
-2.2. Medical Research
-The objective of this paper is to test the compositional hypothesis described in Section 2.1.
-We do so in the context of medical research.
-We argue that in this setting we are able to decompose research units into coarse types.
-Moreover, for each type, institutional context allows us to infer absolute and relative productivity and capital-intensity.
-Measures of quantity can be recovered from existing data.
-An additional advantage of this setting is that medical research has an intuitive notion of output: improvement in mortality.
-To keep our focus on the composition of inputs—the right-hand side of (2.2)—we note that it is a well-established empirical regularity that life expectancy has risen at a steady rate since 1840 (see e.g., Oeppen and Vaupel,, 2002; Bloom et al.,, 2020).
-We proceed assuming that the left-hand side of (2.2) is constant, as is standard in this literature.
-[^9] We make the assumption that life expectancy is improving at a steady rate, given the stability of these trends for more than 150 years.
-It is of course unlikely that medical innovations developed in the early 2010s are clearly discernible in recent mortality data, especially given the impact of the COVID-19 pandemic.
-Medical research takes many forms.
-We focus on two types: clinical trials, and other forms of clinical research that aim to inform the provision of medical care.
-Clinical trials are expensive, time-consuming, highly-regulated experiments, which have as their goal the production of causal and mortality-relevant evidence (Alsan et al.,, 2024; Wouters et al.,, 2020).
-The contours of the second set, other forms of clinical research, are less precise.
-This set includes observational studies, meta-analyses and systematic reviews, cost-effectiveness studies, etc.
-(Food and Drug Administration,, 2024).
-In contrast to clinical trials, these other study types are less capital-intensive and less productive, where productivity is defined as their efficiency in turning research inputs into changes in mortality (Jones and Podolsky,, 2015; Chavez-MacGregor and Giordano,, 2016; Kones et al.,, 2014).
-[^10] Of course, one can identify specific studies that deviate from our categorization.
-See Kones et al., (2014) for discussion of a setting in which, the authors argue, observational data are as informative as randomized controlled trial evidence.
-However, clinical trials remain the “gold standard” for medical research (Jones and Podolsky,, 2015).
-Non-randomized evidence is consistently excluded from prescribing guidelines, on the grounds that it provides evidence of lesser quality.
-Chavez-MacGregor and Giordano, (2016) and Kones et al., (2014) discuss hierarchies of medical evidence in more detail.
-3. Identifying Clinical Trials in Publication Data
-We construct a census of clinical trials.
-No existing dataset, used off-the-shelf, is sufficient.
-Although dollar-denominated expenditures are an obvious metric, available records of investment in clinical trials are self-reported, in some form, by pharmaceutical firms and are essentially impossible to validate independently.
-[^11] There are two frequently used sources of data on dollar-denominated expenditures.
-DiMasi et al., (2016) provide one set of estimates of drug research and development costs, based on confidential data obtained from 10 firms.
-Myers and Pauly, (2019) and Scannell et al., (2012) use research and development expenditure figures published by the Pharmaceutical Research and Manufacturers of America, a trade organization representing the industry.
-Administrative databases—such as ClinicalTrials.gov—aim to provide an index of trials, but suffer from widespread non-compliance with reporting requirements, with substantial variation in enforcement over time (DeVito et al.,, 2020).
-Proprietary databases pose two challenges: first, it is, again, difficult to validate their contents and, second, leading databases take as inputs data in public registries, thus adopting the same shifts in reporting over time (see e.g., the list of registry inputs for one frequently-used database, Cortellis,, 2024).
-For each of these options, there is no symmetric way to identify the quantity of other clinical research.
-3.1. Publication Data
+Although each element of this framework is simple, constructing their empirical analogues is challenging.
+Decades of scholarship—summarized in Griliches, (1990, 1994) and, more recently, Bryan and Williams, (2021)—have grappled with alternative ways of measuring outputs, $A_{t}$.
+[^17] Pakes and Griliches, (1980) consider this challenge and employ slightly different notation than we use here.
+They focus on the quantification of an output variable $K$, which is the “level of economically valuable technological knowledge.” In this paper, our focus is on the comparable challenges associated with quantifying the resources that go into their production, $S_{t}$.
+An ideal measure of resources $S_{t}$ should capture the total quantity of economic resources devoted to increasing the stock of useful knowledge.
+Not all economic activities related to scientific research are equally relevant.
+Many forms of scientific output have as their goal the synthesis or diffusion of existing information.
+In medicine, for example, literature reviews and meta-analyses allow physicians to rapidly review the empirical evidence supporting particular treatment options.
+[^18] For one discussion of the appropriate interpretation of meta-analyses in clinical practice, see Thompson and Pocock, (1991).
+Medical guidelines, produced by entities like the American Heart Association (e.g., Smith Jr.
+et al.,, 2001), consolidate these findings into consensus documents that may be viewed as defining the “standard of care” (Moffett and Moore,, 2011).
+These types of literature syntheses are, anecdotally, of increasing importance, as health technologies previously available only in a small number of countries are made available to new patient populations (Ioannidis et al.,, 2013; Cockburn et al.,, 2016).
+It is important to be clear about the distinction we are making between these forms of research.
+This paper, and productivity calculations more generally, are interested in those forms of research that generate new information.
+These are the elements of the quantity $S_{t}$.
+From a welfare perspective, there are many other types of investment that are important, including these types of literature reviews that aid in the diffusion of information.
+These other investments may enable the “resources related by productivity growth .
+.
+.
+[to] find adequate employment,” thus increasing aggregate welfare, but are not themselves contributing to productivity growth (Griliches,, 1994).
+Even if, conceptually, it is clear that certain economic activities are more relevant to the quantity $S_{t}$ than others, available indicators of research—such as scientific publications and patents—may make it challenging to accommodate this heterogeneity empirically.
+This task is somewhat simpler in industries often studied by economists (see, e.g., the discussion of manufacturing firms in Syverson,, 2011), where dollar-denominated expenditures and price indices are available.
+In many settings central to economic growth, however, research occurs in university-based laboratories that do not maintain detailed financial ledgers, [^19] The UMETRICS data are one exception.
+See Babina et al., (2023) for an example of a study employing financial data from a university research setting.
+in countries where high-quality data may not be collected or made available, or in industries where confidentiality restrictions limit the public release of dollar-denominated expenditure records.
+Consequently, measures of research must rely on counts of researchers (who differ in skill), publications (which differ in relevance and resource-intensity), and administrative databases (which differ in completeness).
+The objective of this paper is to address these conceptual and empirical challenges in the context of medical research.
+Measuring the outputs of the production function (2.1) is, in some sense, more straightforward in medicine than in other sectors.
+Medical innovation appears coarsely in records maintained by regulators (e.g., counts of “new molecular entities” approved by the U.S.
+Food and Drug Administration) and, more directly, in measures of life expectancy and life-years saved from disease.
+[^20] Cutler and McClellan, (2001) observe that technological change has accounted for the bulk of increases in health spending over time, but that health has improved alongside these increases.
+In a series of case studies, they argue that the benefits of technological progress exceed their costs.
+See, also, Almond et al., (2010), who estimate that the returns to medical spending are positive on the margin.
+As this paper focuses on quantification of inputs $S_{t}$, we note only that it is a well-established empirical regularity that improvements in life expectancy, on many margins, appear to be “stagnating” (Mehta et al.,, 2020; Ho and Hendi,, 2018).
+In drawing conclusions about research productivity, we will follow Bloom et al., (2020) and assume that the left-hand side of (2.1) is constant.
+[^21] It is of course unlikely that medical innovations developed in the early 2010s are clearly discernible in recent mortality data, especially given the impact of the COVID-19 pandemic.
+Our task, then, is to identify a measure for inputs $S_{t}$ that best explains potential shifts in this output measure.
+Institutional features of medicine are helpful in doing so: medical decision-makers enforce a clear hierarchy of statistical evidence (Jones and Podolsky,, 2015).
+As guidelines to medical journal authors and editors from the American Medical Association state, “[p]rospective studies generally yield more reliable conclusions than retrospective studies.
+Cause and effect cannot be established by observational studies; consequently, study findings from observational studies should not be presented using causal language.” [^22] In 2024, the Journal of the American Medical Association—the “trunk journal” (see Section 4) of the American Medical Association and among the highest ranked journals in medicine—adopted a framework to allow studies other than randomized experiments to use this type of “causal” language.
+In doing so, editors of JAMA noted that “well-conducted randomized clinical trials remain the preferred approach” for the production of this type of evidence (Dahabreh and Bibbins-Domingo,, 2024).
+In decisions about the approval of new medicines (Food and Drug Administration,, 2024), the terms of insurance coverage (Choudhry,, 2017), and the appropriateness of a particular prescription (Martí-Carvajal,, 2020; Alsan et al.,, 2024), decision-makers defer to evidence generated via clinical trial.
+Other forms of medical evidence, generated through observational studies and surveys of the literature, receive less weight.
+Although detailed cost information on medical research is not available, available evidence suggests that these other forms of medical research are lower cost [^23] Recent estimates suggest that the average clinical trial may cost upwards of \$20 million dollars (Martin et al.,, 2017).
+Available estimates suggest that the cost of completing a systematic review or meta-analysis can reach \$150,000, where the bulk of the cost stems from the quantity of labor required to identify trials in publication data (Michelson and Reuter,, 2019).
+and involve fewer regulatory hurdles (Congressional Budget Office,, 2021).
+[^24] For example, while the approval of an institutional review board is necessary for any randomized clinical trial to proceed, many observational studies are exempt from this type of review.
+See guidance to researchers at Stanford University here: https://stanfordmedicine.box.com/shared/static/ase46ixmatdeahj7qdfgp4ydmy5qc3ei.pdf.
+Thus, we treat clinical trials as the appropriate measure of research inputs $S_{t}$.
+Clinical trials are discrete, regulated experiments.
+Research proceeds according to a pre-defined experimental protocol under the supervision of a regulator and across various phases.
+Trial phases differ in their scale (number of participants and duration) and, thus, in their cost (Food and Drug Administration,, 2024).
+Standardized measures (e.g., statistical objects like p-values) are recorded at various intervals.
+Trials are sufficiently expensive (Congressional Budget Office,, 2021) as to reflect “real” and substantial investment, and decisions to progress a study across phases (e.g., to move from a small Phase II study to a large Phase III study) capture decisions to pursue continued, directed research in an area of study (for an application of this insight, see Cunningham et al.,, 2021).
+As we approach this exercise with the ultimate goal of making progress on the question of research productivity, it is useful to note an additional feature about clinical trials, which suggests that sharp increases in quantities year-on-year (as documented in Bloom et al., (2020)) are unlikely.
+Not only are clinical trials sufficiently expensive as to require access to a source of financing, but trials also require an often scarce “input:” eligible patients.
+Malani and Philipson, (2019) collect empirical evidence of this type of “subject market effect,” and Alsan et al., (2024) collect anecdotal evidence on the difficulties of patient recruitment.
+[^25] The popular press has described this as a conundrum for cancer research: “too many drug trials, too few patients” (Kolata,, 2017).
+3. Identifying Clinical Trials
+We construct a new census of clinical trials disclosed in scientific publications.
+We begin, in Section 3.1, by detailing the deficiencies of existing, or alternative, approaches to measuring this form of medical research.
+In Sections 3.2 and 3.3 we propose a method, based on fine-tuning a large language model, for identifying relevant records at scale and with high fidelity.
+The performance of the method is assessed in Section 3.4.
+3.1. Existing Approaches
+Although many datasets report information on clinical trials—and have been increasingly used by economists to measure innovation (see e.g., Budish et al.,, 2015; Cunningham et al.,, 2021)—no existing dataset, used off-the-shelf, yields accurate, transparent quantity measures for medical research in aggregate.
+3.1.1. Dollar Denominated Data
+Although disaggregated, dollar-denominated trial costs are a natural measure, available records of investment in clinical trials are self-reported, in some form, by pharmaceutical firms and are essentially impossible to validate independently (for a discussion, see Morten et al.,, 2024).
+There are, to our knowledge, three such sources of data that have been used in research.
+DiMasi et al., (2003, 2016) produce one widely cited (and contested, see Light and Warburton, (2005)) set of estimates of drug research and development costs, based on confidential data obtained from 10 firms.
+Myers and Pauly, (2019) and Scannell et al., (2012) collect aggregate estimates of annual research and development expenditures published by the Pharmaceutical Research and Manufacturers of America, a trade organization representing the industry.
+Both sets of estimates are often characterized as an upper bound (Morten et al.,, 2024) and cannot be produced independently by external researchers.
+Wouters et al., (2020) generates estimates of research and development costs, drawing on the information available in financial filings collected by regulators in the United States.
+Although these estimates are important and informative in a context that lacks precise measures of costs, expenditures can be defined in this way for only roughly 18 percent of the new drugs approved by the FDA between 2009 and 2018.
+These estimates, then, are not sufficient to answer the question this paper sets out to address.
+Even if dollar-denominated measures were available, it is not obvious how these measures should be scaled to allow for comparison across time periods.
+See Berndt and Cockburn, (2013) for a discussion of the difficulties associated with constructing a price index for clinical trial research.
+[^26] Berndt and Cockburn, (2013) construct a preliminary price index using data on trial budgets from a private firm.
+This firm no longer discloses data to academic researchers.
+3.1.2. Trial Registries and Proprietary Databases
+Nearly all clinical trials are conducted under the auspices of a regulator that requires both pre-registration and results disclosure.
+In the United States, the Food and Drug Administration Amendments Act (FDAAA), as passed in 2007 and revised in a ”Final Rule” that took effect in 2017, imposes these requirements on sponsors of regulated trials.
+Nominally, effected sponsors are required to register their studies on “ClinicalTrials.gov.” For nearly all studies, sponsors must also report the results of a trial within a designated period after completion.
+Moreover, increasingly, medical journals require registration as a pre-condition of publication (Laine et al.,, 2007).
+In practice, there is widespread noncompliance with these regulations and requirements.
+Deborah Zarin, then-director of ClinicalTrials.gov, notes that “funders, sponsors, and institutional review boards continue to allow unregistered trials or trials with late registration to be conducted, and some journals continue to allow the results of such trials to potentially be published” (Zarin et al.,, 2017).
+Noncompliance with federal reporting requirements is similarly widespread.
+In a series of papers assessing compliance over time DeVito et al., (2020) and DeVito and Goldacre, (2021) find that roughly 60 percent of trials registered on ClinicalTrials.gov fail to report required data elements at the time of their reporting deadline.
+[^27] The FDA is empowered, by statute, to impose fines of up \$10,000 for each day that required disclosures are late.
+To date, no fines have been collected, but the agency now sends notices of noncompliance to delinquent sponsors (Stephenson,, 2021).
+Notices are available here: https://www.fda.gov/science-research/fdas-role-clinicaltrialsgov-information/clinicaltrials gov-notices-noncompliance-and-civil-money-penalty-actions
+From the perspective of our study, there are three issues with the use of registry data.
+First, requirements for study registration have changed over time.
+Table B.1 lists five changes between 2000 and 2020 that altered the set of studies that should appear in a complete version of the database.
+Second, sponsors engage in considerable “back-filling” of past studies.
+Zarin et al., (2017) find that a large proportion of trials are registered after the trial start date.
+[^28] Often back-filling relates to changes in organizational disclosure policy.
+In a version of ClinicalTrials.gov current through 31 May 2023, Boehringer Ingelheim—the largest privately held pharmaceutical company in the world—appears to have posted 804 studies in 2014, only 72 of which were started in 2014.
+In 2013, Boehringer registered 101 studies.
+In the fifteen preceding years, since the creation of the database, Boehringer posted 872 records in total.
+We selected this case study based on a discussion in Zarin et al., (2017).
+Third, not all records in ClinicalTrials.gov and similar registries are, in fact, clinical trials (Tse et al.,, 2018).
+As a result, it is difficult to determine whether changes in the number of records reflect actual growth in scientific research or a change in the propensity to register studies (trials and non-trial studies).
+In Appendix B, we illustrate the difficulty of recovering accurate measures of trends in the quantity of clinical trials from this database.
+We show that small changes in data construction yield meaningfully different estimates of trial levels and that many records are not in fact clinical trials.
+Moreover, due to widespread data gaps resulting from noncompliance in reporting of results and study attributes, there is little reliable information about characteristics of registered studies.
+Other databases, such as the PDQ Cancer Information database maintained by the U.S.
+National Cancer Institute and used in papers including Budish et al., (2015), are disease-specific and, thus, insufficiently general for our purposes.
+A small number of firms (notably, GlaxoSmithKline) have disclosed all clinical trials over a longer period of time.
+[^29] See GlaxoSmithKline’s registry here: https://www.gsk-studyregister.com/ Of course, records from a single firm are not informative about aggregate trends, especially because many investigational drugs change hands frequently during the course of development (Cunningham et al.,, 2021).
+Researchers, instead, often draw on proprietary databases.
+Although these databases are sufficient for certain research questions—for example, related to changes in the ownership of investigational drugs (Cunningham et al.,, 2021, using Pharmaprojects data) or on the link between insurance policy and drug development (Agha et al.,, 2022, using Cortellis data)—they are unsuitable for our particular productivity measurement exercise.
+Our aim is to construct a census that allows for comparison of quantities across points in time.
+First, we are unable to determine, with specificity, how records are identified and catalogued based on available documentation.
+Second, these databases take public registries as inputs and, thus, risk adopting the same shifts in reporting over time (see e.g., the list of registry inputs for one frequently-used database, Cortellis,, 2024).
+3.1.3. Publication Data
 Publication data provide a natural alternative.
+[^30] In Section 4, we discuss the two natural concerns with measures constructed from publication data—publication bias and shifts in reporting norms over time.
+Neither is likely to affect the interpretation of trends in our data.
 The database PubMed / MEDLINE indexes roughly $34$ million publications, a near universe of published biomedical research.
 [^31] See Section A.1 for further details on on our treatment of the PubMed data.
-Scientific publications provide an intuitive proxy for the quantity of research inputs.
-[^13] The existence of publication bias means that scientific publications do not perfectly capture research inputs.
-See Andrews and Kasy, (2019) for a discussion.
-In a setting such as ours—where objects of interest are trends in research over time—we need only assume that changes in publication bias year-on-year do not perfectly offset changes in research investment.
+It is reasonable to expect that the corpus of scientific publications contains a census of research.
+The challenge is in extracting measures of research from this unstructured text.
 Of these $34$ million records, which are clinical trials?
@@
 Both approaches retrieve large numbers of records that are not, in fact, clinical trials, while excluding potentially relevant records.
-Figure 1 contrasts counts of published trials constructed in various ways.
-The light green line (‘Clinical Trial’ NLM Tag) plots the number of publications, in each calendar year, indexed with the publication type “clinical trial.” The teal line (Any NLM Tag) plots counts of records indexed with any of 18 types that are likely to include clinical trials or related medical research.
-[^15] On inspection, it appears as though the NLM shifted from using the tag “Clinical Trial” to “Randomized Controlled Trial”, as well as more specific tags (e.g., “Clinical Trial, Phase 1”) .
+Figure 3.1 contrasts counts of published trials constructed in various ways.
+The light green line (‘Clinical Trial’ NLM Tag) plots the number of publications, in each calendar year, indexed with the publication type “clinical trial.” This light green line is the measure reported in Bloom et al., (2020).
+The teal line (Any NLM Tag) plots counts of records indexed with any of 18 types that are likely to include clinical trials or related medical research.
+[^33] On inspection, it appears as though the NLM shifted from primarily using the tag “Clinical Trial” to “Randomized Controlled Trial” and added more specific tags (e.g., “Clinical Trial, Phase 1”) .
+There are 19,937 records tagged with “Clinical Trial” in 2005 and 7,518 with this same tag in 2006.
 Alternative approaches to identifying trials are similarly imprecise.
@@
 The purple line (Potential Universe) plots this trend over time.
-Figure 1 highlights that alternative approaches to data construction, in this setting, yield meaningfully different conclusions about levels, trends, and composition.
+Figure 3.1 highlights that alternative approaches to data construction, in this setting, yield meaningfully different conclusions about levels, trends, and composition.
+That few of these series suggest “smooth” trends in the quantity of trials indicates, also, the extent to which the quality of data can affect these estimates.
 See Section A.2 for further details on the construction of these series.
-Figure 1. Universe of Potential Clinical Trials
-Notes: Figure 1 displays counts of the number of clinical trials indexed in PubMed / MEDLINE over time, constructed using alternative search strategies.
+Given these difficulties, we describe a procedure, below, to recover a sample of clinical trials from this corpus of text in a manner that is consistent and transparent.
+Figure 3.1. Universe of Potential Clinical Trials
+Notes: Figure 3.1 displays counts of the number of clinical trials indexed in PubMed / MEDLINE over time, constructed using alternative search strategies.
 The National Library of Medicine (NLM) categorizes each publication into a “pubtype.” The light green line (‘Clinical Trial’ NLM Tag) displays the number of publications in the “Clinical Trial” pubtype.
@@
 Studies involving animals, literature summaries, and re-analyses of existing data, for example, are excluded.
-We take the “potential universe” plotted in Figure 1 as our baseline sample of potential clinical trials.
-We further restrict attention to the 1,821,429 records published in or after 2010.161616There is a trade-off between data coverage and quality.
-Around 2010, a set of policy changes—including trial registration requirements and harmonization in publication formatting—affected trial disclosure in scientific publications.
-See Laine et al., (2007) and Schulz et al., (2010) for examples of two such changes.
+We take the “potential universe” plotted in Figure 3.1 as our baseline sample of potential clinical trials.
+We further restrict attention to the 1,821,429 records published in or after 2010, as there is a trade-off between data coverage and quality in this setting.
+In the 1990s and 2000s, medical experts created a standardized system for clinical trial reporting in scientific publications.
+By the time of the most recent Consolidated Standards of Reporting Trials (CONSORT) in 2010, nearly all medical journals required authors to comply with a standardized structure for all clinical trial publications (Schulz et al.,, 2010).
+As a result, most trial publications have a distinctive format and include a pre-defined set of elements.
+Although our procedure can be applied to any type of scientific publication text, the standardization required by the CONSORT Statement changes the relationship between clinical trial research and publications in the early 2000s.
+To answer the research question that motivates this paper—is the quantity of medical research rising over time?—we elect to proceed with a narrower sample to avoid this issue of interpretability.
 To quantify our ability to identify the clinical trials in this set of publications, we hand-label approximately 3,000 randomly selected publications based on the content of their abstracts.
-[^17] We develop a custom labelling interface that reduces the time required to label 100 records, for the authors, by a factor of seven.
-Section B.1 describes this tool and its application to this context in detail.
-Of the hand-labelled publications, 11.2% meet the criteria for inclusion in our sample.
-The hand-labelled data are split into three subsets—validation, training, and testing—based on their eventual use.
+We develop a custom labeling interface that reduces the time required to label 100 records, for the authors, by a factor of seven.
+Section C.1 describes this tool and its application to this context in detail.
+Of the hand-labeled publications, 11.2% meet the criteria for inclusion in our sample.
+The hand-labeled data are split into three subsets—validation, training, and testing—based on their eventual use.
 As a baseline, we assess the performance of several standard machine learning algorithms.
-The results are displayed in Figure 2.
-Further details on the construction of this figure are given in Section B.2.
+The results are displayed in Figure 3.2.
+Further details on the construction of this figure are given in Section C.2.
 Each algorithm is estimated on the hand-labels assigned to the publications in the training and validation sample splits.
@@
 At a 90% true positive rate, the best performing model identifies 50 true positives for every 50 false positives.
-Figure 2. Performance of Standard Machine Learning Methods
-Notes: Figure 2 displays the receiver operating characteristic of several machine learning models, trained on two types of embeddings, for classifying whether a publication satisfies the restrictions enumerated in 3.1.
+Figure 3.2. Performance of Standard Machine Learning Methods
+Notes: Figure 3.2 displays the receiver operating characteristic of several machine learning models, trained on two types of embeddings, for classifying whether a publication satisfies the restrictions enumerated in 3.1.
 The hyper-parameters of each model are determined with 10-fold cross-validation.
@@
 In addition, we display the performance of our, ensemble, fine-tuned large language model in purple.
-This is the same curve labelled “Ensemble” in Figure 3.
+This is the same curve labelled “Ensemble” in Figure 3.3.
 3.3. Prompt Design and Fine-Tuning
@@
 First, we iteratively construct a set of prompts that exhibit good performance when posed to proprietary models—OpenAI’s GPT-3.5 and GPT-4.
+A priori, we expect GPT-4 to output labels of slightly lower quality than hand-labeled output and GPT-3.5 to produce slightly noisier labels than GPT-4.
+The extraction of GPT-3.5 labels is substantially faster and cheaper.
 Second, we extract noisy labels for a large number of publications in our sample by querying the proprietary models.
@@
 3.3.1. Prompt Design
-We devise three types of prompts and measure their performance on samples of abstracts taken from our validation dataset.
-We measure model performance with the hand-labeled data, conduct error analyses, and iteratively revise prompts.
-When modifications to the prompt cease to yield improvements in performance, we finalize prompts best suited for GPT-3.5 and GPT-4, respectively.
-Section B.3 provides more detail on this process.
-Figure 3 displays estimates of the true and false positive rates computed in the test data, for both proprietary models, with black dots.
-Figure 3. Receiver Operating Characteristic
-Notes: Figure 3 displays estimates of the receiver operating characteristic of several models used to classify the publications indexed by PubMed / MEDLINE according to whether they satisfy the restrictions enumerated in 3.1.
+To produce these labels at scale, each of the two models must be appropriately prompted.
+That is, each pre-trained model must be provided with a block of text as an input and asked to return the appropriate completion of this input.
+A priori, it is not clear what prompt structure will work well.
+The relative infancy of this area of research renders it difficult to identify a set of “best practices.” [^34] See, for one example of this evolving area of work, a guide to “prompt engineering” from OpenAI: https://platform.openai.com/docs/guides/prompt-engineering.
+All of the prompts that we consider are displayed as figures in Appendix E.
+We identify three general prompt formats, which differ both in the amount of detail provided about our classification task and in the structure of the requested model completion.
+The simplest prompt provides a version of our sample definition, 3.1, and the text of an abstract, and asks the model to return ‘TRUE’ if the abstract satisfies these criteria.
+Otherwise, it returns ‘FALSE’.
+The text of this prompt is displayed in Figure E.1.
+We refer to this prompt as Prompt 1.0.
+A second, more complicated prompt provides the same definition, with a set of examples of publication characteristics that do and do not satisfy the definition.
+Here, the prompt asks the model to return either ‘TRUE’ or the name of a specific excluded category.
+The text of this prompt is displayed in Figure E.7.
+We refer to this prompt as Prompt 2.0.
+A third, more complex prompt provides the same definitions and examples, but asks the model to return either ‘TRUE’ or an explanation of why the record does or does not satisfy our sample definition.
+The text of this prompt is displayed in Figure E.10.
+We refer to this prompt as Prompt 3.0.
+We devise initial language for each prompt iteratively, using small numbers of records from our hand-labelled validation data.
+We test each of these three prompts in our 1000-record hand-labelled validation dataset, using both GPT-3.5 and GPT-4.
+We conduct a detailed error analysis, reported in LABEL:tab:Error-Analysis.
+For each instance in which a model returns a label (TRUE/FALSE) that differs from that in the hand-labelled dataset, we inspect the record.
+We categorize errors into types and sub-types.
+Details are given in Section C.3.
+This exercise is instructive on three margins.
+First, it highlights the highest performing prompts.
+Second, it indicates particular types of errors in categorization—which suggest opportunities for more precise language in a prompt.
+Third, it draws attention to differences in the performance of the two models.
+Observe that, in the “Other” error type, we include a sub-type called “overly literal interpretation of inclusion criteria.” We primary record errors of this type for Prompt 3, which asked the model to return a completion that described why, or why not, a record was classified as being in our sample.
+Here, GPT-3.5 makes 30 such errors—classifying records as FALSE by recapitulating the sample definition provided.
+GPT-4 makes two errors.
+The insight that emerges from this exercise, then, is that different prompt structures may be preferable depending on the model used, and that, by and large, GPT-4 is better suited for tasks that require (the resemblance of) more sophisticated reasoning.
+We revise each class of prompt based on these findings.
+For Prompt 1.0, the simplest true/false prompt, we consider three variants.
+We refer to these revised prompts as Prompts 1.1, 1.2, and 1.3.
+The text of these prompts is displayed in Figures E.2, E.3, and E.5.
+For Prompts 2 and 3, we consider one variant each.
+We refer to these revised prompts as Prompts 2.1 and 3.1.
+The text of these prompts is displayed in Figures E.8 and E.11.
+These changes reflect the differences in performance catalogued in LABEL:tab:Error-Analysis.
+Table 3.1 reports estimates of the true positive rate and false positive rate for each prompt, in both models, computed in the validation sample.
+Observe—for example, with Prompts 1.1, 1.2, and 1.3 queried to GPT-4—that even small changes in the text of a prompt yield substantial differences in performance.
+We attempt a second round of prompt iteration (Prompt 1.3).
+We observe that performance deteriorates with even small modification.
+Thus, we select the two highest performing prompts—one for each proprietary model.
+We use Prompt 2.0 to extract weak labels using GPT-3.5, and Prompt 1.2 to extract weak labels using GPT-4.
+Figure 3.3 displays estimates of the true and false positive rates computed in the test data, for both proprietary models, with black dots.
+These prompts are used to extract 64,000 labels—TRUE/FALSE flags—using GPT-3.5 and GPT-4.
+We refer to these as our set of “noisy labels.”
+Table 3.1. Prompt Performance in Validation Data
+Panel A: GPT-3.5
+Prompt Type Prompt Sub-Type False Positive Rate True Positive Rate
+1 0 0.247 0.788
+1 0.172 0.898
+2 0.037 0.584
+3 0.037 0.489
+2 0 0.162 0.876
+1 0.176 0.905
+3 0 0.248 0.722
+1 0.171 0.883
+Panel B: GPT-4
+Prompt Type Prompt Sub-Type False Positive Rate True Positive Rate
+1 0 0.202 0.971
+1 0.065 0.949
+2 0.049 0.934
+3 0.056 0.912
+2 0 0.081 0.964
+1 0.072 0.964
+3 0 0.167 0.971
+1 0.068 0.956
+Notes: Table 3.1 records the performance of each of eight prompt variants in a sample of 1,000 validation dataset records.
+Panel A reports performance associated with the proprietary model GPT-3.5.
+Panel B reports analogous statistics for model GPT-4.
+Prompt 1 asks the model to return TRUE or FALSE.
+Prompt 2 asks the model to return TRUE or the name of a specific, excluded category.
+Prompt 3 asks the model to return TRUE or an explanation of why the record should be excluded.
+Prompt sub-types correspond to various iterations.
+Sub-type 0 is the initial version of the prompt.
+Sub-types 1-3, where applicable, are subsequent iterations.
+Appendix E records the text of each prompt.
+We use Prompt 2, Sub-Type 0, to extract weak labels using GPT-3.5, and Prompt 1, Sub-Type 2 to extract weak labels using GPT-4.
+The performance measurements for these prompts are displayed in boldface.
+Figure 3.3. Receiver Operating Characteristic
+Notes: Figure 3.3 displays estimates of the receiver operating characteristic of several models used to classify the publications indexed by PubMed / MEDLINE according to whether they satisfy the restrictions enumerated in 3.1.
 All metrics are computed in the testing split of the hand-labelled publications.
@@
 We observe no large differences in performance stemming from these hyper-parameters.
-Sections B.4 and B.5 give further details the fine-tuning process and the results of these experiments, respectively.
+Sections C.4 and C.5 give further details concerning the fine-tuning process and the results of these experiments, respectively.
 3.4. Performance
@@
 Publications whose probabilities fall above a chosen threshold are classified as belonging in our sample.
-Figure 3 displays estimates of the true and false positive rates, computed with the test data, as we vary this threshold for both sizes of the RoBERTa and BioMedBERT models.
-[^20] Each of the models whose results are displayed in Figure 3 are trained with noisy labels obtained with GPT-4.
-We document in Section B.5 that models trained with noisy labels obtained from GPT-3.5 or with the 1000 hand-labeled records in the training data perform significantly worse.
+Figure 3.3 displays estimates of the true and false positive rates, computed with the test data, as we vary this threshold for both sizes of the RoBERTa and BioMedBERT models.
+[^37] Each of the models whose results are displayed in Figure 3.3 are trained with noisy labels obtained with GPT-4.
+We document in Section C.5 that models trained with noisy labels obtained from GPT-3.5 or with the 1000 hand-labeled records in the training data perform significantly worse.
 The figure additionally displays the performance of an ensemble model estimated in the training data.
-This model is obtained by estimating a logistic regression of the hand-labels on the probabilities output by all four models displayed Figure 3.
+This model is obtained by estimating a logistic regression of the hand-labels on the probabilities output by all four models displayed Figure 3.3.
 The ensemble model is used to produce our final sample.
@@
 We choose three thresholds according to the stringency with which they enforce the sample restrictions.
-The estimated true and false positive rates associated with these points are displayed in Figure 3 and labeled as “Conservative,” “Moderate,” and “Liberal.” Our preferred sample is associated with the conservative threshold.
+The estimated true and false positive rates associated with these points are displayed in Figure 3.3 and labeled as “Conservative,” “Moderate,” and “Liberal.” Our preferred sample is associated with the conservative threshold.
 [^38] The liberal model may be particularly useful for conducting literature reviews, where a near-perfect true positive rate is needed.
@@
 We conduct an error analysis.
-See Section B.6 for further details.
+See Section C.6 for further details.
 In 13 cases, there is a clear error.
@@
 Our final, conservative, sample consists of 152,027 publications classified as satisfying the restrictions enumerated in 3.1.
-Appendix A compares the contents of our sample to the counts of potential clinical trials plotted in Figure 1.
+Each table displayed in Appendix A compares the contents of our sample to the counts of potential clinical trials plotted in Figure 3.1.
 Although certain elements of PubMed metadata—including the union of all NLM tags—capture many of the records in our final sample, we confirm that they miss many records that we flag as trials and include many records that do not satisfy 3.1.
 No combination of existing search strategies, then, classifies records with the same accuracy or precision as our final sample.
-4. Medical Research as a Productivity Indicator
-Figure 1 suggests that, across metrics, there has been a substantial increase in the quantity of medical research over time, including since 2010.
-This is consistent with documented increases in the quantity of scientific papers in Bloom et al., (2020) and Park et al., (2023), as well as economy-wide increases in research effort reported in Goldin et al., (2024).
-4.1. Decomposing Medical Research
-We decompose this pattern, into changes attributable to high-productivity, high-capital investments—clinical trials—and changes attributable to other forms of clinical research.
-Our census of clinical trials provides measures for the first category.
-Although there are many reasonable ways to construct samples of other clinical research, our preferred approach defines this category as those scientific publications that cite clinical trials.
-[^23] Alternative approaches to construction of this sample, as in Figure 1, also suggest an increase in the quantity of non-clinical trial research.
-We prefer to use this citation-based definition because it allows for consistency in sample definition.
-Other approaches that rely on keyword searches, NLM tags, etc., as Section 3 and its appendices make clear, capture different sets of records over time.
-These are records of research investment that necessarily engage with findings from trials (such as meta-analyses and re-analyses of existing data), but, as discussed in Section 2, are essentially always lower productivity and less capital-intensive.
-Figure 4 documents that the quantity of clinical trials—the black line—has remained constant over the past ten years.
-In contrast, other clinical research has increased substantially.
-Observe that we display the percent growth in the quantity of clinical trials in our sample published in each calendar year, alongside the growth in the set of papers that cite clinical trials.
+4. Disentangling Trends in Medical Research
+Figure 3.1 indicates that, across measures, the total quantity of medical research appears to be increasing, including since 2010.404040We find similar patterns using alternative cuts of records indexed in PubMed.
+The total number of unique authors listed on scientific papers, the total number of citations given and received by scientific papers, the total number of scientific journals, and the total number of pages in scientific journals all exhibit similar patterns.
+These patterns are consistent with documented increases in the quantity of scientific papers in Bloom et al., (2020) and Park et al., (2023), as well as economy-wide increases in research effort reported in Goldin et al., (2024).
+In this section, we use our census of clinical trials to decompose this trend.
+We contrast trends in the production of clinical trials with those documented for other forms of medical research.
+Differences between these series are informative in two ways.
+First, existing estimates of the quantity of medical research, such as those plotted in Figure 3.1, capture some combination of these two series (e.g., Bloom et al.,, 2020; Park et al.,, 2023; Chu and Evans,, 2021).
+Second, any proxy for the quantity of medical research is difficult to interpret if its relationship to the underlying research changes over time.
+This decomposition highlights the limitations of empirical strategies that treat counts of scientific publications as a proxy for research quantity.
+There are many potential ways to construct samples of other (i.e., non-clinical trial) medical research.
+Our preferred approach defines this category as the collection of publications that cite clinical trials.
+(We confirm in Section 4.3.1 that this set includes only medical research.) Conceptually, of course, this “citing” sample is a subset of non-clinical trial medical research.
+Alternative methods of constructing a measure of “other” research introduce additional, complicating conceptual problems.
+Approaches that rely on keyword searches, NLM tags, etc., as Section 3 and its appendices make clear, capture different sets of records over time, as a consequence of changes in database structure.
+Approaches that examine other publications in the same journals as clinical trials will mechanically suggest shifts in total quantities, if the number of pages in journals change over time or if certain journals become online-only.
+[^41] Ioannidis et al., (2023) document that the number of medical journals is increasing.
+In 2024, the Annals of Emergency Medicine, the highest impact journal in the medical sub-field of emergency medicine, shifted to an online-only format.
+Brainard, (2020) discusses dozens of instances in which medical journals transitioned to online-only formats, then ceased publication entirely.
+The primary challenge in using a citation-based measure is one of truncation: constructing measures of papers that cite clinical trials requires determining the window in which such citation must have occurred.
+To account for concerns about citation delays, we construct various measures, using 2–6 year citation counts.
+Two-year citation counts, thus, count publications that cite clinical trials published in the preceding two years.
+As our clinical trial sample begins in the year 2010, these measures are available only beginning in 2012.
+It is possible, and common, for clinical trials to cite other trials.
+We drop from our “citing” sample any records that we also flag as reporting the results of a trial.
+4.1. Decomposing Types of Medical Research
+Figure 4.1 documents that the quantity of clinical trials—the black line—has remained essentially constant over the past ten years.
+In contrast, the quantity of other medical research—the colored lines—has increased substantially.
+Observe that we display the percent growth in the quantity of clinical trials in our sample published in each calendar year, alongside the growth in the set of papers that cite clinical trials, constructed in various ways to account for the truncation discussed above.
 Each series is normalized to begin at zero.
-To account for concerns about citation delays and truncation, we construct various measures of papers that cite clinical trials, using 2–6 year citation counts.
-Two-year citation counts, thus, are available starting in the year 2012 (two years into our sample) and count publications that cite clinical trials published in the preceding two years.
-[^24] Figure C.3 and Figure C.4, displayed in the supplemental appendix, document that this trend is robust to alternative sample definitions, using either our liberal or moderate samples, and to citation-weighting count data.
 This presentation masks substantial differences in the levels of these series.
-There are 10,903 clinical trials in our census in 2010, and 73,253 publications that cite clinical trials (using two-year citations) in 2012.
-Figure 4. Growth in Clinical Research, Stability of Clinical Trials
-Notes: Figure 4 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year.
+There are 10,903 clinical trials in our census in 2010, and 30,841 publications that cite clinical trials (using three-year citations) in 2013.
+Figure 4.1. Growth in Clinical Research, Stability of Clinical Trials
+Notes: Figure 4.1 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year.
 Each series is reported in terms of the percent change relative to its initial value.
@@
 To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
-4.2. Stable Composition of Clinical Trials
-To contextualize the trends displayed in Figure 4, it is helpful to return to the simple model from Section 2.
-We document that the quantity of capital-intensive, high-productivity research inputs remains unchanged over our period of interest, while low-capital, low-productivity inputs increase.
-As changes in outputs in this sector (mortality) are assumed to remain constant (see Section 2), there are thus three ways of reconciling these facts in our stylized framework: (i) other clinical research is sufficiently low capital or low productivity, across the board, as to have essentially no impact on growth, (ii) increases in other clinical research have, heterogeneously, low capital or low productivity, or (iii) clinical trial research productivity is declining over this time period, fully offsetting the growth in other types of research.
-If either Case (i) or Case (ii) were an accurate characterization, rising observed research effort—increases in the number of scientific publications or the number of records in PubMed that the previous methods have identified as clinical trials—need not imply diminishing returns to high-capital investment.
-Our data allow us to reject the hypothesis that we are in Case (iii).
-Figure 5 documents that—on four dimensions—measures of trial productivity, across clinical trials, have remained stable over the past decade.
-Panel A of Figure 5 considers the share of publications in our census with the following characteristics: any funding from a government agency (purple), any (three-year) citations from a leading journal in medicine (teal), or any citations from a leading journal conditional on having public funding (green).
-[^25] We designate a publication as having public funding if it appears in the National Institutes of Health RePORTER data, as linked to a funded research grant, or if PubMed reports a source of research funding, including funding from non-U.S.
-agencies.
-To designate a set of the “top” journals in medicine, we follow Angrist et al., (2020) We collect all citations originating from the “trunk journals” in medicine, the Journal of the American Medical Association and the New England Journal of Medicine, to records in PubMed between 2010 and 2022.
-We designate a journal as being “leading” of it received at least 100 citations from a trunk journal over this time period.
+These trends are robust to alternative forms of data construction.
+In Figure D.1, we construct these plots using our liberal and moderate sample definitions and find, qualitatively, no change.
+In Figure D.2, we explore a potential concern with a citation-based measure.
+We document in Figure 3.1 that the quantity of scientific publications is rising over time.
+It is reasonable to expect that this rise, coupled with the ease of citing papers enabled by the internet, coincides with an increase in the total number of citations in the aggregate and per paper.
+We confirm, empirically, that both hypotheses are borne out by our data.
+To account for the fact that the “cost” of citing a clinical trial may have decreased over time, thus generating the trend documented in Figure 4.1, we construct a version of this plot in which each citing paper is weighted by the total number of papers that it cites.
+That is, a citing paper that cites a trial and one other paper receives a weight of 1/2, and a citing paper that cites a trial and nine other papers receives a weight of 1/10.
+In doing so, we “penalize” papers that are especially generous in their tendency to cite other papers.
+Figure D.2 suggests, again, that this trend is robust to this alternate construction.
+4.2. Clinical Trials as a Productivity Indicator
+The trends in Figure 4.1, on their own, allow us to draw two conclusions.
+First, existing estimates that suggest sharply rising quantities of clinical trials (e.g., Bloom et al.,, 2020) are a consequence of technical challenges in the measurement of research.
+Second, existing estimates of rising quantities of medical and scientific research (e.g., Bloom et al.,, 2020; Chu and Evans,, 2021; Park et al.,, 2023) are not driven by shifts in the quantity of clinical trials.
+Recall the very simple conceptual framework from Section 2.
+We assume, following e.g., Bloom et al., (2020), that the left-hand side of the “idea production function” (2.1) is constant in this setting.
+If clinical trials, as captured in our data, are a stable proxy for inputs $S_{t}$—in the sense that published trials capture the same amount of underlying research in each period—then an unchanging number of clinical trials implies that research productivity is constant.
+We return to the appropriate interpretation of rising trends in other medical research below.
+Here, we collect evidence that clinical trials are a stable proxy for inputs.
+Figure 4.2 documents that—on two additional dimensions—clinical trial production has been stable over the past decade.
+In particular, Figure 4.2 displays the share of publications in our census with the following characteristics: any funding from a government agency (purple), any (three-year) citations from a leading journal in medicine (teal), or any citations from a leading journal conditional on having public funding (green).
+Figure 4.2. Public Funding and Never-Cited Research
+Notes: Figure 4.2 displays three time series: the proportion of clinical trials that are publicly funded, the proportion of clinical trials that are ever cited by a leading journal, and the proportion of publicly funded clinical trials that are ever cited by a leading journal.
+The sample of published clinical trials is constructed with the conservative model.
+We designate a publication as having public funding if it appears in the National Institutes of Health RePORTER data, as linked to a funded research grant, or if PubMed reports a source of U.S.-based research funding.
+[^42] Technically, PubMed indicates whether a publication received research funding from government agencies and private philanthropies outside of the United States.
+We restrict consideration to those sources of funding based in the United States because there appears to be a shift in the reporting of certain sources of non-U.S.
+funding beginning in 2017.
+The total number of publications linked to a source of public funding increases sharply—by roughly 50 percent from 2016—in 2017.
+On inspection, other trends are smooth during this time, including trends in public funding for the United States.
+We are unable to determine if certain sets of authors began to disclose sources of funding in 2017, if certain journals changed their reporting requirements, or if PubMed changed its data construction process.
+We thus focus on the United States, where we find no evidence of large changes in data reporting practices.
+To designate a set of the “top” journals in medicine, we follow Angrist et al., (2020).
+We collect all citations originating from the “trunk journals” in medicine, the Journal of the American Medical Association and the New England Journal of Medicine, to records in PubMed between 2010 and 2022.
+We designate a journal as being “leading” if it received at least 100 citations from a trunk journal over this time period.
 This yields a list of 84 journals.
-See Section A.3 for details.
-Appendix C reproduces this figure using five-year citations.
+[^43] See Section A.3 for details.
+Appendix D reproduces this figure using five-year citations.
 The trend is unchanged.
-The trends in Panel A communicate two findings.
-First, all three proxies for research productivity are stable, essentially unchanging, over our time period.
-Second, nearly 55 percent of clinical trials are never cited by a leading journal, suggesting substantial, but time-invariant, heterogeneity across trials.
-Figure 5. Stability of Heterogeneity
-Panel A: Public Funding and Low-Quality Research
-Panel B: The Composition of Quality over Time
-Notes: Figure 5 illustrates the stability of the heterogeneity in the composition and quality of published clinical trials between 2010 and 2019.
-Panel A displays three time series: the proportion of clinical trials that are publicly funded, the proportion of clinical trials that are ever cited by a leading journal, and the proportion of publicly funded clinical trials that are ever cited by a leading journal.
-Panel B displays two heat maps measuring the distribution of citations received by clinical trials in each calendar year from leading journals and from all publications, respectively.
-The $y$-axis has been stretched to elongate the right-tail of the citation distribution.
-Colors are displayed in a log scale.
-For both panels, the sample of published clinical trials is constructed with the conservative model.
-These proxy variables capture slightly different conceptual objects.
-Interpreting each, separately, provides suggestive evidence on different aspects of the research production function.
+Two facts are worth highlighting.
+First, all three measures are stable, essentially unchanging, over our time period.
+Second, nearly 70 percent of clinical trials are never cited by a leading journal, suggesting substantial, but time-invariant, heterogeneity across trials.
+These proxy variables capture slightly different conceptual objects, which are helpful in understanding different aspects of the production function (2.1).
 Stability in public funding suggests minimal changes, at least in the cross section, in allocation of resources to clinical trials over time.
@@
 Note, however, that linking research grant dollars to specific scientific papers is an imprecise exercise.
-See Li, (2017) for an extended discussion of this disconnect.
-Stability in citation patterns, in contrast, suggests that the “usefulness” of information produced by trials also remains stable.
-Panel B of Figure 5 provides a more granular view of this heterogeneity.
-For each publication in our census, we collect citations from publications in leading journals and from all of PubMed.
-We plot the total (three-year) citation counts to publications at each citation quantile over time.
-[^27] Appendix C presents these results using alternative cuts of our data and five-year citations.
+See Li, (2017) for an extended discussion of this measurement challenge.
+Stability in citation patterns, in contrast, suggests that the “usefulness” of information produced by trials has been stable across this period.
+Figure 4.3 provides a more granular view of this heterogeneity.
+We plot quantiles of the distribution of the number of (three-year) citations to publications in our census from publications in leading journals, for each year between 2010 and 2019.454545Appendix D presents these results using alternative cuts of our data and five-year citations.
 The patterns remain unchanged.
@@
 This distribution is stable over our time period.
-[^28] Reassuringly, Panel B indicates that this exercise does capture large shifts in the research environment: citation counts appear especially high in the later years of data, when three-year citation counts (beginning for year 2017) capture research written during the COVID-19 pandemic.
-Specifically, we document that the tails of the citation distribution are unchanging.
-Observe that we elongate the y-axis, to highlight the right-tail.
-Citations received at the top of this distribution are constant across periods.
+Observe that we elongate the y-axis to highlight the right-tail of the citation distribution.
+The quantiles corresponding to the right tail of the citation distribution are constant across years.
 The left-tail is similarly stable: the share of records receiving zero citations is essentially constant, year-on-year.
-This result is consistent with the idea that, for clinical trials, the right-hand side of the production function, (2.1), in our simple framework—i.e., $\alpha_{t}S_{t}$—is both heterogeneous and stable over time.
-For this pattern to be consistent with a systematic decrease in the research productivity of clinical trials, it must be the case that there is an exactly offsetting increase in the distribution of research capital.
-Although we cannot reject that there are shifts in upstream, basic inputs into clinical trials or large changes in factor prices during this period, there is little credible evidence consistent with either during this time period.
+Figure 4.3. Citation Distribution Across Time, Clinical Trials
+Notes: Figure 4.3 displays a heat map measuring the quantiles of the distribution of citations received by clinical trials in each calendar year from leading journals.
+The $y$-axis has been stretched to elongate the right-tail of the citation distribution.
+Colors are displayed in a log scale.
+The sample of published clinical trials is constructed with the conservative model.
+Figure D.4 is constructed analogously, where instead we display quantiles of the distribution of the number of citations received from all publications indexed in PubMed.
+There, we find that the right-tail of this distribution is increasing over time.
+In other words, as the total number of publications is increasing, the total number of citations received by clinical trials is increasing.
+However, these citations are not originating in leading journals.
+We interpret this set of facts as evidence that our measured quantity of clinical trials captures an object that, though heterogeneous, is stable over time.
+Thus, suggestively, research productivity is similarly stable.
+We cannot reject that there have been shifts in upstream, basic inputs into clinical trials or changes in factor prices during this period, which might suggest that the underlying quantity of resources directed toward research is, in fact, changing over time.
+In our view, however, there is little rigorous evidence that supports such an argument.
 A small number of papers examine frictions in markets for basic research.
 See, for example, Hill and Stein, (2021) and Myers, (2020).
-To our knowledge, this literature has not provided evidence of substantial shifts in these frictions over time, which in turn would increase the cost or decrease the productivity of basic science.
-[^29] Anecdotal evidence suggests that there may have been substantial decreases in the cost of basic science during this time period.
+To our knowledge, this literature has not provided evidence of substantial shifts in inputs or factor prices over time, which in turn would increase the cost or decrease the productivity of basic science.
+[^46] Anecdotal evidence suggests that there may have been substantial decreases in the cost of basic science during this time period, which would suggest the opposite conclusion in our simple framework—that research might be more productive over this period.
 See Scannell et al., (2012).
@@
 Any measured increases in these expenditures that cannot be independently verified should be viewed with caution, especially in light of our findings of stability.
-More generally, any type of cancellation—an argument against the conclusion that research productivity across clinical trials is stable—would appear to be a knife-edge case.
-In our view, it is far more likely that the distributions of both research productivity and research capital within clinical trials have been constant over this time period.
-5. Discussion
-All research investments are not alike.
-Certain investments are more likely to yield innovations that have welfare-relevant impacts.
-Others, while potentially valuable contributions to a research ecosystem, may have insignificant impacts on growth.
-We argue that this type of heterogeneity is quantitatively significant in medical research.
-We use an open-source, fine-tuned large language model to construct a novel census of clinical trials from 2010 to the present.
-We contrast stable trends in clinical trials—in quantity, quality, and composition—with substantial growth in the quantity of other forms of medical research.
-Our findings rationalize, in this field, two stylized facts about research investments and innovative output: the quantity of research inputs has grown substantially, while improvements in mortality continue at the same stable rate.
-A second dimension of heterogeneity is helpful in contextualizing our findings: geography.
-Figure 6 displays trends in the annual quantity of published clinical trials, and papers citing clinical trials, disaggregated by the location of the first-listed author.
+[^47] Herper, (2022) collects anecdotal evidence, which has been widely cited, on sharply rising costs of clinical trials.
+Although the magnitudes included in this discussion are striking, they capture the impressions of a small number of industry stakeholders and, to our knowledge, have not been verified in a more detailed analysis.
+For clinical trials to be an informative proxy for the quantity of medical research—and for the black line in Figure 4.1, then, to capture the relevant measure of inputs—we must make two additional assumptions.
+First, the existence of publication bias means that scientific publications do not perfectly capture research inputs (Andrews and Kasy,, 2019).
+In our setting—where the object of interest is the trend in research over time—we need only make a relatively weak assumption about the extent of publication bias: we assume that changes in publication bias, year-on-year, do not perfectly offset changes in the quantity of actual research.
+Second, we must assume that the number of publications associated with each clinical trial is constant across our period of interest (i.e., that the amount of research “embedded” in each publication is not systematically changing over time).
+We view this assumption as reasonable for three reasons.
+First, by 2010, nearly all scientific journals in our sample had adopted rules requiring authors of clinical trial publications to adhere to a standardized format when reporting trial results (Schulz et al.,, 2010).
+Under these rules, the structure of a paper and its contents are pre-determined; there is little scope for authors to, for example, split results across several publications.
+Second, when constructing our sample of clinical trials, we have aimed to exclude those publications that re-analyze data generated by an existing study.
+Third, more suggestively, we use data on clinical trial publications supported by National Institutes of Health grants to examine the quantity of grant dollars per reported clinical trial publication over this period.
+We find no evidence of a change in this relationship between funding and publication during our period of interest.
+[^48] We inspect this relationship by collecting records from the NIH RePORTER database, for years 2010 to 2019.
+Specifically, we collect records of the total award associated with funded projects and the quantity of any publications flagged by our procedure as clinical trials.
+For the median NIH grant, we find that an award of roughly \$100,000 per clinical trial, and a mean of roughly \$200,000 per trial.
+This is stable across years of data.
+4.3. Why is the Quantity of Research Rising?
+The quantity of papers citing clinical trials has increased, per Figure 4.1, by nearly a factor of two over our period of interest.
+In this section, we examine various measures of publication composition and quality to shed light on the source of this increase.
+Our aim, in doing so, is to determine whether this measured increase corresponds to a “real” increase in the quantity of underlying research or whether it captures changes in the ease and frequency of publication.
+Evidence consistent with the latter suggests, then, that counts of scientific publications may provide little insight on questions concerning research productivity.
+4.3.1. Publication Content
+We inspect the technical content of “non-clinical-trial” papers in three ways.
+We begin by hand-labeling random samples of 100 records, drawn from the beginning (publication year 2013) and end (publication year 2019) of our sample of papers that cite clinical trials.
+In this pooled set of 200 records, we identify one record that should have been in our sample of clinical trials.
+[^49] The erroneously labelled publication is PMID 23797691, which reports the results of a clinical trial studying a drug to treat HIV.
+Note that this exercise suggests a false negative rate of 0.005, consistent with estimates reported previously.
+The remainder of the records can be split into seven sets: editorial records (e.g., comments on other papers, errata, editorials, 2.5%), publications reporting preclinical studies (e.g., in-laboratory and in-animal tests, 21.5%), clinical trial protocols (1%), randomized trials studying something other than a drug (2.5%), case studies and case reports (4.5%), observational studies (30.5%), and review articles (e.g., meta-analyses and literature reviews, 37%).
+[^50] In three cases, review articles appeared with the heading “Comment” or “Editorial.” We categorize these records as falling into the “Review” category.
+We find a 30 percent increase in the total number of review articles between 2013 and 2019, while the quantity of observational studies remains essentially constant (31 in 2013, and 30 in 2019).
+The increase in the quantity of meta-analyses is consistent with increases documented elsewhere (Ioannidis et al.,, 2013).
+Of course, in our small sample of records, it is difficult to draw authoritative conclusions about the role of growth in this type of literature.
+We collect measures of changes in the quantity of meta-analyses, based on the hypothesis that this may be a quantitatively important driver of the changes in our sample, in two ways.
+First, we implement a simple search of the abstracts of publications in this “citing” sample.
+We inspect the set of hand-labeled records, used as the “training data” for our fine-tuning procedure in Section 3.3, as many were flagged as out-of-sample because they were literature reviews or meta-analyses.
+From this set, we collect a list of keywords that were indicators when we hand-labeled these records that the publication was a meta-analysis.
+Our search terms are intentionally specific, as the aim of this procedure is to yield a plausible lower bound for any trend.
+[^51] We searched for the following keywords: meta-analysis, metaanalysis, metaanalyses, systematic review, systematic reviews, systematically review, systematic search, review of published data, literature review, literature search, search of databases, review all literature, reviewed all literature, narrative review, systemic review.
+We also included any record that mentioned at least two of the following databases, which—based on our hand-labeling experience—are often referenced as part of the methods involved in a literature review or meta-analysis: MEDLINE, EMBASE, CINAHL, PubMed, Cochrane Central Register of Controlled Trials, BioMedCentral.
+More liberal search terms (e.g., “review”) risk capturing a large number of extraneous records.
+As expected, this yields a small number of records: in our three-year citing trials sample, there are 1,606 record in our citing sample that satisfy these criteria in 2013 and 3,372 in 2019, corresponding to a 109 percent increase.
+Our second method generates—as the discussion in Section 3 suggests—what might be viewed as an upper bound on the quantity of meta-analyses and literature reviews in our data.
+We collect all records indexed in PubMed with the NLM tags “meta-analysis,” “systematic review” and “review.” There are 9,407 records in our citing sample with one of these tags in 2013 and 14,991 in 2019, corresponding to a 59 percent increase.
+[^52] Figure D.8 displays a time series of the number of ‘citing’ papers that satisfy both our keyword and NLM-tag based methods for identifying meta-analyses.
+From these three procedures, we draw a suggestive conclusion: there has been a quantitatively important increase in the number of meta-analyses over our period of interest.
+This increase is on the order of 30 to 100 percent, but various measurement strategies yield alternative estimates.
+The specific quantity of meta-analyses is not of central interest for this paper, except insofar as it allows us to conclude that the quantity of medical research that summarizes existing papers—rather than producing new evidence—appears to be growing over time.
+This trend provides support for our decision to focus on clinical trials as a measure of productivity-relevant research and sheds some light on the source of documented increases in the quantity of research in other papers.
+As we note in Section 2, meta-analyses and similar forms of literature review can provide important information to physicians and can be vital in the diffusion of existing information, especially as the stock of knowledge grows (Jones,, 2009).
+Here, then, the distinction between those forms of scientific effort that are relevant to productivity metrics and those that are relevant to more expansive welfare metrics is especially important.
+4.3.2. Publication Geography
+Changes in the production of meta-analyses have been linked (e.g., Ioannidis et al.,, 2013) to shifts in the geographic distribution of medical research.
+Publication data allow us to inspect geographic trends in our sample.
+Figure 4.4 displays trends in the annual quantity of published clinical trials, and papers citing clinical trials, disaggregated by the location of the first-listed author.
 [^53] In most domains of scientific and medical publication, last-listed authors are senior investigators.
@@
 In our sample, the top four producers of published clinical trials, and papers citing clinical trials, are, in order, the United States, China, Germany, and Japan.
-Clinical trial investment—across geographies—is stable.
-[^31] Note that Panel A of Figure 6 displays measurements of the location of clinical trial publication authors, not the location of clinical trial sites.
+Figure 4.4. Composition of Medical Research Across Countries
+Panel A: Clinical Trials Panel B: Papers Citing Clinical Trials
+Refer to caption Refer to caption
+Notes: Figure 4.4 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, by country.
+The sample of published clinical trials is constructed with the conservative model.
+To address truncation, we report the number of publications that cite clinical trials published in the preceding three years.
+Clinical trial quantities—across geographies—are stable over our period of interest.
+[^54] Note that Panel A of Figure 4.4 displays measurements of the location of clinical trial publication authors, not the location of clinical trial sites.
 We cannot rule out that, for example, trials authored by researchers with United States mailing addresses were conducted elsewhere.
@@
 This small change is dwarfed by increases—on the order of 30-50 percent—in the quantity of other medical research from China and from countries outside of the top four producers of medical research.
-Figure 6. Composition of Medical Research Across Countries
-Panel A: Clinical Trials Panel B: Papers Citing Clinical Trials
-Refer to caption Refer to caption
-Notes: Figure 6 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, by country.
-The sample of published clinical trials is constructed with the conservative model.
-To address truncation, we report the number of publications that cite clinical trials published in the preceding three years.
-These compositional facts are, in many ways, unsurprising.
-In recent years, editorials in prominent journals in medicine have observed that changes in the incentive structures for physicians, scientists, and even medical students, across countries, have led to “an extreme of quantity at the expense of quality” in medical publishing (Siegel et al.,, 2018).
-[^32] See, as one example, program certification requirements, which describe publication expectations for medical students and clinical faculty, from the Accreditation Council for Graduate Medical Education in Accreditation Council for Graduate Medical Education, (2017).
-Recent work documents a substantial increase in the quantity of meta-analyses and systematic reviews—summaries of existing clinical trials—driven by scientists from outside of the United States (Ioannidis, 2016a, ).
-In parallel, there has been a rapid increase in the number, and scale, of “mega-journals”—scientific journals published frequently, with a large and growing number of pages (Ioannidis et al.,, 2023).
-Nonetheless, substantial constraints—financial, ethical, and practical—keep firms, physicians, and scientists from increasing the frequency of clinical trials.
-Clinical trials are infrastructure-intensive.
-Trials are run at specific sites, which must identify patients who satisfy all relevant eligibility criteria and monitor them for periods ranging from six months to twenty years (Piantadosi,, 2024).
-Best available estimates suggest that trials can cost on the order of 500 million to one billion dollars to complete (Qiao et al.,, 2019).
-Perhaps most importantly, sites must be capable of producing data sufficient to persuade regulators.
+Institutional details are helpful in making sense of the differential patterns across these cuts of data.
+Regulatory requirements and local capabilities are key factors influencing where clinical trials are conducted and published.
+Sites must be capable of producing data sufficient to persuade regulators in high-value markets.
 Certain regulators require that sponsors of new drug applications submit trial evidence collected from a domestic population.
@@
 See Namba et al., (2024) for one discussion.
-In recent years, however, the U.S.
-Food and Drug Administration (FDA)—responding to concerns from patient groups—has indicated a limited willingness to approve drugs on the basis of evidence collected from exclusively foreign sites (for a longer discussion, see Alsan et al.,, 2024).
+In recent years, the U.S.
+Food and Drug Administration (FDA)—responding to concerns from patient groups—has indicated that it may not be willing to approve drugs on the basis of evidence collected from exclusively foreign sites (for a longer discussion, see Alsan et al.,, 2024).
 In one especially high-profile example, a cancer drug was rejected by the FDA after being tested exclusively in China (Kolata,, 2022).
 As long as the U.S.
-market continues to have outsized value for pharmaceutical firms—a consequence of the especially high prices paid by American consumers—it may be unsurprising to find that evidence disproportionately originates in the United States.
+market continues to have outsized value for pharmaceutical firms—a consequence of the especially high prices paid by American consumers—it may be unsurprising to find that trial evidence disproportionately originates in the United States, even as global production of other forms of medical research continues to grow.
 Observational studies, meta-analyses, and case reports are, by contrast, cheap and not subject to the same financial and regulatory pressures.
-Digital tools, including search engines and language models, will likely further depress these costs.
+Digital tools, including search engines and the large language models deployed in this paper, will likely further depress these costs, potentially further widening this gap.
+Four countries are the most prolific producers of clinical trials and other forms of medical research.
+Of interest here, however, are those countries that experienced the largest growth in their production of medical research over this period.
+On this measure, China, Poland, Japan, and Spain top the ranking in our data.
+In Section D.2, we provide more details on the set of countries with the largest growth in this form of medical research over our period of interest.
+Several details are worth noting.
+First, Chinese production of records in our citing sample increases by roughly 225 percent, producing 5,620 papers in our citing set in 2019.
+The next country in this ranking, Poland, experiences 96 percent growth and produces 311 papers in our citing set in 2019.
+The top 23 countries in terms of growth, listed in the appendix, produce 93 percent of all 2019 papers in this set.
+Per the World Bank income classification scheme, nearly every country in this fast-growing group is high income.
+The only countries categorized differently are China, India, Brazil, Iran, and Turkey, which the World Bank classifies as either upper (China, Brazil, Iran, Turkey) or lower (India) middle-income.
+[^56] https://datatopics.worldbank.org/world-development-indicators/the-world-by-income-and-region.html That is, this trend captures, in part, an increase in the quantity of diffusion-focused literature in high-income countries that have long had robust health infrastructure.
+Below, we establish that on several quality measures, these publications appear to be declining in value over time—consistent with the idea that some of this increase in production may be a response to changing scientific incentives for publication, rather than “real” invention or diffusion effort.
+4.3.3. Publication Quality
+Next, we collect a set of facts about the quality of this growing set of “citing” papers.
+We begin in Figure 4.5 by replicating Figure 4.3 for the sample of papers that cite clinical trials.
+There are several elements of this plot that are informative.
+First, the share of “citing” papers with a small number of citations appears to grow over time.
+That is—the left tail of the citation distribution for this set is both long and growing.
+Note also that the right tail of the citation distribution appears to shrink over time.
+That is, the proportion of highly cited papers appears to be shrinking.
+These patterns provide one indication that the central tendency of the quality of non-clinical-trial medical research may be decreasing over time.
+[^57] Figure D.5 gives an analogous figure, displaying the distribution in citations from all publications to the sample of publications that cite clinical trials.
+Here, the right tail of the distribution of citations is growing.
+Figure 4.5. Citation Distribution Across Time, Papers that Cite Clinical Trials
+Notes: Figure 4.5 displays a heat map measuring the quantiles of the distribution of citations received by papers that cite clinical trials, in each calendar year, from leading journals.
+The $y$-axis has been stretched to elongate the right-tail of the citation distribution.
+Colors are displayed in a log scale.
+The sample of published clinical trials is constructed with the conservative model.
+The sample of papers that cite clinical trials is constructed using a three-year window.
+Citations are imperfect measures of the ex post quality of a scientific publication.
+Our data allow us to examine a variety of other measures that provide insight into the ex ante quality of a particular instance of research.
+Unfortunately, shifting geographic trends in this sample and the low quality of measures of public funding for countries outside of the United States render it difficult to use a measure of public funding as a proxy for quality.
+We do find a 10 percentage point decrease in the share of publications in this set that record having received (U.S.-based) public funding, but this is likely best interpreted as a mechanical consequence of the geographic heterogeneity displayed in Figure 4.4.
+Instead, we focus on author characteristics, constructed using publication data, that also shed some light on research quality.
+Scientific author norms typically mean that first-listed authors are junior scientists and last-listed authors are senior, lead scientists.
+We collect facts about these groups separately.
+We begin by examining the productivity of senior scientists who authored papers in our citing sample.
+We collect all publications indexed in PubMed authored by lead scientists (last-authors) who write at least one paper in our citing sample.
+The median scientist in this set writes five publications per year.
+We observe opposing trends in annual publication quantity and citation patterns during this time.
+We find that the median quantity of papers written by a scientist per year increases from 3 to 5 over this period.
+The median number of citations to those papers from leading journals drops, over the same period, from 3 to 1.
+At the 90th percentile, the quantity of per scientist publications per year is stable, at 9 per scientist between 2013 and 2019.
+Citations from leading journals are similarly stable, at 10 per publication.
+[^58] We observe sharp declines in the likelihood that a scientist has ever received funding from the NIH or from another U.S.
+federal agency.
+As noted above, however, we hesitate to place much weight on this fact, as it likely reflects changes in the composition of the sample.
+For the median senior scientist who appears in our citing sample in 2013, approximately 80 percent of their previously published papers (i.e., published before 2013) had received at least one citation.
+For scientists who first appear in our citing sample in 2019, the share of previous publications with a non-zero number of citations falls to roughly 70 percent.
+For junior scientists, we focus on a measure of attrition: after authoring a paper in our citing sample, do we ever observe another publication written by that junior scientist?
+We collect all publication records in the Web of Science associated with these junior (first-listed) authors.
+In 2013, the likelihood that a junior scientist published another paper in 2014 was roughly 15 percent.
+By 2019, this falls to roughly 8 percent.
+We observe similar trends when two- and three-year publication patterns are observed.
+We interpret this attrition as evidence that junior scientists who author papers in our citing sample may be less likely to remain research productive in the future.
+Taken together, we interpret these facts about scientist research productivity as suggestive evidence of declining average research quality in our citing sample over time.
+When interpreted in light of the citation distribution trends in Figure 4.5, we conclude that measured increases in the quantity of this “other” medical research may be accompanied by a decline in its average quality.
+It is precisely this shift that suggests that quantities of scientific publications in this sector, when not restricted to more precise categories of research such as clinical trials, may be a poor indicator of research productivity.
+5. Discussion
+Studies of innovation and productivity have long grappled with the prospect of diminishing returns to research (e.g., Merton,, 1935; Stafford,, 1952; Schmookler,, 1952, 1954, 1966; Griliches,, 1990, 1994; Bloom et al.,, 2020).
+Answers to the question at the center of this literature—is the cost of sustained technological progress increasing over time?—have been unsatisfying, as fundamental challenges in the measurement of technological innovation have limited our ability to do more than determine that available metrics do not yield conclusive answers.
+In one sense, this paper fits in this tradition, in that we find that, in the context of medical research, recent concerns about declining productivity are an artifact of long-standing measurement challenges.
+On the other hand, enabled by advances in generative AI, we introduce a new technique for annotating unstructured text that allows us to construct a proxy for medical research that, we argue, closely approximates the object of interest in an analysis of medical research productivity.
+We fine-tune an open-source large language model, which allows us to extract records of clinical trials from a corpus of scientific publications.
+The resulting census corrects for errors in previous measures of clinical trial research and approximates the quality of hand-labeled data.
+We show that trends in the quantity, quality, and composition of clinical trials have been stable since 2010.
+Our data allow us to contrast these trends with analogous patterns in the production of other forms of medical research.
+We find that, since 2010, the quantity of other forms of medical research has roughly doubled, a pattern consistent with estimates reported in other papers.
+On inspection, the increase in quantity includes a large increase in the number of papers authored by scientists in China and in the number of papers that synthesize existing literature (e.g., literature reviews and meta-analyses).
+Growth in these forms of research coincides with an average decline in various measures of publication quality.
+Taken together, we interpret these facts as providing no support for claims of declining research productivity, at least since 2010.
+These patterns are, in many ways, unsurprising.
+In recent years, editorials in prominent medical journals (Harvey,, 2020) have observed that changes in the incentive structures for physicians, scientists, and even medical students, across countries, have led to “an extreme of quantity at the expense of quality” in medical publishing (Siegel et al.,, 2018).
+[^59] See, as one example of how these incentives have shifted, program certification requirements from the Accreditation Council for Graduate Medical Education in Accreditation Council for Graduate Medical Education, (2017).
+The ACGME provides formal accreditation for training programs and, as part of this process, lays out expectations of minimum quantities of publication for medical students and clinical faculty.
+Note that certain researchers have put forth alternative explanations for the rapid increase in the total number of publications.
+A trend toward evidence-based decision-making in medicine has increased demand for articles that synthesize large bodies of research (Lohr,, 2004).
+Recent work documents a substantial increase in the quantity of meta-analyses and systematic reviews—summaries of existing clinical trials—driven by scientists outside of the United States (Ioannidis, 2016a, ).
+In parallel, there has been a rapid increase in the number, and scale, of “mega-journals”—scientific journals published frequently, with a large and growing number of pages (Ioannidis et al.,, 2023).
+Despite these changing incentives, substantial constraints—financial, ethical, and practical—keep firms, physicians, and scientists from increasing the quantity of clinical trials.
+Clinical trials are infrastructure-intensive.
+Trials are run at specific sites, which must identify patients who satisfy all relevant eligibility criteria and monitor them for periods ranging from six months to twenty years (Piantadosi,, 2024).
+Sites may struggle to find patients who meet these criteria, especially in small disease markets (Kolata,, 2017), and the costs of accruing large samples for certain studies can be prohibitive (Alsan et al.,, 2024).
 Existing work documents patterns that are not inconsistent with those in this paper.
-However, we offer an alternative interpretation.
+We differ, however, in interpretation.
 Like Park et al., (2023) and Bloom et al., (2020), we find a substantial increase in the size of the medical research ecosystem.
@@
 Technically, we find the same.
-As increases in non-clinical trials outpace (non-existent) changes in clinical trials, the average productivity of studies in Figure 3 mechanically decreases.
-In debates on productivity, however, we argue that this measure of central tendency is not the policy-relevant object.
-To be sure, these findings raise a host of other important questions—about misaligned incentives, wasteful spending, and the productive efficiency of science.
-But canonical frameworks, such as Romer, (1990), center on the simple idea that it is high-intensity investments that generate meaningful shifts in outcomes.
-An explosion of low-capital research that has a negligible impact on outcomes, then, is not cause for concern from a growth perspective.
-That is, growth-oriented studies may prefer to focus—like much of the research on science and innovation—on the right-tail of the productivity distribution (e.g., Kelly et al.,, 2021).
-Our findings do not speak to shifts in upstream, basic science inputs into clinical trials, or to shifts in factor prices.
-In our view, there is little credible evidence indicating changes on either dimension during this time period.
-This setting allows us, only, to note that any shift in inputs or factor prices must be perfectly offset by a countervailing force, to account for the complete stability of clinical trial counts, citations, journal placements, and reliance on public funding in our data.
-Our findings, also, apply to roughly one decade: 2010 to 2022.
-We impose this restriction because policy changes in the early 2000s substantially shifted incentives for both publication and standardized formatting of clinical trial results.
-See Laine et al., (2007) and Schulz et al., (2010) for details.
-Given our methods, these changes mean that data construction over a longer period risks decreasing the quality of our data.
+As increases in the quantity of non-trial publications outpace (non-existent) changes in clinical trials, the average productivity of studies in Figure 3 mechanically decreases.
+We find, also, that increases in the quantity of non-trials are coincident with declines in various measures of publication quality.
+From the perspective of a productivity calculation, however, we argue that measures of average quality provide little insight.
+We cannot rule out, for example, that the same amount of research is being produced in these non-trial publications, but that what had previously been reported in a single paper is now spread across several distinct records.
+To be sure, these findings raise a host of other important questions about scientific research—about misaligned incentives, wasteful spending, and both productive and allocative efficiency.
+But on the issue of research productivity, these shifts provide very little insight.
+Our findings do not speak to changes in upstream, basic science inputs into clinical trials, or to changes in factor prices.
+As Section 3 discusses, in our view, there is very little credible evidence that might allow us to investigate changes on either dimension during this time period.
+We can note, only, that any shift in inputs or factor prices must be perfectly offset by a countervailing force, to account for the complete stability of clinical trial counts, citations, journal placements, and reliance on public funding in our data.
+Our findings, also, apply to roughly one decade.
+We impose this restriction because a set of policy changes in the early 2000s substantially shifted incentives for disclosure and publication of clinical trial results (Laine et al.,, 2007; Schulz et al.,, 2010).
+To be clear, there is nothing about our method that limits its applicability to data reported in the particular form required by these policy changes.
+Our concern here is about the interpretability of the resulting series.
+As noted previously, proxies for scientific research provide very little insight if the relationship between one unit of the proxy and units of underlying, unobserved research shift over time.
+Policy changes in this sector are such that we risk introducing this source of bias with a longer series.
 Thus, we cannot speak to trends that precede this period.
-That clinical trial investment is remarkably stable in the past decade, however, is likely the relevant consideration for contemporary public policy.
-What our findings, and the simple framework borrowed from the growth literature, do suggest is that there is likely value in efforts to increase the productivity and decrease the costs of clinical trials.
+From the perspective of contemporary public policy, however, trends from 2010 to the present are likely to be most relevant.
+We reconcile two stylized facts about research and inventive output in a manner that is both optimistic and pessimistic.
+On the one hand, we find that, in the most recent period, both the inputs and outputs of an “idea production function” (2.1) are constant across periods.
+This should alleviate the concern, in this sector, that “ideas are getting harder to find” (Bloom et al.,, 2020).
+On the other hand, it is worth remarking, briefly, on the implications of completely stable clinical trial trends.
+The early twenty-first century is often described by scientists as a “golden age” for medicine.
+Breakthroughs, such as the completion of the Human Genome Project and the diffusion of DNA sequencing technologies one billion times faster than their predecessors, rapidly expanded the set of available scientific opportunities (e.g., Scannell et al.,, 2012).
+Technological advances were matched by institutional support, including the doubling of the budget of the U.S.
+National Institutes of Health between 1998 and 2003.
+Against these factors that seemingly forecast a period of rapid innovation, the two facts documented in this paper—stability in productivity-relevant research investments, but rising quantities of comparatively low-impact forms of research—highlight the importance of understanding frictions that affect the translation of basic scientific insights into medical technologies relevant to patients.
+Specifically, these patterns suggest that there is likely value in efforts to increase the productivity and decrease the costs of clinical trials, and in efforts to reconcile incentives in academic publishing with social value.
 A long-standing discussion in medicine and statistics, spurred by Altman, (1994), centers on the argument that “we need less research, better research, and research done for the right reasons.” Our data indicate that nearly half of all clinical trials published in our sample are never cited by leading medical journals, and roughly 15 percent are never cited at all.
+Outside of clinical trials, our findings suggest that the quantity of low-quality research might be growing over time.
 This heterogeneity is consistent with findings in Ioannidis, 2016b and van Zwet et al., (2023), who imply that there may exist trade-offs between quantity and data quality in this setting.
 If increasing the scale of research infrastructure is a policy objective, such efforts should invest in tools, like patient registries, that enable the production of high-quality, representative information (Alsan et al.,, 2024; Durvasula,, 2023).
-Supplemental Appendix to: Medical Research as a Productivity Indicator [^1] Date: May 12, 2024
+Supplemental Appendix to: Medical Research as a Productivity Indicator [^1] Date: August 3, 2024
 $\star$ Stanford Department of Economics and Stanford Law School, maya.durvasula@stanford.edu
@@
 Maya M. Durvasula Sabri Eyuboglu David M. Ritzwoller
-Contents  appendix.Asubsection.A.1subsection.A.2subsubsection.A.2.1subsubsection.A.2.2subsubsection.A.2.3subsubsection.A.2.4subsubsection.A.2.5subsection.A.3subsection.A.4appendix.Bsubsection.B.1subsection.B.2subsection.B.3subsection.B.4subsection.B.5subsection.B.6appendix.Cappendix.D
+Contents  appendix.Asubsection.A.1subsection.A.2subsubsection.A.2.1subsubsection.A.2.2subsubsection.A.2.3subsubsection.A.2.4subsubsection.A.2.5subsection.A.3subsection.A.4appendix.Bappendix.Csubsection.C.1subsection.C.2subsection.C.3subsection.C.4subsection.C.5subsection.C.6appendix.Dsubsection.D.1subsection.D.2subsection.D.3subsubsection.D.3.1subsubsection.D.3.2subsubsection.D.3.3appendix.E
 Appendix A Publication Data
@@
 We give a general overview of the PubMed data in Section A.1.
-In Section A.2, we describe the process of identifying our initial sample of abstracts, i.e., the “Potential Universe” displayed in Figure 4.
+In Section A.2, we describe the process of identifying our initial sample of abstracts, i.e., the “Potential Universe” displayed in Figure 4.1.
 We discuss the process of determining the set of leading medical journals in Section A.3.
@@
 Although, in principle, registry identifiers may be reported in publication full-texts, rather than abstracts, the low frequency of identifiers in our sample suggests non-compliance with requirements, consistent with patterns in ClinicalTrials.gov documented in DeVito et al., (2020).
-Second, Table A.1 sheds light on the differences in trial composition that we observe in Figure 1: NLM tags that, ostensibly, capture clinical trials have little overlap with one another and do not fully capture the records in our sample.
+Second, Table A.1 sheds light on the differences in trial composition that we observe in Figure 3.1: NLM tags that, ostensibly, capture clinical trials have little overlap with one another and do not fully capture the records in our sample.
 Of the roughly 150,000 records in our census, approximately 100,000 could be identified using NLM tags for “clinical trial” (and its variants) and “randomized controlled trial.” Observe, however, that use of such flags captures a much larger number of records than one would intend to include.
@@
 A.3. Determining the Set of Leading Medical Journals
-Angrist et al., (2020) propose a strategy to identify leading journals in a field, which we adapt to this context.
+Angrist et al., (2020) propose a strategy to identify the leading journals in a field, which we adapt to this context.
 In medicine, the “trunk journals,” per Angrist et al., (2020), are the Journal of the American Medical Association and the New England Journal of Medicine.
-We collect all citations originating from these journals to records in PubMed between 2010 and 2022.
-Angrist et al., (2020) collect initial journal lists for their citation exercise of the journal is one of the top fifty most cited by a trunk journal.
-As medicine includes a large number of subfields, we instead collect a list of journals that receive at least 100 citations from a trunk journal over this time period, which yields a list of 84 journals.
+We collect all instances in which papers published in these journals cite other records in PubMed, between 2010 and 2022.
+In Angrist et al., (2020), the authors select the fifty journals receiving the most cites from these trunk journals.
+As medicine contains a large number of subfields, we instead collect a list of journals that receive at least 100 citations from a trunk journal over this time period, which yields a list of 84 journals.
+In our data, papers published in trunk journals cite records in 3,136 unique journals.
+[^61] We use unique entries in the field medlineta—which correspond to distinct MEDLINE abbreviations for journal names—to identify distinct journals.
+See https://www.nlm.nih.gov/bsd/serfile_addedinfo.html for details on name abbreviations in this database.
+We confirm that the count of journals is identical if we, instead, use journals identified by the National Library of Medicine’s unique journal identification number.
+The median journal is cited twice by these two trunk journals over our time period.
+The most frequently cited journals are, perhaps unsurprisingly, the trunk journals themselves: JAMA receives 3,143 citations, and the New England Journal of Medicine receives 4,302 citations.
+The next ten journals, by frequency of citations, are: Lancet (1,457), Circulation (789), Journal of Clinical Oncology (734), Nature (660), Blood (651), PLoS One (645), Journal of the American College of Cardiology (593), Clinical Infectious Diseases (589), BMJ (586), and Annals of Internal Medicine (586).
 A.4. The Web of Science
@@
 Thus, we impute to each record the country associated with the first-listed author.
-In Figure 6, we plot geographic trends over time.
-Appendix B Prompt Design, Fine-Tuning, and Performance Assessment
+In Figure 4.4, we plot geographic trends over time.
+We use the Web of Science to collect details about author career and publication histories, which allow us to investigate patterns in the production of medical research.
+The Web of Science assigns to authors a unique, persistent identifier (DAIS ID), which allows us to collect information on authors’ output before and after the publication focal records in our sample.
+To our knowledge, there have been no large-scale efforts to validate the quality of these researcher identifiers.
+On inspection, it appears as though the identifiers perform best, perhaps unsurprisingly, for researchers with distinctive names.
+When researchers have names that are more frequent in the population, our spot-checking (against online publication records) suggests that there are more likely to be errors.
+Nonetheless, tracking researchers using these identifiers is standard in this literature, as there are no obviously superior alternatives.
+Especially given that researcher productivity over time is not a centerpiece of our analysis, we do not undertake additional validation exercises.
+Appendix B ClinicalTrials.gov
+ClinicalTrials.gov is a registry of clinical trials, run by the United States National Library of Medicine.
+As of 1 August 2024, ClinicalTrials.gov lists more than 500,000 studies in all 50 states and in 222 countries.
+[^63] See https://clinicaltrials.gov/about-site/trends-charts.
+ClinicalTrials.gov was available for study registration beginning 29 February 2000.
+At this time, there were 1,255 registered studies.
+[^64] In 1988, in response to public pressure surrounding perceived lags in the development of therapeutics and preventives to treat HIV/AIDS, U.S.
+Congress passed the Health Omnibus Programs Extension Act (Public Law 100-607), which required the creation of a database of AIDS Clinical Trials Information Services (ACTIS).
+Publicly-funded clinical trials that were associated with these previous data registration efforts were included in ClinicalTrials.gov at its inception.
+By the beginning of 2006, there were 24,822 studies and by the beginning of 2012, 118,020.
+In January 2024, there were 477,227.
+The geographic composition of these studies has shifted considerably over time, as have reporting patterns from various firms and universities.
+In this Appendix, we use data from ClinicalTrials.gov to illustrate its limitations as a source for the type of high-quality records necessary for studies of productivity.
+Given the parallels between this paper and a long literature that studies the usefulness of patent data for studies of innovation, it is worth highlighting the important, perhaps obvious, distinction between ClinicalTrials.gov as a government database and records collected by the U.S.
+Patent and Trademark Office (USPTO).
+Records of U.S.
+patents are collected by the USPTO itself.
+Disclosure is widely considered a pivotal component of the patent system: in exchange for a short-term quasi-monopoly, information on patented inventions is placed in the public domain.
+Although ClinicalTrials.gov is described in similarly forceful terms in the laws that authorize its creation, disclosure has been—and continues to remain—effectively voluntary, in the sense that legal tools have never been used by a federal agency to compel the registration of a study or disclosure of its findings.
+Institutional details suggest that the “voluntariness” of this reporting has shifted over time, in ways that likely affected sponsors’ propensity to register.
+[^65] In guidance to researchers on the use of ClinicalTrials.gov, Tse et al., (2018) collect a list of the ten challenges associated with using these data, chief among which is the fact that changes in medical journal reporting requirements and federal law make it extremely difficult to determine if increases in the size of the database are related to actual increases in the amount of scientific research.
+A coalition of medical journal editors began encouraging registration in 2005 (Laine et al.,, 2007), registration requirements were imposed for certain FDA-regulated trials in 2008, and the FDA began threatening to impose fines on noncompliant sponsors in 2021.
+In Table B.1, we collect a list of five changes in the past 25 years that changed sponsors’ requirements for trial registration.
+Table B.1. Policy Changes Affecting the Composition of ClinicalTrial.gov
+Policy (Effective Date) Registration Requirement
+Food and Drugs Administration Modernization Act (21 November 1997; ClinicalTrials.gov registry available starting on 29 February 2000) Clinical trials of investigational new drugs for serious or life threatening conditions and diseases
+International Committee of Medical Journal Editors (1 July 2005 for newly initiated trials and 13 September 2005 for ongoing trials) All clinical trials as a condition for consideration of publication of results
+Food and Drugs Administration Amendments Act (FDAAA) (27 September 2007) Non-phase 1 clinical trials of FDA regulated drug and biological products, and non-feasibility trials of FDA regulated device products
+Final rule implementing FDAAA in Title 42, Part 11 of the Code of Federal Regulations (42 CFR Part 11) (18 January 2017) Non-phase 1 clinical trials of FDA regulated drug and biological products, and non-feasibility trials of FDA regulated device products
+Notes: Table B.1 collects a list of five policy changes in the past 25 years that changed sponsors’ requirements for trial registration. This table is adapted from Tse et al., (2018).
+We inspect a version of the Aggregate Analysis of ClinicalTrials.gov (AACT) data, a publicly available relational database that contains information on all studies registered on ClinicalTrials.gov, downloaded on 2023 May 31.
+Of interest, for us, is the question of whether ClinicalTrials.gov can provide an accurate count of the total quantity of clinical trials and, if so, at what date such an accurate count might be available.
+Figure B.1 displays the results of our investigation.
+Figure B.1. Number of Entries in Clinical Trials.gov, by Sample Restrictions
+Notes: Figure D.1 displays the counts of the number of records in ClinicalTrials.gov that satisfy various sample restrictions.
+The $x$-axis of the top panel is the year in which the clinical trial was completed.
+The $x$-axis of the bottom panel is the year in which the clinical trial was registered on ClinicalTrials.gov.
+We plot counts of records, selected in various ways, against two x-axes.
+In the top panel of Figure B.1, we plot counts associated with the year in which clinical trials are marked complete.
+[^66] These completion dates are somewhat imprecise.
+If a trial is registered, but no results are reported, the completion date is an estimate, based on the projected completion date in the study’s protocol.
+If the trial is registered and has associated results, the completion date is the actual date on which data collection ended.
+In the bottom panel, we plot counts based on the year in which the study was posted on ClinicalTrials.gov.
+Measures of trial completion are informative, insofar as they capture the date on which information production is finished.
+However, trials vary greatly in length—the key observation in Budish et al., (2015)—and as a result represent investments “sunk” at very different earlier points in time.
+Measures of trial posting are analogous, in some ways, to publication, in that they capture decisions to disclose information.
+Even without considering the specific series in Figure B.1, observe that the bottom panel includes a variety of sharp changes in the count of records.
+These correspond to several of the policy changes in Table B.1.
+We see a sharp change in all four series in 2005, when medical journal policies were changed, a sharp change in 2008, when the FDAAA took effect, and a sharp (albeit smaller) change in 2016, in the lead-up to the effective dates of the final two policy changes in Table B.1.
+We observe a smaller, but discernible, change in the slope of each series in 2008 when measures of trial completion are used.
+We draw attention to these differences, in part, because an important challenge for the use of these records in research is determining the appropriate measure of time.
+Turning to the content of these series, we begin by examining the total number of records indexed in ClinicalTrials.gov over time (light green).
+We define a record as a unique “national clinical trial identifier,” although Tse et al., (2018) note that one study may be assigned multiple NCTs.
+There is no method, to our knowledge, that further distinguishes between records.
+In both panels, this series begins at zero in the year 2000 and increases sharply, to roughly 30,000 records by the end of our data, 2022.
+On inspection, it is clear that many of the 30,000 records included in the 2022 bin are not clinical trials.
+Among those that are clinical trials, many do not study medicines, but rather randomize exposure to interventions involving psychological treatments and exercise.
+We impose three sets of restrictions, which aim to bring the data closer to a consistent sample of clinical trials studying the effects of a drug.
+We select these restrictions based on examination of various fields in ClinicalTrials.gov and with the aim of providing an illustration of our concerns with the use of this dataset “off the shelf.” There are alternative ways of constructing a sample of records that may be necessary to yield relevant data for different research questions.
+We begin by dropping a set of records for which the study “status” variable in the database suggests that the study was either never started or was terminated.
+Specifically, we drop 40,622 records for which the status is listed as “Withdrawn,” “Suspended,” or “Terminated.” This status restriction is plotted in teal and has little impact on the trends in either panel.
+Next, we observe that not all records in ClinicalTrials.gov are interventional studies.
+We drop any study with a type other than “interventional.” This removes an additional 98,138 observations from our data.
+This status and type restriction is plotted in blue.
+We observe that the sharply increasing trend in the later years of our data flattens out slightly.
+That is—the growth in the quantity of records is driven, in part, by growth in the number of registered non-interventional studies.
+We implement a third restriction.
+We observe that clinical trial “phases” are a hallmark of drug trials: medical devices, for example, are approved via a “stage” process.
+Any record for which the trial phase is either missing or “Not Applicable” is then less likely to be a drug.
+We drop these records.
+This removes 151,541 observations from our data.
+We plot the resulting series in purple.
+When all three restrictions are imposed, we find steady increases in the quantity of trials leading up to 2008 and a stable series from 2008 forward.
+Reassuringly, this trend mirrors that documented in Figure 4.1.
+These plots suggest a level difference of roughly 2,000 observations, where Figure 4.1 is higher.
+This is consistent with the idea that ClinicalTrials.gov is not a global registry and, as Zarin et al., (2017) notes, many sponsors continue not to register studies in ClinicalTrials.gov
+We emphasize that it is not obvious—in the absence of corroborating evidence, similar to that in Figure 4.1—that the version of this series with all three restrictions in Figure B.1 provides an accurate measure of trends in research, especially given the sharp changes in the early part of the series and the somewhat haphazard way that we identified a set of sample restrictions.
+A series of recent papers—including DeVito et al., (2020) and DeVito and Goldacre, (2021)—document limitations with the quality of particular data elements, which further suggests the difficulty of using this database as a standalone source of estimates.
+Appendix C Prompt Design, Fine-Tuning, and Performance Assessment
 In this appendix, we detail the construction of the sample of clinical trials studied in Section 4.
-We begin in Section B.1 by describing the process we used to hand-label a sample of publications according to whether they satisfy the restrictions enumerated in 3.1.
-In Section B.2, we document poor performance of several approaches for classifying publications according to whether they belong in our sample based on standard machine learning methods.
-Section B.3 overviews the process of constructing prompts that perform well when used to query proprietary language models.
-Details associated with fine-tuning an open-source language model are given in Section B.4.
-Additional measurements of the performance of our model are given in Section B.5.
-B.1. Hand-Labeling
+We begin in Section C.1 by describing the process we used to hand-label a sample of publications according to whether they satisfy the restrictions enumerated in 3.1.
+In Section C.2, we document poor performance of several approaches for classifying publications according to whether they belong in our sample based on standard machine learning methods.
+Section C.3 overviews the process of constructing prompts that perform well when used to query proprietary language models.
+Details associated with fine-tuning an open-source language model are given in Section C.4.
+Additional measurements of the performance of our model are given in Section C.5.
+C.1. Hand-Labeling
 A standard concern when using large language models is the quality of the output.
@@
 We devise an interface for abstract labeling, which allows us to review records quickly.
-Figure B.1 provides an example of this interface.
+Figure C.1 provides an example of this interface.
 A labeller reads in a set of publication records and is shown one PubMed identifier (PMID) and one abstract at a time.
@@
 We do not add extra records to the testing dataset.
-Figure B.1. Abstract Labeling Interface
-Notes: Figure B.1 displays a screenshot of the application we developed to hand-label abstracts.
+Figure C.1. Abstract Labeling Interface
+Notes: Figure C.1 displays a screenshot of the application we developed to hand-label abstracts.
 Each abstract was labelled according to whether it satisfied the sample criteria enumerated in 3.1.
@@
 The user interface was developed with the “Meerkat” Python package available at the link: https://github.com/HazyResearch/meerkat.
-B.2. Benchmark Comparison to Standard Machine Learning Methods
+C.2. Benchmark Comparison to Standard Machine Learning Methods
 We measure the performance of a variety of standard machine learning algorithms for classifying publications according to whether they meet our sample restrictions.
@@
 A balanced loss is used to measure the performance of each hyper-parameter.
-Figure 2 displays receiver operating characteristics for the each class of models estimated with both types of embedding.
+Figure 3.2 displays receiver operating characteristics for the each class of models estimated with both types of embedding.
 Error rates are estimated in the testing data.
-The performance of the receiver operating characteristic for the “ensemble” model whose performance is displayed in Figure 3 is given in purple.
+The performance of the receiver operating characteristic for the “ensemble” model whose performance is displayed in Figure 3.3 is given in purple.
 The fine-tuned large language model developed in this paper substantially out-paces the performance of standard machine learning methods.
-B.3. Prompt Design and Error Analysis
+C.3. Prompt Design and Error Analysis
 We extract weak labels for 64,000 randomly selected abstracts with two proprietary large language models: OpenAI’s GPT-3.5 and GPT-4 (Nori et al.,, 2023; Bubeck et al.,, 2023).
-A priori, we expect GPT-4 to output labels of slightly lower quality than hand-labeled output and GPT-3.5 to produce slightly noisier labels than GPT-4.
-The extraction of GPT-3.5 labels is substantially faster and cheaper.
-To produce these labels at scale, each of the two models must be appropriately prompted.
-That is, each pre-trained model must be provided with a block of text as an input and asked to return the appropriate completion of this input.
-A priori, it is not clear what prompt structure will work well.
-The relative infancy of this area of research renders it difficult to identify a set of “best practices.” The proprietary nature of these models makes them difficult to study.
-See, for details, a guide to prompt engineering from OpenAI here: https://platform.openai.com/docs/guides/prompt-engineering.
-All of the prompts that we consider are displayed as figures in Appendix D.
-We identify three general prompt formats, which differ both in the amount of detail provided about our classification task and in the structure of the requested model completion.
-The simplest prompt provides a version of our sample definition, 3.1, and the text of an abstract, and asks the model to return ‘TRUE’ of the abstract satisfies these criteria.
-Otherwise, it returns false.
-The text of this prompt is displayed in Figure D.8.
-We refer to this prompt as Prompt 1.0.
-A second, more complicated prompt provides the same definition, with a set of examples of publication characteristics that do and do not satisfy the definition.
-The text of this prompt is displayed in Figure D.14.
-We refer to this prompt as Prompt 2.0.
-A third, more complex prompt provides the same definitions and examples, but asks the model to return an explanation of why the record does or does not satisfy these criteria.
-The text of this prompt is displayed in Figure D.17.
-We refer to this prompt as Prompt 3.0.
-We devise initial language for each prompt iteratively, using small numbers of records from our hand-labelled validation data.
+As discussed in the main text, we identify three general prompt formats, which differ both in the amount of detail provided about our classification task and in the structure of the requested model completion.
+We refer to these prompts as Prompt 1.0, Prompt 2.0, and Prompt 3.0, respectively.
+The text of these prompts are displayed in Figure E.1, Figure E.7, and Figure E.10, respectively.
 We test each of these three prompts in our 1000-record hand-labelled validation dataset, using both GPT-3.5 and GPT-4.
 We conduct a detailed error analysis, reported in LABEL:tab:Error-Analysis.
-For each instance in which a model returns a label (TRUE/FALSE) that differs from that in the hand-labelled dataset, we inspect the record.
-We categorize errors into types and subtypes.
 To make this concrete, suppose that GPT-3.5, given Prompt 1.0 and an abstract that reports the results of a meta-analysis, erroneously classifies the record as clinical trial, per our definition.
@@
 We devise error types based on our inclusion/exclusion criteria, in 3.1, and select sub-types based on common categories of errors.
-This exercise is instructive on three margins.
-First, it highlights the highest performing prompts.
-Second, it indicates particular types of errors in categorization—which suggest opportunities for more precise language in a prompt.
-Third, it draws attention to differences in the performance of the two models.
-Observe that, in the “Other” error type, we include a subtype called “overly literal interpretation of inclusion criteria.” We primary record errors of this type for Prompt 3, which asked the model to return a completion that described why, or why not, a record was classified as being in our sample.
-Here, GPT-3.5 makes 30 such errors—classifying records as FALSE by recapitulating the sample definition provided.
-GPT-4 makes two errors.
-Insight that emerges from this exercise, then, is that different prompt structures may be preferable depending on the model used.
 We revise each class of prompt based on these findings.
@@
 We refer to these revised prompts as Prompts 1.1, 1.2, and 1.3.
-The text of these prompts is displayed in Figures D.9, D.10, and D.12.
+The text of these prompts is displayed in Figures E.2, E.3, and E.5.
 For Prompts 2 and 3, we consider one variant each.
 We refer to these revised prompts as Prompts 2.1 and 3.1.
-The text of these prompts is displayed in Figures D.15 and D.18.
+The text of these prompts is displayed in Figures E.8 and E.11.
 These changes reflect the differences in performance catalogued in LABEL:tab:Error-Analysis.
-Table B.1 reports estimates of the true positive rate and false positive rate for each prompt in both models computed in the validation sample.
-Observe—for example, with Prompts 1.1, 1.2, and 1.3 queried to GPT-4—that even small changes in the text of a prompt yield substantial differences in performance.
-We attempt a second round of prompt iteration (Prompt 1.3).
-We observe that performance deteriorates with even small modification.
-Thus, we select the two highest performing prompts—one for each proprietary model.
-We use Prompt 2.0 to extract weak labels using GPT-3.5, and Prompt 1.2 to extract weak labels using GPT-4.
-These prompts are used to extract 64,000 labels—TRUE/FALSE flags—using GPT-3.5 and GPT-4.
-We refer to these as our set of “noisy labels.”
-Table B.1. Prompt Performance in Validation Data
-Panel A: GPT-3.5
-Prompt Type Prompt Sub-Type False Positive Rate True Positive Rate
-1 0 0.202 0.971
-1 0.065 0.949
-2 0.049 0.934
-3 0.056 0.912
-2 0 0.081 0.964
-1 0.072 0.964
-3 0 0.167 0.971
-1 0.068 0.956
-Panel B: GPT-4
-Prompt Type Prompt Sub-Type False Positive Rate True Positive Rate
-1 0 0.247 0.788
-1 0.172 0.898
-2 0.037 0.584
-3 0.037 0.489
-2 0 0.162 0.876
-1 0.176 0.905
-3 0 0.248 0.722
-1 0.171 0.883
-Notes: Table B.1 records the performance of each of eight prompt variants in a sample of 1,000 validation dataset records.
-Panel A reports performance associated with the proprietary model GPT-3.5.
-Panel B reports analogous statistics for model GPT-4.
-Prompt 1 asks the model to return TRUE or FALSE.
-Prompt 2 asks the model to return TRUE or the name of a specific, excluded category.
-Prompt 3 asks the model to return TRUE or an explanation of why the record should be excluded.
-Prompt sub-types correspond to various iterations.
-Sub-type 0 is the initial version of the prompt.
-Sub-types 1-3, where applicable, are subsequent iterations.
-Appendix D records the text of each prompt.
-We use Prompt 2, Sub-Type 0, to extract weak labels using GPT-3.5, and Prompt 2, Sub-Type 1 to extract weak labels using GPT-4.
-Table B.2. Baseline Prompt Error Analysis
+Table C.1. Baseline Prompt Error Analysis
 Model GPT-3 GPT-4
@@
 Notes: LABEL:tab:Error-Analysis categorizes errors documented in the first round of prompt iteration, using Prompts of sub-type 0.
-See Table B.1 for details on model performance.
+See Table 3.1 for details on model performance.
 For each model error—an instance in which a model returned a classification that deviated from hand-labelled data—we examined the associated record.
@@
 We report counts of errors, by model and prompt, of each Type and Sub-type.
-B.4. Fine-tuning
+C.4. Fine-tuning
 This appendix gives technical details related to fine-tuning open-source large language models with noisy labels extracted from GPT-3.5 and GPT-4.
@@
 All models were trained on a single $V100$ GPU with 16GB of HBM.
-B.5. Performance Assessment
+C.5. Performance Assessment
 In this appendix, we assess the performance of various fine-tuned language models for classifying whether a publication satisfies the restrictions enumerated in 3.1.
-Figure B.2 displays estimates of the receiver operating characteristic of several fine-tuned large language models.
+Figure C.2 displays estimates of the receiver operating characteristic of several fine-tuned large language models.
 Here, we consider only the performance of the RoBERTa base model.
@@
 The performance of each language model is measured in the testing dataset.
-Figure B.2. Receiver Operating Characteristic by Data Size and Origin
-Notes: Figure B.2 displays estimates of the the receiver operating characteristic of several fine-tuned language models.
+Figure C.2. Receiver Operating Characteristic by Data Size and Origin
+Notes: Figure C.2 displays estimates of the the receiver operating characteristic of several fine-tuned language models.
 The models differ in terms of the source of their training data.
@@
 There appears to be a threshold, where performance dramatically improves, at around 8,000 training labels.
-The models displayed in Figure 3 were each obtained by using all 64,000 training labels.
-B.6. Final Model: Error Analysis
+The models displayed in Figure 3.3 were each obtained by using all 64,000 training labels.
+C.6. Final Model: Error Analysis
 The Conservative model incorrectly labels 27 papers.
@@
 It is excluded from both the conservative and moderate model-generated samples, but a human-labeller flagged it as satisfying our criteria.
-Appendix C Additional Figures and Further Analyses
+Appendix D Additional Figures and Further Analyses
 This appendix collects additional figures and analyses.
-Figure C.3 displays series analogous to the series displayed in Figure 4, constructed with the Moderate and Liberal sample restrictions.
-The results are very similar.
-One concern with the comparison made in Figure 4 is that researchers may cite a growing number of publications over time.
+D.1. Robustness of Trends
+Figure D.1 displays series analogous to the series displayed in Figure 4.1, constructed with the Moderate and Liberal sample restrictions. The results are very similar.
+One concern with the comparison made in Figure 4.1 is that researchers may cite a growing number of publications over time.
 That is, more papers may cite clinical trials over time because citations are becoming, in some sense, “cheaper.” The relevant clinical literature, then, may not be any larger.
-To alleviate such concerns, Figure C.4 reports counts, where we scale each citation assigned to a paper by the total number of citations in the originating paper.
+To alleviate such concerns, Figure D.2 reports counts, where we scale each citation assigned to a paper by the total number of citations in the originating paper.
 That is, a paper that receives two citations, each from a paper that cites two papers, will have a weighted citation count of 1.
-Figure C.5 displays series analogous to Panel A of Figure 5 with the Moderate and Liberal sample restrictions.
-Figures C.6 and C.7 display heat maps analogous to Panel B of Figure 5, with the Moderate and Liberal sample restrictions and both 3-Year and 5-Year citation counts.
+Figure D.3 displays series analogous to Figure 4.2 with the Moderate and Liberal sample restrictions.
+Figure D.4 gives a displays analogous Figure 4.3 for the distribution of citations received by clinical trials in each calendar year originating from all publications.
+Figure D.5 gives a displays analogous to Figure 4.5 for the distribution of citations received by publications citing clinical trials (three year sample) originating from all publications.
+Figures D.6 and D.7 display heat maps analogous to Figure 4.3 and Figure D.4, with the Moderate and Liberal sample restrictions and both 3-Year and 5-Year citation counts.
 In each case, the results are very similar across specifications.
-Figure C.3. Growth in Clinical Research, by Sample Stringency
-Notes: Figure C.3 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” Each series is reported in terms of the percent change relative to its initial value.
+Figure D.1. Growth in Medical Research, by Sample Stringency
+Notes: Figure D.1 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” Each series is reported in terms of the percent change relative to its initial value.
 To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
-Figure C.4. Growth in Cost-of-Citation Weighted Clinical Research
-Notes: Figure C.4 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” Here, we report the number of scientific papers using a citation-weighted metric.
+Figure D.2. Growth in Cost-of-Citation Weighted Medical Research
+Notes: Figure D.2 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” Here, we report the number of scientific papers using a citation-weighted metric.
 In particular, we scale each citation assigned to a paper by the total number of citations in the originating paper.
@@
 To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
-Figure C.5. Public Funding and Low-Quality Research, by Sample Stringency
-Notes: Figure C.5 illustrates the stability of the heterogeneity in the composition and quality of published clinical trials between 2010 and 2019 in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” See Figure 3 for the true positive and false positive rates of these models.
+Figure D.3. Quality Measures, by Sample Stringency
+Notes: Figure D.3 illustrates the stability of the heterogeneity in the composition and quality of published clinical trials between 2010 and 2019 in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” See Figure 3.3 for the true positive and false positive rates of these models.
 The figure displays three time series: the proportion of clinical trials that are publicly funded, the proportion of clinical trials that are ever cited by a leading journal, and the proportion of publicly funded clinical trials that are ever cited by a leading journal.
-Figure C.6. Quality over Time, 3-Year Cites, by Sample Stringency
+Figure D.4. Citation Distribution Across Time, All Publications to Clinical Trials
+Notes: Figure D.4 displays a heat map measuring the distribution of citations received by clinical trials in each calendar year from all publications.
+The $y$-axis has been stretched to elongate the right-tail of the citation distribution.
+Colors are displayed in a log scale.
+The sample of published clinical trials is constructed with the conservative model.
+Figure D.5. Citation Distribution Across Time, All Publications to Citing Publications
+Notes: Figure D.4 displays a heat map measuring the distribution of citations received by papers that cite clinical trials, in each calendar year, from all publications.
+The $y$-axis has been stretched to elongate the right-tail of the citation distribution.
+Colors are displayed in a log scale.
+The sample of published clinical trials is constructed with the conservative model.
+The sample of papers that cite clinical trials is constructed using a three-year window.
+Figure D.6. Distribution of 3-Year Cites, Clinical Trials, by Sample Stringency
 Panel A: Citations From Publications in Leading Journals
 Panel B: Citations From All Publications
-Notes: Figure C.6 displays heat maps measuring the distribution of three-year citations received by clinical trials in each calendar year from leading journals and from all publications, respectively.
+Notes: Figure D.6 displays heat maps measuring the distribution of three-year citations received by clinical trials in each calendar year from leading journals and from all publications, respectively.
 The $y$-axis has been stretched to elongate the right-tail of the citation distribution.
@@
 We consider the sample of clinical trials constructed with the conservative, moderate, and liberal model.
-Figure C.7. Quality over Time, 5-Year Cites, by Sample Stringency
+Figure D.7. Distribution of 5-Year Cites, Clinical Trials, by Sample Stringency
 Panel A: Citations From Publications in Leading Journals
 Panel B: Citations From All Publications
-Notes: Figure C.7 displays heat maps measuring the distribution of five-year citations received by clinical trials in each calendar year from leading journals and from all publications, respectively.
+Notes: Figure D.7 displays heat maps measuring the distribution of five-year citations received by clinical trials in each calendar year from leading journals and from all publications, respectively.
 The $y$-axis has been stretched to elongate the right-tail of the citation distribution.
@@
 We consider the sample of clinical trials constructed with the conservative, moderate, and liberal model.
-Appendix D Prompt Repository
-This appendix serves as a repository for figures displaying each of the prompts considered in Section B.3.
+D.2. Growth in Citing Papers by Geography
+In Section 4.3.2, we document differential trends in the production of publications by type.
+Specifically, we find essentially stable geographic patterns in the production of clinical trials, whereas we find that large increases in the quantity of papers in our citing sample are driven by research produced outside of the United States.
+Here, we provide additional details on this geographic fact.
+In Section A.4, we provide details on the data that are used to construct these plots.
+There are two especially striking aspects of Figure 4.4: the quantity of citing papers published by researchers in China is high and sharply rising, and the quantity of (pooled) citing papers from what we refer to as the “rest of the world” set is also high and rising.
+Table D.1 breaks down the publication of citing papers by country.
+We elect to focus on countries that produced at least 200 citing papers in 2019—a cut that captures 93 percent of the citing papers in our data.
+That is, nearly all of the papers in the citing sample are produced by this relatively small number of countries.
+We highlight several patterns in Table D.1.
+First—consistent with its presentation as a stand-alone series in Figure 4.4—the quantity of citing papers produced by Chinese researchers, the change in the size of this set, and its growth over our sample period are distinctive.
+While the United States continues the produce the largest number of papers in the citing sample, the growth of this set ranks last among the countries listed in Table D.1.
+Whereas there is 11 percent change in the quantity of citing papers produced by United States-based researchers, there is more a more than 200 percent change for Chinese papers.
+Table D.1. Changes in the Number of Papers Citing Clinical Trials by Country
+Country Published, 2013 Published, 2019 Change, 2013 to 2019 % Change, 2013 to 2019
+China 1724 5620 3896 226.0
+Poland 158 311 153 96.8
+Japan 1167 2159 992 85.0
+Spain 487 892 405 83.2
+Italy 1241 2205 964 77.7
+Austria 191 319 128 67.0
+Korea 635 1045 410 64.6
+Netherlands 617 1011 394 63.9
+Belgium 211 335 124 58.8
+France 695 1075 380 54.7
+Denmark 223 343 120 53.8
+Austria 597 907 310 51.9
+Czech Republic 320 485 165 51.6
+Brazil 286 431 145 50.7
+Greece 158 228 70 44.3
+Sweden 276 391 115 41.7
+Germany 1555 2148 593 38.1
+Canada 791 1044 253 32.0
+India 302 393 91 30.1
+Iran 167 215 48 28.7
+Great Britain 1680 2104 424 25.2
+Turkey 231 259 28 12.1
+U.S.A. 11207 12536 1329 11.9
+Notes: Table D.1 records the number of papers that cite clinical trials published in the preceding three years in 2013 and 2019, by country.
+We include all countries that publish more than 200 papers that cite clinical trials in 2019.
+The fourth column displays the change in the number of ‘citing’ papers between 2013 and 2019.
+The fifth column displays the percent change in the number of ‘citing’ papers between 2013 and 2019.
+D.3. Quantifying Meta-Analyses
+In Section 4.3.1, we present three methods that allow us to identify “review” publications—publications that summarize the findings of existing studies—in our set of citing papers.
+Each approach is imperfect (for reasons that mirror the motivation for the technical contribution of this paper).
+Here, we provide additional details on these three approaches.
+D.3.1. Hand-Labeling Meta-Analyses
+From the set of publications that cite clinical trials (three-year sample), we drew 100 observations randomly from each publication year.
+We inspected records in the first publication year (2013) and last (2019).
+Each record was reviewed twice.
+During the first round of review, we collected notes about the content of the publication.
+We standardized these notes and observed that these notes could be coded into seven distinct categories.
+We reviewed the records a second time and assigned these standardized codes.
+Table D.2 reports the composition of these 100-record samples.
+The three largest categories in both samples are pre-clinical studies, observational studies, and meta-analyses.
+We estimate a decrease in the proportion of pre-clinical studies and an increase in the proportion of meta-analyses.
+By contrast, the proportion of observation studies appears stable.
+Table D.2. Composition of Publications that Cite Clinical Trials
+Published in 2013 Published in 2019
+Clinical Trial 1 0
+Editorial 3 2
+Pre-Clinical 26 17
+Non-Drug RCT 1 4
+Protocol 1 1
+Case Study 5 4
+Observational 31 30
+Meta-Analysis 32 41
+Notes:Table D.2 records the results of an exercise aimed at characterizing the composition of the set of publications that cite clinical trials (three-year sample).
+We randomly draw 100 publications from this sample that were published in 2013 and 2019, respectively.
+We categorize each publication into one of seven categories.
+We report the number of publications, in each year, that fall into each category.
+D.3.2. Text Searches for Meta-Analyses
+Recall that, in the process of constructing the training data that were used to fine-tune our large language models, we identified many records in PubMed that were meta-analyses, literature reviews, etc., based on the contents of their text.
+We reviewed all hand-labeled records from this training process and identified a set of keywords and phrases that appeared most frequently.
+We constructed a list of phrases and keywords that consistently indicated, to a human labeler, that a record is a meta-analysis or review article.
+This list is reproduced in Footnote 51.
+Using this list of search terms, we search the abstracts of all records in our citing sample.
+We flag any record with any of the keywords listed in Footnote 51 as a meta-analysis and flag any record that mentions at least two of the databases listed in Footnote 51 as well.
+The objective of this exercise is to yield a conservative estimate of the change in the quantity of meta-analyses over our period of interest.
+In Figure D.8, we plot this trend.
+Panel A plots the raw count of meta-analyses flagged by this procedure.
+Panel B plots the proportion of records in our citing sample that are designated as meta-analyses via this search.
+As expected, this approach captures a small number of records: in 2013, there are 1,606 records in our citing sample flagged as meta-analyses.
+This rises to 3,372 records in 2019.
+Although the baseline level this small, this corresponds to a 109 percent increase over this period.
+Figure D.8. Trends in Review Articles and Meta-Analyses
+Panel A: Counts Panel B: Proportions
+Refer to caption Refer to caption
+Notes: Figure D.8 displays the results of two exercises aimed at characterizing the change in the quantity of meta-analyses in the set of publications that cite clinical trials (three-year sample) over time.
+Estimates associated with the keyword search discussed in Section D.3.2 are displayed in purple.
+Estimates associated with the ‘NLM’ tag query discussed in Section D.3.3 are displayed in purple.
+Panel A displays counts of the number of publications that satisfy both sets of criteria.
+Panel B displays proportions of publications that satisfy both sets of criteria.
+D.3.3. NLM Tag Queries for Meta-Analyses
+Next, we observe that NLM tags, added when publication records are indexed in PubMed, also record information that may signal whether a record is a meta-analysis.
+On inspection, three of the 77 NLM tags are relevant: “literature review,” “meta-analysis,” and “systematic review.” Of course, as Section 3.2 makes clear, NLM tags may overcount records.
+As Section D.3.2 generates what is likely a lower bound, we proceed with this strategy, noting that it is likely to yield an upper bound.
+We flag all records in our citing sample indexed with one of these three NLM tags.
+We plot the resulting trend—as a count in Panel A and as a proportion of records in our citing sample in Panel B—of Figure D.8.
+In 2013, there are 9,407 records satisfying these criteria.
+In 2019, there are 14,991 such records.
+Over our period of interest, there is, thus, a 59 percent increase.
+Appendix E Prompt Repository
+This appendix serves as a repository for figures displaying each of the prompts considered in Section C.3.
 Throughout, the dummy text “{abstract}” is placed in the location where the abstract of a publication would be input.
-The baseline “True/False” Prompt 1.0 is displayed in Figure D.8.
-Refinements to this prompt are displayed in Figures D.9, D.10, and D.12.
-The prompt that asks the model to categorize excluded publications, i.e., Prompt 2.0, is displayed in Figure D.14.
-A refinement to this prompt is given in Figure D.15.
-The prompt that asks the model to explain why it has excluded a publication, i.e., Prompt 3.0, is displayed in Figure D.17.
-In turn, a refinement to this prompt is given in Figure D.18.
-Figure D.8. Prompt 1.0: True/False, Base
-Figure D.9. Prompt 1.1: True/False, Initial Examples
-Figure D.10. Prompt 1.2: True/False, Extended Examples, Short
-Figure D.11. Prompt 1.2: True/False, Extended Examples, Short (Cont.)
-Figure D.12. Prompt 1.3: True/False, Extended Examples, Long
-Figure D.13. Prompt 1.3: True/False, Extended Examples, Long (Cont.)
-Figure D.14. Prompt 2.0: Categorize Exclusion Restriction
-Figure D.15. Prompt 2.1: Categorize Exclusion Restriction, Examples
-Figure D.16. Prompt 2.1: Categorize Exclusion Restriction, Examples (Cont.)
-Figure D.17. Prompt 3.0: Provide Reason for Exclusion
-Figure D.18. Prompt 3.1: Provide Reason for Exclusion, Examples
-Figure D.19. Prompt 3.1: Provide Reason for Exclusion, Examples (Cont.)
+The baseline “True/False” Prompt 1.0 is displayed in Figure E.1.
+Refinements to this prompt are displayed in Figures E.2, E.3, and E.5.
+The prompt that asks the model to categorize excluded publications, i.e., Prompt 2.0, is displayed in Figure E.7.
+A refinement to this prompt is given in Figure E.8.
+The prompt that asks the model to explain why it has excluded a publication, i.e., Prompt 3.0, is displayed in Figure E.10.
+In turn, a refinement to this prompt is given in Figure E.11.
+Figure E.1. Prompt 1.0: True/False, Base
+Figure E.2. Prompt 1.1: True/False, Initial Examples
+Figure E.3. Prompt 1.2: True/False, Extended Examples, Short
+Figure E.4. Prompt 1.2: True/False, Extended Examples, Short (Cont.)
+Figure E.5. Prompt 1.3: True/False, Extended Examples, Long
+Figure E.6. Prompt 1.3: True/False, Extended Examples, Long (Cont.)
+Figure E.7. Prompt 2.0: Categorize Exclusion Restriction
+Figure E.8. Prompt 2.1: Categorize Exclusion Restriction, Examples
+Figure E.9. Prompt 2.1: Categorize Exclusion Restriction, Examples (Cont.)
+Figure E.10. Prompt 3.0: Provide Reason for Exclusion
+Figure E.11. Prompt 3.1: Provide Reason for Exclusion, Examples
+Figure E.12. Prompt 3.1: Provide Reason for Exclusion, Examples (Cont.)

### v3

@@
 JEL: C81, O32
-A recent literature contends that the returns to research in medicine are declining.
-This conclusion is premised on two empirical findings: the quantity of medical research is rising, while improvements in health outcomes are stagnant.
-We argue that the first finding is an artifact of conceptual and empirical problems in the measurement of research.
-We propose a method for annotating unstructured text based on fine-tuning a large language model.
+We develop a method for assigning high-quality labels to unstructured text.
+The method is based on fine-tuning an open-source language model with data extracted from a proprietary large language model.
 We apply this method to construct a census of published clinical trials.
-We argue that clinical trials function as a natural proxy for productivity-relevant research in medicine.
-In our data, the quantity, quality, and composition of clinical trials have been stable since 2010.
-We find that measured increases in the total quantity of medical research are driven by shifts in the composition of scientific publications that do not correspond to real increases in the quantity of productivity-relevant research.
-We document large increases in the growth of non-trial publications, driven by publications from China and publications that review existing literature.
-This increase in quantity is coincident with a decline in various measures of average publication quality.
-Taken together, we find no support for declining medical research productivity over this period.
-Keywords: Medical Innovation, Productivity, Clinical Trials, Large Language Models, Generative AI
+We revisit a literature that contends that pharmaceutical research productivity is declining, based on measured increases in the quantity of clinical trials, which outpace trends in output.
+In our data, the quantity and composition of trials are stable since 2010.
+Previous measurements are an artifact of biases driven by shifts in the composition of other forms of research.
+Keywords: Large Language Models, Medical Research Productivity, Clinical Trials
 JEL: C81, O32
+Perhaps the most serious task facing empirical work in the area of “technological change” and “invention and innovation” is the construction and interpretation of measures (indices) of advances in knowledge.
+— Pakes and Griliches,, 1980
 1. Introduction
-Two stylized facts support the argument that the returns to medical research are diminishing.
-[^1] Goldin et al., (2024) survey evidence related to declining productivity across the economy, including evidence of diminishing returns to research investment in sectors beyond medicine.
-The quantity of resources directed toward medical research, measured in dollars and scientific publications, appears to be increasing (Scannell et al.,, 2012; DiMasi et al.,, 2016; Bloom et al.,, 2020; Park et al.,, 2023).
-[^2] See Park et al., (2023), Bloom et al., (2020), and Scannell et al., (2012) for case studies examining the quantity of investment in science, medicine, and drug development, respectively.
-Bloom et al., (2020) collect case studies from other research-intensive fields.
-See, also, Chu and Evans, (2021), who document large increases in the quantity of papers in certain scientific fields and examine associated patterns in citation networks.
-Meanwhile, improvements in human health (e.g., increases in life expectancy) continue at the same steady pace (Mehta et al.,, 2020; Ho and Hendi,, 2018).
-If more resources are necessary to yield the same rates of technological progress, simple models of “idea production” imply that research productivity—the efficiency with which research is converted into new and useful knowledge—must be falling (Jones,, 1995; Bloom et al.,, 2020).
-[^3] In some models, these patterns can be reconciled without diminishing returns to research.
-For instance, Young, (1998) investigates the idea that investment in new varieties of existing products can yield the same trends.
-These facts, and their forecast of technological stagnation in a sector closely linked to health and welfare (Cutler and McClellan,, 2001; Mokyr et al.,, 2002; Acemoglu and Johnson,, 2007), have influenced prominent policy debates.
-[^4] These facts were most prominently documented in Bloom et al., (2020).
-Concerns about technological stagnation in modern policy discourse predate Bloom et al., (2020) and respond in part to Cowen, (2011) and Gordon, (2017).
-As we discuss below, the prospect of diminishing returns to scientific research has long been a subject of concern.
-Proposals that would restructure public subsidies administered by the National Institutes for Health (NIH) (e.g., Executive Office of the President,, 2019; Moss et al.,, 2020; The Economist,, 2022), relax the regulatory standards for approval of new medicines by the U.S.
-Food and Drug Administration (FDA) (e.g., Roy,, 2021; Chertman,, 2023; Graboyes and Reimschisel,, 2017), and prevent the Centers for Medicare and Medicaid Services (CMS) from negotiating drug prices (e.g., York,, 2023; Long,, 2023; Inflation Reduction Act,, 2022) are premised, in no small part, on the concern that science and medicine are in a state of decline.
-[^5] This concern has been explored extensively in the popular press.
-Collison and Nielsen, (2018), Broad, (2023), Thompson, (2021), and Piper, (2023), among others, interpret these facts in the popular press as a “warning sign” for the future of technological progress (Piper,, 2023).
-In this paper, we find that one of these stylized facts—the increase in the quantity of resources directed toward medical research—is an artifact of conceptual and empirical problems in the measurement of medical research.
-An ideal, and infeasible, measure of the quantity of research captures the amount of “effort” directed toward increasing the stock of useful (medical) knowledge.
-Informative proxies for this measure have two characteristics.
-First, as researchers differ in skill and research designs differ in quality, the proxy must account for heterogeneity in the likelihood that research projects yield “productivity-relevant” information.
-Second, units of the proxy must be comparable over time.
-If the quantity of research embedded in, say, a patent or paper decreases with time, increases in the quantity of patents or papers provide very little insight into research productivity.
-We find that existing measures of medical research fall short on both accounts.
-The primary contribution of this paper is the construction of a measure of medical research—a census of published clinical trials—that addresses both issues and fills this gap.
-We show that standard data construction tools classify research imprecisely, and that this imprecision yields inaccurate conclusions about trends in productivity.
-Thus, we introduce a method for annotating unstructured text at scale.
-When applied to a corpus of scientific text, this method allows us to identify and classify heterogeneous forms of research, relaxing a long-standing data constraint in the economics of innovation.
-We fine-tune an open-source large language model and devise a workflow that allows us to identify clinical trial records from the unstructured text of scientific publications with an accuracy comparable to that of a human labeler.
-In our data, the quantity, quality, and composition of clinical trials have been stable since 2010.
-We find evidence that measured increases in the quantity of scientific publications capture shifts in scientific incentives that are not indicative of real increases in the quantity of (productivity-relevant) research.
-Taken together, we find no support for claims of declining medical research productivity over this period.
-We begin by arguing that clinical trials are a natural proxy for the research object of interest in this setting.
-Trials are discrete, regulated experiments that generate useful statistical information about new and existing technologies (Chavez-MacGregor and Giordano,, 2016; Alsan et al.,, 2024).
-These experiments are designed in consultation with regulators after sufficiently robust preliminary evidence has been produced but before health technologies can be marketed.
+Considerable progress in the measurement of innovation has been made in the half-century since the publication of Pakes and Griliches, (1980).
+This progress has been enabled in large part by the proliferation of large-scale corpora of patents and scientific publications.
+Yet measurement remains an active constraint for empirical research (Bryan and Williams,, 2021).
+Patents and publications are attractive proxies for innovation, as they necessarily relate to instances of invention and research.
+Their utility is limited, however, by heterogeneity in their technical content and value (see e.g., Pakes,, 1986; Griliches,, 1990, 1994).
+Consequently, trends in counts of documents are often confounded by trends in their composition (Lerner and Seru,, 2022).
+Correction of this confounding can be viewed as a classification problem: how accurately can unstructured text be mapped to specific quantities of interest?
+[^1] The most prominent example of such a classification effort is the construction of the National Bureau of Economic Research (NBER) Patent Data, which, among other things, maps patent records to technology classes (Hall et al.,, 2001).
+Earlier, Scherer, (1984) classified, by hand, over 15,000 patents according to their technological category.
+In this paper, we demonstrate that large language models can facilitate classification tasks of this kind, achieving an accuracy and precision comparable to hand-made labels at a research-relevant scale and budget.
+We develop a procedure for constructing task-specific large language models, particularly suited to data construction problems that involve classification.
+The class of methods that we consider has two steps.
+First, a comparatively small number of labels are extracted from a large, typically proprietary, language model.
+[^2] By proprietary large language model, we are referring, at present, to models such as OpenAI’s GPT 3.5 and GPT-4.
+Alternative models include Anthropic’s Claude 3.5 Opus, Google’s Gemini, among many others.
+Second, these labels are used as training data for a smaller, more efficient, open-source language model.
+[^3] Here, we are referring to classes of open-source language models such as RoBERTa, based on Google’s BERT model, among others.
+This process is an instance of model distillation, which is a growing area of methodological research in computer science (Taori et al.,, 2023; Xu et al.,, 2024).
+This procedure allows researchers to capture the benefits of complex, frontier models, while retaining the resource efficiency and transparency of lighter-weight open-source models.
+The primary contribution of this paper is the application of this procedure in a context that is emblematic of the econometric challenges faced by studies of innovation: measurement of trends in clinical trial production.
+We show that standard methods for data classification identify clinical trials in publication data imprecisely.
+These imprecise measurements indicate, spuriously, that the quantity of clinical trials has increased sharply since 2010.
+In our data, the quantity, quality, and composition of clinical trials have been stable year-on-year.
+When considered alongside recent evidence on the stability of drug development costs during this period (Sertkaya et al.,, 2024), these new measurements of clinical trial production call into question the conclusions of an influential literature that finds diminishing returns to pharmaceutical research (Scannell et al.,, 2012; DiMasi et al.,, 2016; Bloom et al.,, 2020).
+[^4] Goldin et al., (2024) survey evidence related to declining productivity across the economy and note that, in the pharmaceutical sector, evidence for declining productivity primarily relies on measures of total numbers of clinical trials and estimates of R&D expenditures.
+The conclusions of this literature are based on measures that suggest that the quantity of clinical trials is increasing substantially, in a way unmatched by trends in measures of outputs.
+In light of our new data, we find that there is little rigorous empirical support for falling pharmaceutical research productivity in the most recent period.
+This finding is directly relevant to a set of active policy debates, which were initiated, in no small part, in response to concerns that the production of innovative output in this sector is slowing down.
+These debates include proposals that would restructure public subsidies administered by the National Institutes for Health (NIH) (e.g., Executive Office of the President,, 2019; Moss et al.,, 2020; The Economist,, 2022), relax the regulatory standards for approval of new medicines by the U.S.
+Food and Drug Administration (FDA) (e.g., Roy,, 2021; Chertman,, 2023; Graboyes and Reimschisel,, 2017), and prevent the Centers for Medicare and Medicaid Services (CMS) from negotiating drug prices (e.g., York,, 2023; Long,, 2023; Inflation Reduction Act,, 2022).
+Clinical trials are an especially attractive measure of research in medicine (Bryan and Williams,, 2021).
+In the pharmaceutical industry, clinical trials are discrete, regulated experiments that generate useful statistical information about new and existing medicines (Chavez-MacGregor and Giordano,, 2016).
 Trials top the hierarchy of medical evidence employed by firms, regulators, physicians, and patients (Jones and Podolsky,, 2015).
-Since the early 2000s, clinical trials have been reported in a consistent, standardized format across scientific publications.
-[^6] In the 1990s and 2000s, medical experts created a checklist that provides authors of papers reporting the results of clinical trials with a standardized structure and list of elements for publication.
-The most recent Consolidated Standards of Reporting Trials (CONSORT) Statement was published in 2010 and is now the reporting standard for most scientific journals that publish the results of trials.
-See Laine et al., (2007); Schulz et al., (2010) for details.
-To our knowledge, this type of standardization is unique to this form of science.
-No existing dataset allows us to track quantities of clinical trials over time.
-Although dollar-denominated expenditures are an obvious metric, available records of investment in trials are self-reported, in some form, by pharmaceutical firms and are essentially impossible to validate independently.
-Administrative databases—such as ClinicalTrials.gov—aim to provide an index of trials, but suffer from widespread noncompliance with reporting requirements (e.g., DeVito et al.,, 2020).
-Proprietary databases (e.g., Cortellis,, 2024) take as an input the contents of publicly-maintained registries and, thus, are biased by these same noncompliance patterns.
-We construct a new census of clinical trials indexed in scientific publications.
-We collect all records in the National Library of Medicine’s PubMed / MEDLINE database, from 2010 forward, that disclose the results of a clinical trial that studied the effects of a medicine in human subjects.
-This definition is intentionally restrictive and aims to include only those records that clearly report trials.
-Our goal is to identify clinical trials in our sample of interest—from the set of roughly 35 million PubMed records—with an accuracy and precision close to that of a human labeler.
-Manually identifying clinical trial records, at this scale, is infeasible.
-[^7] A small number of papers, including Oostrom, (2024), examine medical investment using small samples of clinical trial records that are constructed by hand.
-Standard methods perform poorly.
-In particular, we show that documented increases in the quantity of clinical trials during this time period, as reported by Bloom et al., (2020), are driven by a categorization error caused by keyword-based searches.
+Since the early 2000s, clinical trials have been reported in a consistent, standardized format across scientific publications and, with less consistency, in administrative databases (DeVito et al.,, 2020; DeVito and Goldacre,, 2021).
+Trials have, thus, been widely used to measure relative and absolute differences in the composition of research investment (see Finkelstein, (2004) and Budish et al., (2015) for two influential examples).
+Studies of information production and disclosure have examined the substantive content of trials (see e.g., Alsan et al., (2024), Oostrom, (2024), and Kao and Oostrom, (2024)).
+Recently, and prominently, measures of clinical trial quantities have been used to support arguments that medical and pharmaceutical research productivity are declining over time (Scannell et al.,, 2012; Bloom et al.,, 2020).
+Nevertheless, it is remarkably difficult to generate accurate measures of trends in clinical trial production.
+Administrative databases—such as ClinicalTrials.gov—have experienced large shifts in compliance with reporting regulations over time, which make it difficult to distinguish genuine increases in research effort from greater adherence to reporting requirements (see e.g., DeVito et al.,, 2020).
+Proprietary databases (e.g., Cortellis,, 2024) take as an input the contents of publicly-maintained registries and, thus, are confounded by the same shifts in compliance.
+Given these limitations, scientific publication data have served as a natural alternative or complement in recent empirical studies (e.g., Lichtenberg,, 2018; Bloom et al.,, 2020; Kao,, 2024; Guzman et al.,, 2024).
+With unstructured publication text data, the challenge becomes, again, one of classification: of the millions of records in large-scale publication databases, which disclose the results of a clinical trial?
+Manually identifying clinical trial records at a scale sufficient to study aggregate trends is infeasible.
+[^5] A small number of papers curate select samples of clinical trial data, for specific research questions.
+For example, Oostrom, (2024) collects records of psychiatric clinical trials as identified in prominent meta-analyses for a study of the relationship between sources of clinical trial funding and reported drug efficacy.
+We show that standard methods—based on publication metadata and machine learning tools—perform poorly.
+In particular, we show that keyword-based searches, including methods employed by Bloom et al., (2020), are very imprecise and generate spurious measurements of increases in the quantity of clinical trials.
 [^6] Specifically, Bloom et al., (2020) categorize many records as clinical trials that discuss trials (e.g., meta-analyses and literature reviews).
 We find that changes in the scientific publishing landscape, discussed in Section 4, increase the size of this categorization error over time.
-In our period of interest, this generates a spurious increase in the measured quantity of trials.
-Machine learning methods used in the health services literature similarly lack the precision necessary for our purposes (Thomas et al.,, 2021).
-Instead, we leverage recent advances in generative AI, which allow for data extraction from unstructured text at scale.
-We devise a workflow for extracting precisely defined objects from a large corpus of text, and train a bespoke large language model that is specifically optimized for this task.
+Specialized classifiers developed for use in the health services literature, as in Thomas et al., (2021), similarly lack the precision necessary to recover meaningful trends.
+Instead, we leverage recent advances in generative AI, which enable data extraction from unstructured text at scale.
+Our objective is to identify all records in the National Library of Medicine’s PubMed / MEDLINE database, from 2010 forward, that disclose the results of a clinical trial that studied the effects of a medicine in human subjects.
+We aim to classify records based only on the contents of the publication’s abstract, with an accuracy close to that of a human labeler.
+[^7] Many records in PubMed / MEDLINE have full-texts available for download, via the PubMed Central database.
+However, these full-texts are subject to copyright restrictions.
+Roughly one quarter of records in PubMed—$8$ million of the $34$ million total records—have full-texts that are available under licenses that would permit this type of research use (the PubMed Central Open Access Subset, at https://www.ncbi.nlm.nih.gov/pmc/tools/openftlist/).
+Thus, we restrict our focus to abstracts, for which no such restrictions apply.
 We proceed in stages.
-We devise an interface to iteratively revise model prompts.
-With the highest-performing prompts, we use proprietary language models, OpenAI’s GPT-3.5 and GPT-4, to classify a large number of records (Achiam et al.,, 2023; Bubeck et al.,, 2023).
-These noisy labels are used as training data to fine-tune a set of open-source models.
-An ensemble of the fine-tuned models matches the performance of the best proprietary model, with false positive and false negative rates below 5 percent, and produces a sample of approximately 150,000 publications that report the results of clinical trials from 2010 to 2022.
-These data allow us to document two sets of facts about the quantity of medical research over time.
+Given the abstract of a scientific publication, we prompt two proprietary language models, OpenAI’s GPT-3.5 and GPT-4 (Ouyang et al.,, 2022; Achiam et al.,, 2023; Bubeck et al.,, 2023), to label whether the record satisfies our desired sample restrictions.
+We find considerable variation in performance with seemingly small differences in prompt text.
+We iteratively revise prompt text until small changes yield no improvement in performance.
+GPT-4 performs quite well off-the-shelf.
+The performance of GPT-3.5 is unsuitable for our application.
+However, use of GPT-4 at scale—to classify millions of records—is computationally and financially impractical.
+We classify a moderate number of abstracts with the best-performing prompts.
+The labels extracted from the proprietary language models are used as training data to fine-tune a set of open-source models.
+Our preferred sample is constructed using an ensemble of fine-tuned models, which matches the performance of the best proprietary model.
+False positive and false negative rates are below 5 percent.
+When this model is used to label the full sample of PubMed records, it yields a sample of approximately 150,000 publications that report the results of clinical trials from 2010 to 2022.
+Three technical features of this fine-tuning process are worth highlighting.
+First, we document a phase-transition in model quality in the number of training labels used.
+Put differently, we observe a threshold in the number of training labels used in fine-tuning, beyond which the performance of the fine-tuned model increases dramatically.
+Second, models fine-tuned with labels extracted from GPT 3.5 exceed the performance of GPT 3.5.
+Third, models fine-tuned with labels extracted from GPT-4 match the performance of GPT-4 itself.
+That is, we show that researchers can distill the full benefits of large proprietary models into light-weight, resource-efficient models that can be realistically applied to standard data construction tasks.
+The resulting data allow us to document two sets of facts about the quantity of medical research over time.
 First, we find that the quantity of clinical trials has been effectively constant, year-on-year, across our period of interest.
 We collect a set of ex-ante and ex-post quality measures, derived from scientific publication data.
-On these dimensions, we find considerable heterogeneity.
+On these dimensions, we find considerable heterogeneity that is, to our knowledge, new to the literature.
 Roughly half of clinical trial publications in our sample are never cited by a leading medical journal.
@@
 However, we find that the distribution across clinical trials of measures of quality, importance, and geography is unchanging.
-Taken together, we interpret these patterns as evidence that “productivity-relevant” research in medicine is essentially constant in quantity and stable in composition over our time period.
-This contradicts existing findings, at least for our period of interest, and calls into question the conclusion that the productivity of medical research is falling.
-Second, our data allow us to examine the source of measured increases in the quantity of scientific publications (Bloom et al.,, 2020; Park et al.,, 2023; Chu and Evans,, 2021).
-We contrast clinical trial production with trends in the publication of other forms of medical research.
-While the quantity of clinical trials is constant across our period of interest, the quantity of other medical research—defined, in our preferred version, as the set of papers citing clinical trials—has increased by nearly a factor of two.
+Second, our data allow us to shed light on the source of the classification errors that drive the trends documented in Bloom et al., (2020).
+[^8] These data are relevant outside of this particular context, for studies of medical innovation and for studies in the health services literature that use clinical trial publications to construct prescribing guidelines for physician practice.
+Identifying trial records in publication data has been a challenge for this health services literature, where existing best practices have both high true and false positive rates.
+See Thomas et al., (2021).
+As similar trends in the total quantity of scientific publications have been documented elsewhere, including in Park et al., (2023) and Chu and Evans, (2021), these insights may be more generally applicable to discussions about how large increases in the quantity of publications should be interpreted.
+We contrast trends in clinical trial production with trends in the production of medical research that is similar—conceptually and in its textual presentation in scientific publications—to clinical trials.
+In our preferred version, this set of other medical research is defined as the set of papers that cite clinical trials.
 [^9] In Section 4, we inspect the contents of this set and document that nearly all records fall into one of three categories: observational studies, preclinical studies, and meta-analyses/literature reviews.
-In doing so, we confirm that our focus on clinical trials is appropriate.
-These increases in quantity have coincided with shifts in the quality and composition of these other types of research.
-We highlight three patterns.
-First, the large increase in the total number of medical papers in our sample is largely driven by researchers in China.
-Second, this rise in quantity coincides with a decrease (on various measures) in the average quality of publications.
-Third, there is a substantial increase in the quantity of meta-analyses and literature reviews during our period of interest.
-[^10] These patterns are consistent with the geometric increase in the quantity of meta-analyses produced by Chinese researchers, documented in Ioannidis et al., (2013).
-We interpret these patterns as evidence that—for this category of other medical research—the quantity of productivity-relevant research “embedded” in a single document may be declining over time.
-Quantities of publications, then, provide essentially no insight on the question of research productivity.
-[^11] Park et al., (2023) examine the “disruptiveness” of scientific research, using data on the universe of publications across six decades.
-Their findings are similar to ours, in the sense that they document that the absolute counts of highly disruptive papers remain constant over time, while average “disruptiveness” is declining.
-They interpret this as evidence that science is becoming less disruptive over time.
-Our paper differs in two important ways.
-First, we distinguish between types of publications on the basis of the economic value they produce.
-Second, we document that measured increases in the quantity of publications are disproportionately concentrated in types of publication that focus on diffusion, rather than production, of new information.
-Pooled averages, then, may be difficult to interpret.
-Our findings speak directly to ongoing policy discussions and branches of the macroeconomic literature, which consider the hypothesis that the rate of technological progress may be declining.
-Nearly all micro-level evidence on rising levels of research investment comes from studies of scientific publications and, in particular, clinical trials.
-[^12] See Goldin et al., (2024) for a summary.
-Notable exceptions are studies of agriculture and semiconductors.
-We show that these findings result from conceptual and empirical challenges in the quantification of innovation.
-Thus, our work is most closely related to a series of papers—which Syverson, (2017) summarizes as putting forth a “mismeasurement hypothesis”—that view the measured decline in aggregate productivity as illusory (see e.g., Mokyr,, 2014).
-[^13] Additional explanations include market imperfections and failures—including barriers to innovation (see e.g., Aghion et al.,, 2019), inadequate public sector support for research (Gruber and Johnson,, 2019), and productivity imbalances across sectors (Acemoglu et al.,, 2024).
-More directly, Myers and Pauly, (2019), Scannell et al., (2012), and Cockburn, (2006) investigate the determinants of declining productivity in the pharmaceutical sector, with a focus on large increases in factor prices for clinical trials rather than quantities.
-In doing so, we draw on, and contribute to, a long literature on the measurement of invention.
-Pakes and Griliches, (1980) begin a discussion of inventive activity led by U.S.
-firms by remarking: “Perhaps the most serious task facing empirical work in the area of ‘technological change’ and ‘invention and innovation’ is the construction and interpretation of measures (indices) of advances in knowledge.” In a more recent survey, Bryan and Williams, (2021) note that, although much progress has been made over half a century of research, construction of such measures remains an open challenge.
+While the quantity of clinical trials is constant over this period, the quantity of this other non-trial research has increased by a factor of two.
+Three facts about shifts in the composition and quality of these publications provide insight into the source of measured increases in research quantity.
+First, the large increase in the total number of non-trial papers in our sample is largely driven by researchers in China.
+Second, this rise in quantity coincides with a decrease (on available measures) in the average quality of publications.
+Third, there is a substantial increase in the quantity of meta-analyses and literature reviews.
+The main conclusion that emerges from this analysis is that errors in the classification of scientific publications can yield spurious findings about trends in research effort.
+1.1. Related Literature
+We draw on, and contribute to, a long literature on the measurement of invention.
 Decades of work have focused, in particular, on the development of proxies for the outputs of technological innovation.
 An especially prominent approach, due to Hall et al., (2000), uses citation-weighted counts of patents.
-Kogan et al., (2017) and Kelly et al., (2021) provide two recent, novel, alternatives.
 Our focus, by contrast, is on the difficulties associated with measuring the inputs of innovation, which can similarly bias estimates of macroeconomic trends.
-While scientific papers are a natural alternative to patents on the input side, existing methods for the use of patent data are not easily applied to publications (Popp,, 2015; Popp et al.,, 2010).
-[^14] In particular, citations across scientific papers are not analogous to citations across patents.
-Whereas patent citations are legal objects, which are defined in part by patent examiners and delineate the boundaries of a right enforceable by courts, citations across scientific papers reflect norms and networks.
-See Hall et al., (2000).
-The idea that patterns like those that motivate this paper—rising research quantities and stagnant outcomes—signal the “exhaustion of inventive opportunities” (a phrase due to Griliches, (1994)) is, similarly, an old one.
-Early appearances in the literature include Robert K.
+The new methods developed in this paper yield measures that call into question a fact at the center of debates involving health, innovation, and fiscal policy.
+Discussions of declining medical and pharmaceutical research productivity have been especially extensive in the popular press.
+Collison and Nielsen, (2018), Broad, (2023), Thompson, (2021), and Piper, (2023), for example, interpret trends in pharmaceutical research productivity, alongside patterns collected from other fields, as a “warning sign” for the future of technological progress (Piper,, 2023).
+The idea that patterns like those considered in this paper—rising research quantities and stagnant outcomes—signal the “exhaustion of inventive opportunities” (Griliches,, 1994) is old and closely tied to the development of methods for the measurement of innovation.
+Early appearances of this concern in the literature include Robert K.
 Merton’s 1935 discussion of the “possibility of a slackening in the rate of technologic advance” (Merton,, 1935), Alfred B.
 Stafford’s 1952 paper titled “Is the Rate of Invention Declining?” (Stafford,, 1952) and Jacob Schmookler’s various investigations (Schmookler,, 1952, 1954, 1966) into patterns suggestive of diminishing returns to research.
-Our reconciliation of these patterns hews closely to that of this earlier literature, which viewed the evidence base on productivity decline as inconclusive, given the shortcomings of available measures of innovation (Pakes and Griliches,, 1980; Griliches,, 1990, 1994).
-[^15] Griliches, (1990) summarizes his version of this conclusion more prosaically: “One can always worry that the world is coming to an end.
+Our treatment of these patterns hews closely to that of earlier literatures, which view the evidence base on productivity decline as inconclusive, given the shortcomings of available data (e.g., Pakes and Griliches,, 1980; Griliches,, 1990, 1994; Cockburn,, 2006).
+[^10] Griliches, (1990) summarizes his version of the conclusion that any finding of diminishing returns to innovative efforts are limited by the availability of high-quality data more prosaically: “One can always worry that the world is coming to an end.
 Someday it undoubtedly will, but it does not look as if the end is already upon us, at least not yet.” This paper differs on one important dimension.
-We show that advances in data construction technology, if deployed carefully, provide researchers with sufficiently precise measures to draw conclusions.
+We show that advances in data construction technology, if deployed carefully, can allow researchers to resolve some long-standing data construction challenges.
 A nascent literature investigates the use of large language models in economics (Dell,, 2024; Korinek,, 2023).
-We join a small set of papers, e.g., Bartik et al., (2023) and Dell, (2024), that use generative AI as a tool to construct high-quality economic data.
-Our approach is distinct in this literature, in that we fine-tune an open-source language model using data extracted from a larger, proprietary model.
-In doing so, we show that it is feasible for researchers to achieve the quality of hand-constructed data at a practically interesting scale, with a reproducible and transparent methodology, at a feasible cost.
-Our use of these technologies is distinct from other applications considered in e.g., Korinek, (2023), Manning et al., (2024), Bybee, (2023), Horton, (2023), and Mullainathan and Rambachan, (2024), which focus on the use of generative AI for other aspects of the scientific process, including hypothesis identification, survey participation, idea generation, and scientific writing.
+We join a small set of papers—notably, Bartik et al., (2023) and Dell, (2024)—that use generative AI as a tool to construct high-quality economic data.
+Our approach is distinct, in that we fine-tune an open-source language model using data extracted from a larger, proprietary model.
+In doing so, we show that, with model distillation, it is feasible for researchers to achieve the quality of hand-constructed data at a practically interesting scale, with a reproducible and transparent methodology, at a feasible cost.
+Prominent examples of model distillation in the computer science literature focus on the construction of chat bots by fine tuning open source models with data extracted from proprietary models (Taori et al.,, 2023; Chiang et al.,, 2023; Xu et al.,, 2023).
+Other papers have fine-tuned models for simple completion tasks like arithmetic (Liu and Low,, 2023) or querying an API (Patil et al.,, 2023).
+We are distinguished from this literature by considering a problem with, arguably, a higher-degree of complexity: the construction of data of sufficient quality to be of substantive interest to social science.
+Our use of large language models is also distinct from other applications considered in papers such as Korinek, (2023), Manning et al., (2024), Bybee, (2023), Horton, (2023), and Mullainathan and Rambachan, (2024), which focus on the use of generative AI for other aspects of the scientific process, including hypothesis identification, survey participation, idea generation, and scientific writing.
+It is worth emphasizing that, in this paper, we develop a procedure that is applicable to a binary classification problem, albeit one that is highly complex.
+The viability of these methods, and the extent of required modifications from the process outlined here, for application in more complicated classification problems of interest to economists, including those outlined in Dell, (2024), is a valuable direction for further research.
 More generally, we contribute to the large empirical literature that uses text as data (Gentzkow et al.,, 2019).
 We anticipate that our procedure for the construction of new, precise data on scientific research may be applicable to other contexts where unstructured text is available and manual construction of data is infeasible.
-We highlight a set of steps—in validation, prompt design, and fine-tuning of open source models—that yield data of sufficiently high quality at comparatively low cost.
-2. Measuring Research
-Studies of research productivity have, at their core, some notion of an “idea production function”—a mapping from research (e.g., scientists, dollars, research capital) to new, useful information.
-The canonical Romer, (1990) specification, which motivates a recent empirical investigation in Bloom et al., (2020), is given by
-$\frac{\text{d}A_{t}}{\text{d}t}\frac{1}{A_{t}}=\alpha_{t}S_{t}\leavevmode\nobreak\ .$ (2.1)
-Here, the variable $A_{t}$ denotes total factor productivity.
-The left-hand side of (2.1) captures the rate at which “ideas” are produced.
-On the right-hand side, the factor $S_{t}$ measures capital devoted to research.
-The structural parameter $\alpha_{t}$ is interpreted as research productivity—the efficiency with which inputs are transformed into outputs.
-[^16] See e.g., Romer, (1990), Jones, (1995), and Aghion and Howitt, (1992), among many others.
-Although each element of this framework is simple, constructing their empirical analogues is challenging.
-Decades of scholarship—summarized in Griliches, (1990, 1994) and, more recently, Bryan and Williams, (2021)—have grappled with alternative ways of measuring outputs, $A_{t}$.
-[^17] Pakes and Griliches, (1980) consider this challenge and employ slightly different notation than we use here.
-They focus on the quantification of an output variable $K$, which is the “level of economically valuable technological knowledge.” In this paper, our focus is on the comparable challenges associated with quantifying the resources that go into their production, $S_{t}$.
-An ideal measure of resources $S_{t}$ should capture the total quantity of economic resources devoted to increasing the stock of useful knowledge.
-Not all economic activities related to scientific research are equally relevant.
-Many forms of scientific output have as their goal the synthesis or diffusion of existing information.
-In medicine, for example, literature reviews and meta-analyses allow physicians to rapidly review the empirical evidence supporting particular treatment options.
-[^18] For one discussion of the appropriate interpretation of meta-analyses in clinical practice, see Thompson and Pocock, (1991).
-Medical guidelines, produced by entities like the American Heart Association (e.g., Smith Jr.
-et al.,, 2001), consolidate these findings into consensus documents that may be viewed as defining the “standard of care” (Moffett and Moore,, 2011).
-These types of literature syntheses are, anecdotally, of increasing importance, as health technologies previously available only in a small number of countries are made available to new patient populations (Ioannidis et al.,, 2013; Cockburn et al.,, 2016).
-It is important to be clear about the distinction we are making between these forms of research.
-This paper, and productivity calculations more generally, are interested in those forms of research that generate new information.
-These are the elements of the quantity $S_{t}$.
-From a welfare perspective, there are many other types of investment that are important, including these types of literature reviews that aid in the diffusion of information.
-These other investments may enable the “resources related by productivity growth .
-.
-.
-[to] find adequate employment,” thus increasing aggregate welfare, but are not themselves contributing to productivity growth (Griliches,, 1994).
-Even if, conceptually, it is clear that certain economic activities are more relevant to the quantity $S_{t}$ than others, available indicators of research—such as scientific publications and patents—may make it challenging to accommodate this heterogeneity empirically.
-This task is somewhat simpler in industries often studied by economists (see, e.g., the discussion of manufacturing firms in Syverson,, 2011), where dollar-denominated expenditures and price indices are available.
-In many settings central to economic growth, however, research occurs in university-based laboratories that do not maintain detailed financial ledgers, [^19] The UMETRICS data are one exception.
-See Babina et al., (2023) for an example of a study employing financial data from a university research setting.
-in countries where high-quality data may not be collected or made available, or in industries where confidentiality restrictions limit the public release of dollar-denominated expenditure records.
-Consequently, measures of research must rely on counts of researchers (who differ in skill), publications (which differ in relevance and resource-intensity), and administrative databases (which differ in completeness).
-The objective of this paper is to address these conceptual and empirical challenges in the context of medical research.
-Measuring the outputs of the production function (2.1) is, in some sense, more straightforward in medicine than in other sectors.
-Medical innovation appears coarsely in records maintained by regulators (e.g., counts of “new molecular entities” approved by the U.S.
-Food and Drug Administration) and, more directly, in measures of life expectancy and life-years saved from disease.
-[^20] Cutler and McClellan, (2001) observe that technological change has accounted for the bulk of increases in health spending over time, but that health has improved alongside these increases.
-In a series of case studies, they argue that the benefits of technological progress exceed their costs.
-See, also, Almond et al., (2010), who estimate that the returns to medical spending are positive on the margin.
-As this paper focuses on quantification of inputs $S_{t}$, we note only that it is a well-established empirical regularity that improvements in life expectancy, on many margins, appear to be “stagnating” (Mehta et al.,, 2020; Ho and Hendi,, 2018).
-In drawing conclusions about research productivity, we will follow Bloom et al., (2020) and assume that the left-hand side of (2.1) is constant.
-[^21] It is of course unlikely that medical innovations developed in the early 2010s are clearly discernible in recent mortality data, especially given the impact of the COVID-19 pandemic.
-Our task, then, is to identify a measure for inputs $S_{t}$ that best explains potential shifts in this output measure.
-Institutional features of medicine are helpful in doing so: medical decision-makers enforce a clear hierarchy of statistical evidence (Jones and Podolsky,, 2015).
-As guidelines to medical journal authors and editors from the American Medical Association state, “[p]rospective studies generally yield more reliable conclusions than retrospective studies.
-Cause and effect cannot be established by observational studies; consequently, study findings from observational studies should not be presented using causal language.” [^22] In 2024, the Journal of the American Medical Association—the “trunk journal” (see Section 4) of the American Medical Association and among the highest ranked journals in medicine—adopted a framework to allow studies other than randomized experiments to use this type of “causal” language.
-In doing so, editors of JAMA noted that “well-conducted randomized clinical trials remain the preferred approach” for the production of this type of evidence (Dahabreh and Bibbins-Domingo,, 2024).
-In decisions about the approval of new medicines (Food and Drug Administration,, 2024), the terms of insurance coverage (Choudhry,, 2017), and the appropriateness of a particular prescription (Martí-Carvajal,, 2020; Alsan et al.,, 2024), decision-makers defer to evidence generated via clinical trial.
-Other forms of medical evidence, generated through observational studies and surveys of the literature, receive less weight.
-Although detailed cost information on medical research is not available, available evidence suggests that these other forms of medical research are lower cost [^23] Recent estimates suggest that the average clinical trial may cost upwards of \$20 million dollars (Martin et al.,, 2017).
-Available estimates suggest that the cost of completing a systematic review or meta-analysis can reach \$150,000, where the bulk of the cost stems from the quantity of labor required to identify trials in publication data (Michelson and Reuter,, 2019).
-and involve fewer regulatory hurdles (Congressional Budget Office,, 2021).
-[^24] For example, while the approval of an institutional review board is necessary for any randomized clinical trial to proceed, many observational studies are exempt from this type of review.
-See guidance to researchers at Stanford University here: https://stanfordmedicine.box.com/shared/static/ase46ixmatdeahj7qdfgp4ydmy5qc3ei.pdf.
-Thus, we treat clinical trials as the appropriate measure of research inputs $S_{t}$.
-Clinical trials are discrete, regulated experiments.
-Research proceeds according to a pre-defined experimental protocol under the supervision of a regulator and across various phases.
-Trial phases differ in their scale (number of participants and duration) and, thus, in their cost (Food and Drug Administration,, 2024).
-Standardized measures (e.g., statistical objects like p-values) are recorded at various intervals.
-Trials are sufficiently expensive (Congressional Budget Office,, 2021) as to reflect “real” and substantial investment, and decisions to progress a study across phases (e.g., to move from a small Phase II study to a large Phase III study) capture decisions to pursue continued, directed research in an area of study (for an application of this insight, see Cunningham et al.,, 2021).
-As we approach this exercise with the ultimate goal of making progress on the question of research productivity, it is useful to note an additional feature about clinical trials, which suggests that sharp increases in quantities year-on-year (as documented in Bloom et al., (2020)) are unlikely.
-Not only are clinical trials sufficiently expensive as to require access to a source of financing, but trials also require an often scarce “input:” eligible patients.
-Malani and Philipson, (2019) collect empirical evidence of this type of “subject market effect,” and Alsan et al., (2024) collect anecdotal evidence on the difficulties of patient recruitment.
-[^25] The popular press has described this as a conundrum for cancer research: “too many drug trials, too few patients” (Kolata,, 2017).
+We highlight a set of steps—in validation, prompt design, and fine-tuning of open source models—that yield high-quality data at comparatively low cost.
 2. Identifying Clinical Trials
 We construct a new census of clinical trials disclosed in scientific publications.
-We begin, in Section 3.1, by detailing the deficiencies of existing, or alternative, approaches to measuring this form of medical research.
-In Sections 3.2 and 3.3 we propose a method, based on fine-tuning a large language model, for identifying relevant records at scale and with high fidelity.
+We begin, in Section 2.1 and Section 2.2, by providing context on the objectives of our data construction exercise and by detailing the deficiencies of existing, or alternative, approaches to measuring this form of medical innovation.
+In Sections 2.3 and 2.4 we propose a method, based on distilling data from a large language model, for identifying relevant records at scale and with high fidelity.
 The performance of the method is assessed in Section 2.5.
+2.1. Context
+Brief institutional context on the existing evidence related to pharmaceutical research productivity is helpful both in motivating our data construction and in interpreting our findings.
+Declining pharmaceutical research productivity has been a topic of policy concern since the late 1990s.
+In the past thirty years, discussions have largely centered on two sets of empirical patterns, which Cockburn, (2006) summarizes as follows: “More and more money is being invested in R&D, but the rate at which new drugs are introduced is failing to keep pace.” R&D, in this setting, refers to a particular type of biomedical research: investments in preclinical and, especially, clinical studies intended to translate discoveries made in basic science into commercialized medicines suitable for use in human patients.
+That inputs per new drug appear to be rising sharply over time is viewed, generally, as an established fact in the empirical productivity literature (Bloom et al.,, 2020; Goldin et al.,, 2024).
+Measures of productivity in this setting are somewhat crude.
+The measure of outputs in the most influential articulation of this phenomenon, by Scannell et al., (2012), is the number of “new molecular entities”—innovative drugs—approved by the U.S.
+Food and Drug Administration.
+[^11] Technically, a new molecular entity is a drug that contains no previously approved active moiety.
+In practice, this means that every active element of the drug (e.g., acetaminophen and caffeine in the commonly used over-the-counter migraine medicine Excedrin) must be novel.
+As Bloom et al., (2020) note, that this count has been stagnant or declining is viewed as an established fact.
+[^12] In data on new molecular entities approved by the FDA between 2005 and 2021, we observe a small increase in the number of approvals, largely driven by an uptick in approvals in 2020 and 2021.
+In earlier periods of data, we confirm that there is no discernible trend in new molecular entity approvals.
+Measures of inputs are, typically, some estimate of R&D expenditures.
+The simplest construction of a measure of R&D expenditure scales the (inflation-adjusted) cost of a clinical trial by the number of clinical trials.
+Recent evidence on these objects comes from Bloom et al., (2020) and Sertkaya et al., (2024).
+Bloom et al., (2020) find a sharp increase in the number of clinical trials, as measured in publication data over time.
+Using granular data on clinical trial site contracts and internal records of the U.S.
+Food and Drug Administration, Sertkaya et al., (2024) find that the cost of clinical trials and the cost of drug development was roughly constant between 2000 and 2018.
+Arguments of declining productivity in the pharmaceutical industry, then, depend substantially on the finding that the quantity of clinical trials is rising.
 2.2. Existing Approaches
-Although many datasets report information on clinical trials—and have been increasingly used by economists to measure innovation (see e.g., Budish et al.,, 2015; Cunningham et al.,, 2021)—no existing dataset, used off-the-shelf, yields accurate, transparent quantity measures for medical research in aggregate.
-3.1.1. Dollar Denominated Data
-Although disaggregated, dollar-denominated trial costs are a natural measure, available records of investment in clinical trials are self-reported, in some form, by pharmaceutical firms and are essentially impossible to validate independently (for a discussion, see Morten et al.,, 2024).
-There are, to our knowledge, three such sources of data that have been used in research.
-DiMasi et al., (2003, 2016) produce one widely cited (and contested, see Light and Warburton, (2005)) set of estimates of drug research and development costs, based on confidential data obtained from 10 firms.
-Myers and Pauly, (2019) and Scannell et al., (2012) collect aggregate estimates of annual research and development expenditures published by the Pharmaceutical Research and Manufacturers of America, a trade organization representing the industry.
-Both sets of estimates are often characterized as an upper bound (Morten et al.,, 2024) and cannot be produced independently by external researchers.
-Wouters et al., (2020) generates estimates of research and development costs, drawing on the information available in financial filings collected by regulators in the United States.
-Although these estimates are important and informative in a context that lacks precise measures of costs, expenditures can be defined in this way for only roughly 18 percent of the new drugs approved by the FDA between 2009 and 2018.
-These estimates, then, are not sufficient to answer the question this paper sets out to address.
-Even if dollar-denominated measures were available, it is not obvious how these measures should be scaled to allow for comparison across time periods.
-See Berndt and Cockburn, (2013) for a discussion of the difficulties associated with constructing a price index for clinical trial research.
-[^26] Berndt and Cockburn, (2013) construct a preliminary price index using data on trial budgets from a private firm.
-This firm no longer discloses data to academic researchers.
+Although many datasets report information on clinical trials—and have been increasingly used by economists to measure innovation (see e.g., Budish et al.,, 2015; Cunningham et al.,, 2021)—no existing dataset, used off-the-shelf, yields accurate, transparent aggregate quantity measures.
 2.2.1. Trial Registries and Proprietary Databases
@@
 In a series of papers assessing compliance over time DeVito et al., (2020) and DeVito and Goldacre, (2021) find that roughly 60 percent of trials registered on ClinicalTrials.gov fail to report required data elements at the time of their reporting deadline.
-[^27] The FDA is empowered, by statute, to impose fines of up \$10,000 for each day that required disclosures are late.
-To date, no fines have been collected, but the agency now sends notices of noncompliance to delinquent sponsors (Stephenson,, 2021).
+[^13] Note that noncompliance in this industry persists despite the existence of high statutory penalties.
+Sponsors can be fined up to \$10,000 per day for failing to register studies and disclose results.
+See the FDA Amendments Act of 2007 (FDAAA) and its Final Rule issued in 2021, at 42 U.S.C.
+§ 282(j)(5)(C)(i).
+To date, no fines have been collected, but the FDA now sends notices of noncompliance to delinquent sponsors (Stephenson,, 2021).
 Notices are available here: https://www.fda.gov/science-research/fdas-role-clinicaltrialsgov-information/clinicaltrials gov-notices-noncompliance-and-civil-money-penalty-actions
@@
 [^17] See Section A.1 for further details on on our treatment of the PubMed data.
-It is reasonable to expect that the corpus of scientific publications contains a census of research.
-The challenge is in extracting measures of research from this unstructured text.
+The challenge is to classify the specific forms of research captured by this corpus.
 Of these $34$ million records, which are clinical trials?
@@
 Figure 2.1 highlights that alternative approaches to data construction, in this setting, yield meaningfully different conclusions about levels, trends, and composition.
-That few of these series suggest “smooth” trends in the quantity of trials indicates, also, the extent to which the quality of data can affect these estimates.
 See Section A.2 for further details on the construction of these series.
@@
 The sample of interest is composed of all publications that report the results of a prospective, interventional clinical trial that evaluates the effects of investigational or approved drugs in a setting with exclusively human subjects.
-3.1 embeds several restrictions, intended to yield a measure of new investment in information relevant to human-facing medical care.
-We view this set of studies as the set most likely to generate information relevant to human mortality.
+2.1 embeds several restrictions, intended to yield precise measures of clinical trials relevant to the approval of new pharmaceuticals for use in human patients.
 Studies involving animals, literature summaries, and re-analyses of existing data, for example, are excluded.
@@
 Although our procedure can be applied to any type of scientific publication text, the standardization required by the CONSORT Statement changes the relationship between clinical trial research and publications in the early 2000s.
-To answer the research question that motivates this paper—is the quantity of medical research rising over time?—we elect to proceed with a narrower sample to avoid this issue of interpretability.
+To answer the research question at the center of our empirical analysis—is the quantity of clinical trials rising over time?—we elect to proceed with a narrower sample to avoid this issue of interpretability.
 To quantify our ability to identify the clinical trials in this set of publications, we hand-label approximately 3,000 randomly selected publications based on the content of their abstracts.
@@
 Each algorithm is estimated on the hand-labels assigned to the publications in the training and validation sample splits.
+Performance is measured in the testing sample split.
 For feature vectors, we use either TF-IDF embeddings computed in the corpus of abstracts in the hand-labeled sample or the embeddings of each abstract obtained from the SentenceTransformer language model (Reimers and Gurevych,, 2019).
@@
 Figure 2.2. Performance of Standard Machine Learning Methods
-Notes: Figure 3.2 displays the receiver operating characteristic of several machine learning models, trained on two types of embeddings, for classifying whether a publication satisfies the restrictions enumerated in 3.1.
+Notes: Figure 2.2 displays the receiver operating characteristic of several machine learning models, trained on two types of embeddings, for classifying whether a publication satisfies the restrictions enumerated in 2.1.
 The hyper-parameters of each model are determined with 10-fold cross-validation.
@@
 This is the same curve labelled “Ensemble” in Figure 2.4.
-3.3. Prompt Design and Fine-Tuning
+2.4. Model Distillation
 We construct a large language model optimized for our task.
@@
 The extraction of GPT-3.5 labels is substantially faster and cheaper.
-Second, we extract noisy labels for a large number of publications in our sample by querying the proprietary models.
+Second, we extract noisy labels for a moderate number of publications in our sample by querying the proprietary models.
 Third, the noisy labels are used to train an off-the-shelf large language model.
 The resulting model is then used to identify clinical trials in our baseline sample.
-This process of constructing specialized large language models is referred to as “fine-tuning” (see e.g., Taori et al.,, 2023, for a prominent application of this idea).
+This process for building specialized large language models is referred to as model distillation (Xu et al.,, 2024).
 2.4.1. Prompt Design
@@
 We identify three general prompt formats, which differ both in the amount of detail provided about our classification task and in the structure of the requested model completion.
-The simplest prompt provides a version of our sample definition, 3.1, and the text of an abstract, and asks the model to return ‘TRUE’ if the abstract satisfies these criteria.
+The simplest prompt provides a version of our sample definition, 2.1, and the text of an abstract, and asks the model to return ‘TRUE’ if the abstract satisfies these criteria.
 Otherwise, it returns ‘FALSE’.
@@
 We refer to this prompt as Prompt 3.0.
-We devise initial language for each prompt iteratively, using small numbers of records from our hand-labelled validation data.
+We devise initial language for each prompt iteratively, using a small number of records from our hand-labelled validation data.
 We test each of these three prompts in our 1000-record hand-labelled validation dataset, using both GPT-3.5 and GPT-4.
@@
 We use Prompt 2.0 to extract weak labels using GPT-3.5, and Prompt 1.2 to extract weak labels using GPT-4.
-Figure 3.3 displays estimates of the true and false positive rates computed in the test data, for both proprietary models, with black dots.
-These prompts are used to extract 64,000 labels—TRUE/FALSE flags—using GPT-3.5 and GPT-4.
-We refer to these as our set of “noisy labels.”
 Table 2.1. Prompt Performance in Validation Data
@@
 The performance measurements for these prompts are displayed in boldface.
-Figure 3.3. Receiver Operating Characteristic
-Notes: Figure 3.3 displays estimates of the receiver operating characteristic of several models used to classify the publications indexed by PubMed / MEDLINE according to whether they satisfy the restrictions enumerated in 3.1.
-All metrics are computed in the testing split of the hand-labelled publications.
-Estimates of the true positive rate and false positive rate of OpenAI’s proprietary models GPT-3 and GPT-4 are indicated with black dots.
-The curves give the performance of four fine-tuned, open-source large language models, in addition to an ensemble model.
-The blue dots indicate the true positive rate and false positive rate of the models used to construct the “Conservative,” “Moderate,” and “Liberal” samples of clinical trials.
-GPT-4 performs remarkably well.
-For every 71 true positives, GPT-4 identifies 29 false positives, while achieving a true positive rate of approximately 0.974.
-This represents a substantial improvement over existing methods.
+The performance of GPT-4 represents a substantial improvement over existing methods.
 However, practical and substantive considerations mitigate the applicability of proprietary models for classification of the complete baseline sample.
-Practically, it is—and likely will remain—prohibitively expensive to deploy GPT-4 at this scale.
+Practically, it is—and likely will remain—prohibitively expensive, both computationally and financially, to deploy GPT-4 at this scale.
 [^21] We incur a cost of roughly \$4,500 to extract noisy labels for 64,000 abstracts.
@@
 2.4.2. Fine-Tuning
-We compute noisy labels for 64,000 randomly selected publications in our baseline sample using the best performing prompts for both GPT-3 and GPT-4.
-These noisy labels are used as data to train off-the-shelf language models from two architecture classes: (1) RoBERTa (Liu et al.,, 2019) and (2) BioMedBERT (Gu et al.,, 2021).
-Both models are available in either “baseline” or “large” varieties.
+We compute noisy labels for a moderate number of randomly selected publications in our baseline sample using the best performing prompts for both GPT-3.5 and GPT-4.
+These noisy labels are used as data to train off-the-shelf BERT models from two architecture classes: (1) BigBird (Zaheer et al.,, 2021) and (2) BioMedBERT (Gu et al.,, 2021).
 [^22] There are many open-source large language models (e.g., LLaMa, Mistral, Pythia).
-Many are several orders of magnitude (e.g., 7-70 billion parameters) larger than RoBERTA ($\sim$ 350 million parameters).
-We selected RoBERTA and BioMedBERT via trial-and-error that included testing the performance of these larger models.
-In our setting, RoBERTA outperforms LLaMA (both 7 and 70 billion parameter versions, trained with QLoRA).
-Both models exhibited comparable performance, but RoBERTa was substantially faster and simpler to train.
-The BioMedBERT model is an instance of a BERT model (Devlin et al.,, 2019), that has itself been fine-tuned on the abstracts of the publications indexed in PubMed.
-We systematically vary model architecture, size, and pre-training regimen.
-We observe no large differences in performance stemming from these hyper-parameters.
-Sections C.4 and C.5 give further details concerning the fine-tuning process and the results of these experiments, respectively.
+Many are several orders of magnitude (e.g., 7-70 billion parameters) larger than BERT ($\sim$ 350 million parameters).
+We selected BigBird and BioMedBERT via trial-and-error that included testing the performance of these larger models, including LLaMA (both 7 and 70 billion parameter versions, trained with QLoRA).
+Both models exhibited comparable performance, but BigBird was substantially faster and simpler to train.
+Both models use medium-scale Transformer architectures, i.e., between 100 and 300 million parameters, pre-trained with the masked-language modeling (MLM) objective (Vaswani et al.,, 2023; Devlin et al.,, 2019).
+The models differ in their architectural details and pre-training corpora.
+BigBird uses a sparse attention mechanism to reduce the computational cost of processing long text sequences.
+[^23] The computational complexity of regular attention is quadratic in sequence length, while BigBird reduces this to linear complexity.
+This enables efficient handling of long documents up to 4,096 tokens.
+In contrast, standard BERT models can only process 512 tokens.
+BigBird is useful because abstracts of clinical trials regularly exceed 512 tokens.
+Prior to pre-training, the model was warm-started from the RoBERTa checkpoint (Liu et al.,, 2019) and then pre-trained on the standard BERT corpus.242424 The standard BERT corpus consists of the Books (Zhu et al.,, 2015), CC-News (Guu et al.,, 2020), Stories (Trinh and Le,, 2019), and Wikipedia (Wikimedia Foundation,, 2023) datasets.
+We evaluate two model sizes: BigBird Base (125 million parameters) and BigBird Large (355 million parameters).
+BioMedBert uses a domain-specific pre-training corpus sourced from articles on PubMed (Gu et al.,, 2021).
+It uses the same model architecture as RoBERTa, but a different, domain-specific token vocabulary optimized for PubMed articles.
+Like BigBird, we evaluate two model sizes: BioMedBert Base (125 million parameters) and BioMedBert Large (355 million parameters).
+We fine-tune the pre-trained architectures to classify abstracts according to the inclusion and exclusion criteria enumerated in 2.1.
+We replace the language modeling classification head with a binary classification head consisting of a single linear layer and softmax activation.
+We fine-tune the full model (i.e.
+not just the classification head) with cross-entropy loss.
+[^25] We use the Adam optimizer with learning rate $1\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$ (determined via hyperparameter sweep).
+We use a maximum sequence length of 4096 tokens.
+If an abstract does not fit into the maximum sequence length provided by a model, the abstract is truncated.
+All models were trained on a single $V100$ GPU with 16GB of HBM.
 2.5. Performance
-Given the text of an abstract, our fine-tuned language models output a probability that the publication satisfies the restrictions enumerated in 3.1.
+Given the text of an abstract, our fine-tuned language models output a probability that the publication satisfies the restrictions enumerated in 2.1.
 Publications whose probabilities fall above a chosen threshold are classified as belonging in our sample.
-Figure 3.3 displays estimates of the true and false positive rates, computed with the test data, as we vary this threshold for both sizes of the RoBERTa and BioMedBERT models.
-[^37] Each of the models whose results are displayed in Figure 3.3 are trained with noisy labels obtained with GPT-4.
-We document in Section C.5 that models trained with noisy labels obtained from GPT-3.5 or with the 1000 hand-labeled records in the training data perform significantly worse.
+Figure 2.3 displays estimates of true and false positive rates, computed with the test data, as we vary this threshold for labels assigned by fine-tuned version of the base BigBird model.
+The top and bottom panels are trained on noisy labels extracted from GPT-3.5 and GPT-4, respectively.
+We vary the quantity of noisy labels used to train the language model.
+Additionally, in the bottom panel, we display the performance of a fine-tuned model trained with the 1000 hand-labelled observations in the training dataset.
+Figure 2.3. Receiver Operating Characteristic by Data Size and Origin
+Notes: Figure 2.3 displays estimates of the the receiver operating characteristic of several fine-tuned language models, based on the BigBird architecture.
+The models differ in terms of the source of their training data.
+Solid lines display the performance of models trained on noisy labels extracted from either GPT-3.5 or GPT-4.
+The dotted line displays the performance of a model trained on the hand-labeled data in the training dataset.
+Estimates of the true positive rate and false positive rate of GPT-3.5 and GPT-4 are indicated with black dots and are measured using the testing data.
+Fine-tuned models trained with noisy labels extracted from GPT-3.5 outperform GPT-3.5.
+By contrast, fine-tuned models trained with noisy labels extracted from GPT-4 match the performance of GPT-4, and significantly outperform models trained with labels extracted from GPT-3.5 or with hand-labels.
+There appears to be a threshold where performance dramatically improves at around 8,000 training labels.
+Figure 2.4. Receiver Operating Characteristic by Model Architecture
+Notes: Figure 2.4 displays estimates of the receiver operating characteristic of several models used to classify the publications indexed by PubMed / MEDLINE according to whether they satisfy the restrictions enumerated in 2.1.
+All metrics are computed in the testing split of the hand-labelled publications.
+Estimates of the true positive rate and false positive rate of OpenAI’s proprietary models GPT-3 and GPT-4 are indicated with black dots.
+The curves give the performance of four fine-tuned, open-source large language models, in addition to an ensemble model.
+The blue dots indicate the true positive rate and false positive rate of the models used to construct the “Conservative,” “Moderate,” and “Liberal” samples of clinical trials.
+Figure 2.4 displays analogous estimates for both sizes of the BigBird and BioMedBERT models.
+Both models are trained with 64,000 training labels extracted with GPT-4.
+Again, the fine-tuned models are able to match the performance of GPT-4 in the test data.
 The figure additionally displays the performance of an ensemble model estimated in the training data.
@@
 The ensemble model is used to produce our final sample.
-The fine-tuned models are able to match the performance of GPT-4 in the test data.
 We choose three thresholds according to the stringency with which they enforce the sample restrictions.
@@
 In 14 cases, however, errors are associated with records that were difficult to categorize for a human labeller.
-For example, PubMed record 32737793 is flagged as satisfying 3.1 with all three thresholds, but is a literature review.
-By contrast, PubMed record 27880726 was categorized as satisfying 3.1 twice by a human labeller.
+For example, PubMed record 32737793 is flagged as satisfying 2.1 with all three thresholds, but is a literature review.
+By contrast, PubMed record 27880726 was categorized as satisfying 2.1 twice by a human labeller.
 It is excluded from both the conservative and moderate model-generated samples.
@@
 We report results associated with the moderate and liberal thresholds in the appendix as tests of robustness.
-Our final, conservative, sample consists of 152,027 publications classified as satisfying the restrictions enumerated in 3.1.
+Our final, conservative, sample consists of 152,027 publications classified as satisfying the sample restrictions.
 Each table displayed in Appendix A compares the contents of our sample to the counts of potential clinical trials plotted in Figure 2.1.
-Although certain elements of PubMed metadata—including the union of all NLM tags—capture many of the records in our final sample, we confirm that they miss many records that we flag as trials and include many records that do not satisfy 3.1.
+Although certain elements of PubMed metadata—including the union of all NLM tags—capture many of the records in our final sample, we confirm that they miss many records that we flag as trials and include many records that do not satisfy 2.1.
 No combination of existing search strategies, then, classifies records with the same accuracy or precision as our final sample.
-4. Disentangling Trends in Medical Research
-Figure 3.1 indicates that, across measures, the total quantity of medical research appears to be increasing, including since 2010.404040We find similar patterns using alternative cuts of records indexed in PubMed.
-The total number of unique authors listed on scientific papers, the total number of citations given and received by scientific papers, the total number of scientific journals, and the total number of pages in scientific journals all exhibit similar patterns.
+3. Disentangling Trends in Publication Data
+Figure 2.1 indicates that, across measures, the total quantity of potentially-relevant records in PubMed has increased substantially over time, including since 2010.282828Here, we focus on patterns in the production of publications that are similar to those that disclose the results of a clinical trial.
+We find similar patterns in PubMed as a whole, and in alternative cuts of records.
+Across fields, the total number of scientific papers, the total number of unique authors listed on scientific papers, the total number of citations given and received by scientific papers, the total number of scientific journals, and the total number of pages in scientific journals exhibit similar patterns.
 These patterns are consistent with documented increases in the quantity of scientific papers in Bloom et al., (2020) and Park et al., (2023), as well as economy-wide increases in research effort reported in Goldin et al., (2024).
 In this section, we use our census of clinical trials to decompose this trend.
-We contrast trends in the production of clinical trials with those documented for other forms of medical research.
-Differences between these series are informative in two ways.
-First, existing estimates of the quantity of medical research, such as those plotted in Figure 3.1, capture some combination of these two series (e.g., Bloom et al.,, 2020; Park et al.,, 2023; Chu and Evans,, 2021).
-Second, any proxy for the quantity of medical research is difficult to interpret if its relationship to the underlying research changes over time.
-This decomposition highlights the limitations of empirical strategies that treat counts of scientific publications as a proxy for research quantity.
-There are many potential ways to construct samples of other (i.e., non-clinical trial) medical research.
+We contrast trends in the production of clinical trials with those documented for other, similar forms of medical research.
+There are many potential ways to construct a sample to contrast with data on clinical trial production.
 Our preferred approach defines this category as the collection of publications that cite clinical trials.
-(We confirm in Section 4.3.1 that this set includes only medical research.) Conceptually, of course, this “citing” sample is a subset of non-clinical trial medical research.
+This sample includes publications that “look like” clinical trials—a fact that we confirm on inspection—and which necessarily engage with the findings of these studies.
+Thus, in an imprecise classification task, it is precisely this set of records that are likely to be erroneously included.
+For example, a keyword search of publication abstracts for “clinical trials” may capture studies that summarize the findings of clinical trials, without reporting new information.
+Based on our inspection of these records, these types of errors are especially likely for publications in which the writing quality is low, for which it may be challenging—even for a human labeler—to determine whether the record reports novel results or summarizes existing findings.
 Alternative methods of constructing a measure of “other” research introduce additional, complicating conceptual problems.
@@
 We drop from our “citing” sample any records that we also flag as reporting the results of a trial.
-4.1. Decomposing Types of Medical Research
+3.1. Decomposing Trends
 Figure 3.1 documents that the quantity of clinical trials—the black line—has remained essentially constant over the past ten years.
-In contrast, the quantity of other medical research—the colored lines—has increased substantially.
+In contrast, the quantity of other non-trial research—the colored lines—has increased substantially.
 Observe that we display the percent growth in the quantity of clinical trials in our sample published in each calendar year, alongside the growth in the set of papers that cite clinical trials, constructed in various ways to account for the truncation discussed above.
@@
 Second, existing estimates of rising quantities of medical and scientific research (e.g., Bloom et al.,, 2020; Chu and Evans,, 2021; Park et al.,, 2023) are not driven by shifts in the quantity of clinical trials.
-Recall the very simple conceptual framework from Section 2.
-We assume, following e.g., Bloom et al., (2020), that the left-hand side of the “idea production function” (2.1) is constant in this setting.
-If clinical trials, as captured in our data, are a stable proxy for inputs $S_{t}$—in the sense that published trials capture the same amount of underlying research in each period—then an unchanging number of clinical trials implies that research productivity is constant.
-We return to the appropriate interpretation of rising trends in other medical research below.
-Here, we collect evidence that clinical trials are a stable proxy for inputs.
 Figure 3.2 documents that—on two additional dimensions—clinical trial production has been stable over the past decade.
@@
 Second, nearly 70 percent of clinical trials are never cited by a leading journal, suggesting substantial, but time-invariant, heterogeneity across trials.
-These proxy variables capture slightly different conceptual objects, which are helpful in understanding different aspects of the production function (2.1).
-Stability in public funding suggests minimal changes, at least in the cross section, in allocation of resources to clinical trials over time.
-[^44] NIH RePORTER data allow us to tab total expenditures associated with grants acknowledged in these publications.
-These grant-based expenditure measures, too, are stable.
+Stability in public funding suggests, on its own, that the allocation of public funding to clinical trials has not considerably shifted over this time period.
+One can imagine alternative explanations for this stability.
+For example, perhaps the share of all projects receiving public funding has remained stable, but the cost of each trial has increased substantially.
+Alternatively, perhaps public agencies are funding a higher or lower share of studies, but that these changes are perfectly offset by changes in either study success or publication bias.
+We view the first potential explanation as unlikely for two reasons.
+Somewhat suggestively, NIH RePORTER data allow us to tab total expenditures associated with grants acknowledged in these publications.
+These grant-based expenditure measures, too, are stable over this period.
 Note, however, that linking research grant dollars to specific scientific papers is an imprecise exercise.
 See Li, (2017) for an extended discussion of this measurement challenge.
-Stability in citation patterns, in contrast, suggests that the “usefulness” of information produced by trials has been stable across this period.
+More directly, recent work by Sertkaya et al., (2024) suggests that the costs of drug development—measured using the contracts associated with particular study sites in clinical trials—have been essentially constant from 2000 to 2018.
+The second potential explanation is unlikely from a more conceptual perspective.
+We find essentially complete stability in this measure over time: any changes in publication bias or trial failure rates must, then, perfectly offset changes in public funding patterns, to yield the flat series here.
 Figure 3.3 provides a more granular view of this heterogeneity.
-We plot quantiles of the distribution of the number of (three-year) citations to publications in our census from publications in leading journals, for each year between 2010 and 2019.454545Appendix D presents these results using alternative cuts of our data and five-year citations.
+We plot quantiles of the distribution of the number of (three-year) citations to publications in our census from publications in leading journals, for each year between 2010 and 2019.323232Appendix D presents these results using alternative cuts of our data and five-year citations.
 The patterns remain unchanged.
@@
 However, these citations are not originating in leading journals.
-We interpret this set of facts as evidence that our measured quantity of clinical trials captures an object that, though heterogeneous, is stable over time.
-Thus, suggestively, research productivity is similarly stable.
-We cannot reject that there have been shifts in upstream, basic inputs into clinical trials or changes in factor prices during this period, which might suggest that the underlying quantity of resources directed toward research is, in fact, changing over time.
-In our view, however, there is little rigorous evidence that supports such an argument.
+We interpret this set of facts as evidence that trends in clinical trial production have been stable over this time period. There are four potential concerns with this interpretation.
+First, if the costs of either upstream, basic inputs into clinical trials or the costs of clinical trials themselves have increased substantially over this time, the total resources devoted toward research will have increased, even though Figure 3.1 suggests that counts of trials produced are stable.
+In our view, this is unlikely to be the case.
+On the issue of clinical trial costs, the most up-to-date, comprehensive evidence on drug development costs, in our read, is given in Sertkaya et al., (2024), who use internal data from the U.S.
+Food and Drug Administration and contracts associated with clinical trial sites to compute estimates for the period from 2000 to 2018.
+In these data, they find that over this period, “the cost of drug development remained relatively stable or may have even decreased.” [^33] This finding revises a frequently cited claim made in popular press interviews (e.g., Herper,, 2022), in academic articles that rely on proprietary data sources (e.g., DiMasi et al.,, 2016), and in surveys on research productivity (Goldin et al.,, 2024)—which suggest that clinical trial and drug development costs are sharply increasing over time.
+This revision is in the spirit of an argument made by Cockburn, (2006), in response to an earlier debate about drug development costs.
+Thus, we do not consider the prospect of rising study prices a concern for the interpretation of this pattern.
+On the issue of basic science inputs into clinical trials, there is little empirical evidence that increases in cost or decreases in the productivity have occurred.
 A small number of papers examine frictions in markets for basic research.
@@
 To our knowledge, this literature has not provided evidence of substantial shifts in inputs or factor prices over time, which in turn would increase the cost or decrease the productivity of basic science.
-[^46] Anecdotal evidence suggests that there may have been substantial decreases in the cost of basic science during this time period, which would suggest the opposite conclusion in our simple framework—that research might be more productive over this period.
-See Scannell et al., (2012).
-On factor prices, as discussed in Section 3, nearly all estimates of research and development costs are self-reported by the pharmaceutical industry.
-Any measured increases in these expenditures that cannot be independently verified should be viewed with caution, especially in light of our findings of stability.
-[^47] Herper, (2022) collects anecdotal evidence, which has been widely cited, on sharply rising costs of clinical trials.
-Although the magnitudes included in this discussion are striking, they capture the impressions of a small number of industry stakeholders and, to our knowledge, have not been verified in a more detailed analysis.
-For clinical trials to be an informative proxy for the quantity of medical research—and for the black line in Figure 4.1, then, to capture the relevant measure of inputs—we must make two additional assumptions.
-First, the existence of publication bias means that scientific publications do not perfectly capture research inputs (Andrews and Kasy,, 2019).
-In our setting—where the object of interest is the trend in research over time—we need only make a relatively weak assumption about the extent of publication bias: we assume that changes in publication bias, year-on-year, do not perfectly offset changes in the quantity of actual research.
-Second, we must assume that the number of publications associated with each clinical trial is constant across our period of interest (i.e., that the amount of research “embedded” in each publication is not systematically changing over time).
+In fact, changes in the production of basic scientific inputs around the year 2000—enabled by breakthroughs such as the Human Genome Project and the advent of high-throughput screening for drug discovery (see Scannell et al.,, 2012)—suggest, if anything, that costs of basic science inputs have fallen.
+Moreover, it is reasonable that such trends would affect trends in trial quantity, quality, composition, or cost.
+Given the absence of such patterns, it is unlikely that changes in basic science threaten the interpretation of our findings.
+[^34] Constructing a comprehensive measure of medical research productivity, which grapples with these various input markets, remains an open question.
+To our knowledge, the work that is closest in spirit is Berndt and Cockburn, (2013), who aim to develop a price index for clinical trial research, with the eventual goal of developing price indices for biomedical research more generally.
+Second, one might be concerned about how to interpret the x-axis in our plots.
+In particular, the relationship between clinical trial completion dates and clinical trial start dates varies systematically within and across disease classes (Budish et al.,, 2015).
+Here, we follow existing literatures that use dates affixed to patents and publications as measures of the timing of research.
+Practically speaking, as Footnote 14 suggests, records of the start dates of clinical trials are back-filled in administrative databases and, thus, are available only with a considerable lag.
+We focus on the date of publication, then, for ease of interpretation.
+Third, our findings are threatened by publication bias.
+The existence of publication bias means that counts of scientific publications only partially measure the quantity of research (Andrews and Kasy,, 2019).
+Our main finding is that the quantity of clinical trials is stable over time.
+If a trend were in fact present, and our measure were confounded by publication bias, then it would need to be the case that the trend was perfectly offset by the bias.
+It is reasonable to view cancellation, in this way, as unlikely.
+Fourth, and finally, we must assume that the number of publications associated with each clinical trial is constant across our period of interest.
 We view this assumption as reasonable for three reasons.
 First, by 2010, nearly all scientific journals in our sample had adopted rules requiring authors of clinical trial publications to adhere to a standardized format when reporting trial results (Schulz et al.,, 2010).
-Under these rules, the structure of a paper and its contents are pre-determined; there is little scope for authors to, for example, split results across several publications.
-Second, when constructing our sample of clinical trials, we have aimed to exclude those publications that re-analyze data generated by an existing study.
-Third, more suggestively, we use data on clinical trial publications supported by National Institutes of Health grants to examine the quantity of grant dollars per reported clinical trial publication over this period.
-We find no evidence of a change in this relationship between funding and publication during our period of interest.
+Under these rules, there is little scope for authors to, for example, split results across several publications.
+Second, when constructing our sample of clinical trials, we have excluded publications that re-analyze data generated by an existing study.
+Third, more suggestively, we find no evidence of a change in this relationship between the size of NIH grants and the quantity of resulting publications reporting the results of a clinical trial during our period of interest.
 [^35] We inspect this relationship by collecting records from the NIH RePORTER database, for years 2010 to 2019.
@@
 This is stable across years of data.
+Taken together, we view the trends in our data, when set against existing empirical evidence, as indicating that clinical trial publication has been stable over this time period.
 3.3. Why is the Quantity of Research Rising?
@@
 In this section, we examine various measures of publication composition and quality to shed light on the source of this increase.
-Our aim, in doing so, is to determine whether this measured increase corresponds to a “real” increase in the quantity of underlying research or whether it captures changes in the ease and frequency of publication.
-Evidence consistent with the latter suggests, then, that counts of scientific publications may provide little insight on questions concerning research productivity.
+Our aim is to determine the source of the categorization errors that yield measures such as those in Figure 2.1.
+In doing so, we provide suggestive evidence on whether measured increases in the quantity of scientific research correspond to “real” increases in the quantity of underlying research or whether they capture changes in the ease, frequency, and composition of publication.
 3.3.1. Publication Content
 We inspect the technical content of “non-clinical-trial” papers in three ways.
-We begin by hand-labeling random samples of 100 records, drawn from the beginning (publication year 2013) and end (publication year 2019) of our sample of papers that cite clinical trials.
+We hand-label random samples of 100 records, drawn from the beginning (publication year 2013) and end (publication year 2019) of our sample of papers that cite clinical trials.
 In this pooled set of 200 records, we identify one record that should have been in our sample of clinical trials.
-[^49] The erroneously labelled publication is PMID 23797691, which reports the results of a clinical trial studying a drug to treat HIV.
+[^36] The erroneously labeled publication is PMID 23797691, which reports the results of a clinical trial studying a drug to treat HIV.
 Note that this exercise suggests a false negative rate of 0.005, consistent with estimates reported previously.
@@
 Of course, in our small sample of records, it is difficult to draw authoritative conclusions about the role of growth in this type of literature.
-We collect measures of changes in the quantity of meta-analyses, based on the hypothesis that this may be a quantitatively important driver of the changes in our sample, in two ways.
+We collect two measures of changes in the quantity of meta-analyses, based on the hypothesis that this may be a quantitatively important driver of the changes in our sample.
 First, we implement a simple search of the abstracts of publications in this “citing” sample.
-We inspect the set of hand-labeled records, used as the “training data” for our fine-tuning procedure in Section 3.3, as many were flagged as out-of-sample because they were literature reviews or meta-analyses.
-From this set, we collect a list of keywords that were indicators when we hand-labeled these records that the publication was a meta-analysis.
-Our search terms are intentionally specific, as the aim of this procedure is to yield a plausible lower bound for any trend.
+To select keywords for this search, we inspect the set of hand-labeled records, used as the “training data” for our fine-tuning procedure in Section 2.4, as many were flagged as out-of-sample because they were literature reviews or meta-analyses.
 [^38] We searched for the following keywords: meta-analysis, metaanalysis, metaanalyses, systematic review, systematic reviews, systematically review, systematic search, review of published data, literature review, literature search, search of databases, review all literature, reviewed all literature, narrative review, systemic review.
 We also included any record that mentioned at least two of the following databases, which—based on our hand-labeling experience—are often referenced as part of the methods involved in a literature review or meta-analysis: MEDLINE, EMBASE, CINAHL, PubMed, Cochrane Central Register of Controlled Trials, BioMedCentral.
-More liberal search terms (e.g., “review”) risk capturing a large number of extraneous records.
-As expected, this yields a small number of records: in our three-year citing trials sample, there are 1,606 record in our citing sample that satisfy these criteria in 2013 and 3,372 in 2019, corresponding to a 109 percent increase.
+As expected, given the stringency of our search terms, this yields a small number of records: in our three-year citing trials sample, there are 1,606 records that satisfy these criteria in 2013 and 3,372 in 2019, corresponding to a 109 percent increase.
 Our second method generates—as the discussion in Section 2 suggests—what might be viewed as an upper bound on the quantity of meta-analyses and literature reviews in our data.
@@
 From these three procedures, we draw a suggestive conclusion: there has been a quantitatively important increase in the number of meta-analyses over our period of interest.
-This increase is on the order of 30 to 100 percent, but various measurement strategies yield alternative estimates.
+This increase is on the order of 30 to 100 percent, across measurement strategies.
 The specific quantity of meta-analyses is not of central interest for this paper, except insofar as it allows us to conclude that the quantity of medical research that summarizes existing papers—rather than producing new evidence—appears to be growing over time.
-This trend provides support for our decision to focus on clinical trials as a measure of productivity-relevant research and sheds some light on the source of documented increases in the quantity of research in other papers.
-As we note in Section 2, meta-analyses and similar forms of literature review can provide important information to physicians and can be vital in the diffusion of existing information, especially as the stock of knowledge grows (Jones,, 2009).
-Here, then, the distinction between those forms of scientific effort that are relevant to productivity metrics and those that are relevant to more expansive welfare metrics is especially important.
+To be clear, meta-analyses and similar forms of literature review provide important information to physicians and can be vital in the diffusion of existing information, especially as the stock of knowledge grows (Jones,, 2009).
+This change in composition, however, is relevant context when interpreting measured increases in the quantity of publications.
 3.3.2. Publication Geography
@@
 In our sample, the top four producers of published clinical trials, and papers citing clinical trials, are, in order, the United States, China, Germany, and Japan.
-Figure 4.4. Composition of Medical Research Across Countries
+Figure 3.4. Composition of Publications Across Countries
 Panel A: Clinical Trials Panel B: Papers Citing Clinical Trials
@@
 Since 2013, there has been a small increase in the quantity of other medical research published by authors in the United States.
-This small change is dwarfed by increases—on the order of 30-50 percent—in the quantity of other medical research from China and from countries outside of the top four producers of medical research.
+This small change is dwarfed by increases—on the order of 30-230 percent—in the quantity of other medical research from China and from countries outside of the top four producers of medical research.
 Institutional details are helpful in making sense of the differential patterns across these cuts of data.
@@
 [^43] https://datatopics.worldbank.org/world-development-indicators/the-world-by-income-and-region.html That is, this trend captures, in part, an increase in the quantity of diffusion-focused literature in high-income countries that have long had robust health infrastructure.
-Below, we establish that on several quality measures, these publications appear to be declining in value over time—consistent with the idea that some of this increase in production may be a response to changing scientific incentives for publication, rather than “real” invention or diffusion effort.
+Below, we establish that on several quality measures, these publications appear to be declining in value over time—consistent with the idea that some of this increase in production may be a response to changing scientific incentives for publication.
 3.3.3. Publication Quality
@@
 Citations are imperfect measures of the ex post quality of a scientific publication.
-Our data allow us to examine a variety of other measures that provide insight into the ex ante quality of a particular instance of research.
+Our data allow us to examine several other measures that provide suggestive insight into the ex ante quality of these publications.
 Unfortunately, shifting geographic trends in this sample and the low quality of measures of public funding for countries outside of the United States render it difficult to use a measure of public funding as a proxy for quality.
@@
 We collect facts about these groups separately.
-We begin by examining the productivity of senior scientists who authored papers in our citing sample.
+We begin by examining the productivity of lead scientists who authored papers in our citing sample.
 We collect all publications indexed in PubMed authored by lead scientists (last-authors) who write at least one paper in our citing sample.
@@
 For junior scientists, we focus on a measure of attrition: after authoring a paper in our citing sample, do we ever observe another publication written by that junior scientist?
+[^46] Hill and Stein, (2021) use a similar measure of attrition, when considering the impacts of losing a priority race for scientists’ careers.
 We collect all publication records in the Web of Science associated with these junior (first-listed) authors.
@@
 When interpreted in light of the citation distribution trends in Figure 3.5, we conclude that measured increases in the quantity of this “other” medical research may be accompanied by a decline in its average quality.
-It is precisely this shift that suggests that quantities of scientific publications in this sector, when not restricted to more precise categories of research such as clinical trials, may be a poor indicator of research productivity.
 4. Discussion
-Studies of innovation and productivity have long grappled with the prospect of diminishing returns to research (e.g., Merton,, 1935; Stafford,, 1952; Schmookler,, 1952, 1954, 1966; Griliches,, 1990, 1994; Bloom et al.,, 2020).
-Answers to the question at the center of this literature—is the cost of sustained technological progress increasing over time?—have been unsatisfying, as fundamental challenges in the measurement of technological innovation have limited our ability to do more than determine that available metrics do not yield conclusive answers.
-In one sense, this paper fits in this tradition, in that we find that, in the context of medical research, recent concerns about declining productivity are an artifact of long-standing measurement challenges.
-On the other hand, enabled by advances in generative AI, we introduce a new technique for annotating unstructured text that allows us to construct a proxy for medical research that, we argue, closely approximates the object of interest in an analysis of medical research productivity.
-We fine-tune an open-source large language model, which allows us to extract records of clinical trials from a corpus of scientific publications.
-The resulting census corrects for errors in previous measures of clinical trial research and approximates the quality of hand-labeled data.
+Studies of innovation and productivity have long grappled with the prospect of diminishing returns to research (e.g., Merton,, 1935; Stafford,, 1952; Schmookler,, 1952, 1954, 1966; Bloom et al.,, 2020).
+Answers to the question at the center of this literature—is the cost of sustained technological progress increasing over time?—have been unsatisfying, as fundamental challenges in the measurement of technological innovation have limited our ability to do more than determine that available metrics do not yield conclusive answers (Griliches,, 1990, 1994).
+In one sense, this paper fits in this tradition, in that we find that, in the context of pharmaceutical research, recent concerns about declining productivity are an artifact of confounded measurements of research quantities.
+On the other hand, enabled by advances in generative AI, we develop a procedure for labeling unstructured text that allows us to produce data with an accuracy, precision, and scale sufficient to make progress on the measurement challenges inherent to empirical studies of innovation.
+We apply this procedure to identify records of clinical trials from a corpus of scientific publications.
+The resulting census corrects errors in previous measures of clinical trial research.
 We show that trends in the quantity, quality, and composition of clinical trials have been stable since 2010.
-Our data allow us to contrast these trends with analogous patterns in the production of other forms of medical research.
+We contrast these trends with analogous patterns in the production of other forms of medical research.
 We find that, since 2010, the quantity of other forms of medical research has roughly doubled, a pattern consistent with estimates reported in other papers.
-On inspection, the increase in quantity includes a large increase in the number of papers authored by scientists in China and in the number of papers that synthesize existing literature (e.g., literature reviews and meta-analyses).
-Growth in these forms of research coincides with an average decline in various measures of publication quality.
-Taken together, we interpret these facts as providing no support for claims of declining research productivity, at least since 2010.
+On inspection, the increase in quantity is driven by a large increase in the number of papers authored by scientists in China and in the number of papers that synthesize existing literature (e.g., literature reviews and meta-analyses).
+Growth in these forms of research coincides with a decline in various measures of average publication quality.
+These facts provide evidence on the source of classification errors in existing measures of clinical trial production and, more generally, provide context on compositional trends relevant to studies that use publications as measures of research.
 These patterns are, in many ways, unsurprising.
@@
 A trend toward evidence-based decision-making in medicine has increased demand for articles that synthesize large bodies of research (Lohr,, 2004).
-Recent work documents a substantial increase in the quantity of meta-analyses and systematic reviews—summaries of existing clinical trials—driven by scientists outside of the United States (Ioannidis, 2016a, ).
+Recent work documents a substantial increase in the quantity of meta-analyses and systematic reviews—summaries of existing clinical trials—driven by scientists outside of the United States (Ioannidis,, 2016).
 In parallel, there has been a rapid increase in the number, and scale, of “mega-journals”—scientific journals published frequently, with a large and growing number of pages (Ioannidis et al.,, 2023).
@@
 Sites may struggle to find patients who meet these criteria, especially in small disease markets (Kolata,, 2017), and the costs of accruing large samples for certain studies can be prohibitive (Alsan et al.,, 2024).
-Existing work documents patterns that are not inconsistent with those in this paper.
-We differ, however, in interpretation.
-Like Park et al., (2023) and Bloom et al., (2020), we find a substantial increase in the size of the medical research ecosystem.
-Park et al., (2023) document that this increase in scale brings with it a decline in the average importance of published scientific articles.
-Technically, we find the same.
-As increases in the quantity of non-trial publications outpace (non-existent) changes in clinical trials, the average productivity of studies in Figure 3 mechanically decreases.
-We find, also, that increases in the quantity of non-trials are coincident with declines in various measures of publication quality.
-From the perspective of a productivity calculation, however, we argue that measures of average quality provide little insight.
-We cannot rule out, for example, that the same amount of research is being produced in these non-trial publications, but that what had previously been reported in a single paper is now spread across several distinct records.
-To be sure, these findings raise a host of other important questions about scientific research—about misaligned incentives, wasteful spending, and both productive and allocative efficiency.
-But on the issue of research productivity, these shifts provide very little insight.
-Our findings do not speak to changes in upstream, basic science inputs into clinical trials, or to changes in factor prices.
-As Section 3 discusses, in our view, there is very little credible evidence that might allow us to investigate changes on either dimension during this time period.
-We can note, only, that any shift in inputs or factor prices must be perfectly offset by a countervailing force, to account for the complete stability of clinical trial counts, citations, journal placements, and reliance on public funding in our data.
+Our findings focus on one aspect of the larger biomedical research ecosystem: investments in efforts to bring basic science discoveries to market, in the form of new medicines.
+Development of supplemental measures of medical research productivity, which capture the efficiency of basic scientific research, is a useful area for future research.
 Our findings, also, apply to roughly one decade.
@@
 From the perspective of contemporary public policy, however, trends from 2010 to the present are likely to be most relevant.
-We reconcile two stylized facts about research and inventive output in a manner that is both optimistic and pessimistic.
-On the one hand, we find that, in the most recent period, both the inputs and outputs of an “idea production function” (2.1) are constant across periods.
-This should alleviate the concern, in this sector, that “ideas are getting harder to find” (Bloom et al.,, 2020).
+Our revision of estimates of clinical trial production is both, from a policy perspective, optimistic and pessimistic.
+On the one hand, we find evidence consistent with a market in equilibrium: clinical trial production, on a variety of margins, appears stable.
+This directly calls into question the stylized fact of diminishing returns to drug development, which is viewed in some settings (e.g., Scannell et al., (2012) and Goldin et al., (2024)) as an empirical regularity on par with Moore’s Law for semiconductor development.
 On the other hand, it is worth remarking, briefly, on the implications of completely stable clinical trial trends.
@@
 National Institutes of Health between 1998 and 2003.
-Against these factors that seemingly forecast a period of rapid innovation, the two facts documented in this paper—stability in productivity-relevant research investments, but rising quantities of comparatively low-impact forms of research—highlight the importance of understanding frictions that affect the translation of basic scientific insights into medical technologies relevant to patients.
+Against these factors that seemingly forecast a period of rapid innovation, the facts documented in this paper—stability in the publication of clinical-trials, but rising quantities of other forms of translation research—highlight the importance of understanding frictions that affect the translation of basic scientific insights into medical technologies relevant to patients.
 Specifically, these patterns suggest that there is likely value in efforts to increase the productivity and decrease the costs of clinical trials, and in efforts to reconcile incentives in academic publishing with social value.
 A long-standing discussion in medicine and statistics, spurred by Altman, (1994), centers on the argument that “we need less research, better research, and research done for the right reasons.” Our data indicate that nearly half of all clinical trials published in our sample are never cited by leading medical journals, and roughly 15 percent are never cited at all.
-Outside of clinical trials, our findings suggest that the quantity of low-quality research might be growing over time.
-This heterogeneity is consistent with findings in Ioannidis, 2016b and van Zwet et al., (2023), who imply that there may exist trade-offs between quantity and data quality in this setting.
-If increasing the scale of research infrastructure is a policy objective, such efforts should invest in tools, like patient registries, that enable the production of high-quality, representative information (Alsan et al.,, 2024; Durvasula,, 2023).
-Supplemental Appendix to: Medical Research as a Productivity Indicator [^1] Date: August 3, 2024
+Supplemental Appendix to: Distilling Data from Large Language Models:
+An Application to Research Productivity Measurement [^1] Date: September 11, 2024
 $\star$ Stanford Department of Economics and Stanford Law School, maya.durvasula@stanford.edu
@@
 Maya M. Durvasula Sabri Eyuboglu David M. Ritzwoller
-Contents  appendix.Asubsection.A.1subsection.A.2subsubsection.A.2.1subsubsection.A.2.2subsubsection.A.2.3subsubsection.A.2.4subsubsection.A.2.5subsection.A.3subsection.A.4appendix.Bappendix.Csubsection.C.1subsection.C.2subsection.C.3subsection.C.4subsection.C.5subsection.C.6appendix.Dsubsection.D.1subsection.D.2subsection.D.3subsubsection.D.3.1subsubsection.D.3.2subsubsection.D.3.3appendix.E
+Contents  appendix.Asubsection.A.1subsection.A.2subsubsection.A.2.1subsubsection.A.2.2subsubsection.A.2.3subsubsection.A.2.4subsubsection.A.2.5subsection.A.3subsection.A.4appendix.Bappendix.Csubsection.C.1subsection.C.2subsection.C.3subsection.C.4appendix.Dsubsection.D.1subsection.D.2subsection.D.3subsubsection.D.3.1subsubsection.D.3.2subsubsection.D.3.3appendix.E
 Appendix A Publication Data
@@
 Adaptive trial; Clinical conference; Clinical study; Clinical trial; Clinical trial protocol; Clinical trial, Phase 1; Clinical trial, Phase 2; Clinical trial, Phase 3; Clinical trial, Phase 4; Comparative study; Controlled clinical trial; Equivalence trial; Evaluation study; Observational study; Pragmatic clinical trial; Randomized controlled trial; Twin study; Validation study.
-These categories are chosen to include all pubtypes with the potential to contain a publication satisfying the restriction enumerated in 3.1.
+These categories are chosen to include all pubtypes with the potential to contain a publication satisfying the restriction enumerated in 2.1.
 In particular, we follow Feldman et al., (2019), who use multiple pubtype fields to retrieve a sample of records.
@@
 Notes: Table A.1 reports the frequency and percentage of records indexed in PubMed that have been categorized by the NLM as falling into each of 18 categories.
-These categories are selected for their potential to contain a publication satisfying the restriction enumerated in 3.1.
+These categories are selected for their potential to contain a publication satisfying the restriction enumerated in 2.1.
 A.2.2. Clinical Trial Registry Identifiers
@@
 Of the roughly 150,000 records in our census, approximately 100,000 could be identified using NLM tags for “clinical trial” (and its variants) and “randomized controlled trial.” Observe, however, that use of such flags captures a much larger number of records than one would intend to include.
-Finally, Table A.4 suggests that although various text-based flags that may indicate a record is a clinical trial, in the sense of 3.1, are correlated, there is less overlap than one might expect.
+Finally, Table A.4 suggests that although various text-based flags that may indicate a record is a clinical trial, in the sense of 2.1, are correlated, there is less overlap than one might expect.
 Table A.4. Overlap of Publication Attributes
@@
 In this appendix, we detail the construction of the sample of clinical trials studied in Section 3.
-We begin in Section C.1 by describing the process we used to hand-label a sample of publications according to whether they satisfy the restrictions enumerated in 3.1.
+We begin in Section C.1 by describing the process we used to hand-label a sample of publications according to whether they satisfy the restrictions enumerated in 2.1.
 In Section C.2, we document poor performance of several approaches for classifying publications according to whether they belong in our sample based on standard machine learning methods.
 Section C.3 overviews the process of constructing prompts that perform well when used to query proprietary language models.
-Details associated with fine-tuning an open-source language model are given in Section C.4.
-Additional measurements of the performance of our model are given in Section C.5.
 C.1. Hand-Labeling
@@
 The labeller can select that the abstract satisfies our sample criteria, or else can indicate a reason that the abstract should be excluded.
-The buttons for exclusion mirror the inclusion and exclusion criteria in 3.1, using short-hand convenient for labellers.
+The buttons for exclusion mirror the inclusion and exclusion criteria in 2.1, using short-hand convenient for labellers.
 We exclude records that report, only, the protocol for a clinical trial without results (“protocol”), that report the results of an observational (“observational”) or retrospective (“retrospective”) study, that report the results of a study involving animals (“animal”), that do not include human subjects (“no human subjects”), [^55] In practice, “animal” and “no human subjects” catch different sets of records.
@@
 Notes: Figure C.1 displays a screenshot of the application we developed to hand-label abstracts.
-Each abstract was labelled according to whether it satisfied the sample criteria enumerated in 3.1.
+Each abstract was labelled according to whether it satisfied the sample criteria enumerated in 2.1.
 If a publication did not satisfy our sample restrictions, we stipulated a reason.
@@
 If the abstract references that the publication is summarizing existing studies, or searching a database for records and collecting their findings, we categorize this as an implicit error.
-We devise error types based on our inclusion/exclusion criteria, in 3.1, and select sub-types based on common categories of errors.
+We devise error types based on our inclusion/exclusion criteria, in 2.1, and select sub-types based on common categories of errors.
 We revise each class of prompt based on these findings.
@@
 We categorized errors by Type and Sub-Type.
-Types are drawn from the exclusion restrictions in 3.1.
+Types are drawn from the exclusion restrictions in 2.1.
 Sub-types describe consistent characteristics of publications that resulted in such an error.
 We report counts of errors, by model and prompt, of each Type and Sub-type.
-C.4. Fine-tuning
-This appendix gives technical details related to fine-tuning open-source large language models with noisy labels extracted from GPT-3.5 and GPT-4.
-We fine-tune pre-trained BERT models from two different architecture classes: (1) BigBird and (2) BioMedBERT (Gu et al.,, 2021; Zaheer et al.,, 2021).
-Both models us medium-scale Transformer architectures, i.e., between 100 and 300 million parameters, pre-trained with the masked-language modeling (MLM) objective (Vaswani et al.,, 2023; Devlin et al.,, 2019).
-The models differ in their architectural details and pre-training corpora.
-BigBird uses a sparse attention mechanism to reduce the computational cost of processing long text sequences.
-The computational complexity of regular attention is quadratic in sequence length, while BigBird reduces this to linear complexity.
-This enables efficient handling of long documents up to 4,096 tokens.
-In contrast, standard BERT can only process 512 tokens.
-BigBird is useful because abstracts of clinical trials regularly exceed 512 tokens.
-Prior to pre-training, the model was warm-started from the RoBERTa checkpoint (Liu et al.,, 2019) and then pre-trained on the standard BERT corpus consisting of the Books (Zhu et al.,, 2015), CC-News (Guu et al.,, 2020), Stories (Trinh and Le,, 2019), and Wikipedia (Wikimedia Foundation,, 2023) datasets.
-We evaluate two model sizes: BigBird Base (125 million parameters) and BigBird Large (355 million parameters).
-BioMedBert uses a domain-specific pretraining corpus sourced from articles on PubMed Gu et al., (2021).
-It uses the same model architecture as RoBERTa, but a different, domain-specific token vocabulary optimized for PubMed articles.
-Like BigBird, we evaluate two model sizes: BioMedBert Base (125 million parameters) and BioMedBert Large (355 million parameters).
-We fine-tune the pre-trained architectures to classify abstracts according to the inclusion and exclusion criteria enumerated in 3.1.
-We replace the language modeling classification head with a binary classification head consisting of a single linear layer and softmax activation.
-We fine-tune the full model (i.e.
-not just the classification head) with cross-entropy loss.
-We use the Adam optimizer with learning rate $1\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$ (determined via hyperparameter sweep).
-We use a maximum sequence length of $4{,}096$ tokens.
-If an abstract does not fit into the maximum sequence length provided by a model, the abstract is truncated.
-All models were trained on a single $V100$ GPU with 16GB of HBM.
-C.5. Performance Assessment
-In this appendix, we assess the performance of various fine-tuned language models for classifying whether a publication satisfies the restrictions enumerated in 3.1.
-Figure C.2 displays estimates of the receiver operating characteristic of several fine-tuned large language models.
-Here, we consider only the performance of the RoBERTa base model.
-The top and bottom panels are trained on noisy labels extracted from GPT-3.5 and GPT-4, respectively.
-We vary the quantity of noisy labels used to train the language model.
-Additionally, in the bottom panel, we display the performance of a fine-tuned model trained with the 1000 hand-labelled observations in the training dataset.
-The performance of each language model is measured in the testing dataset.
-Figure C.2. Receiver Operating Characteristic by Data Size and Origin
-Notes: Figure C.2 displays estimates of the the receiver operating characteristic of several fine-tuned language models.
-The models differ in terms of the source of their training data.
-Solid lines display the performance of models trained on noisy labels extracted from either GPT-3.5 or GPT-4.
-The dotted line displays the performance of a model trained on the hand-labeled data in the training dataset.
-Estimates of the true positive rate and false positive rate of GPT-3.5 and GPT-4 are indicated with black dots.
-The fine-tuned models trained with noisy labels extracted from GPT-4 significantly outperform the models trained with noisy labels extracted from GPT-3.5 or with the hand-labels.
-There appears to be a threshold, where performance dramatically improves, at around 8,000 training labels.
-The models displayed in Figure 3.3 were each obtained by using all 64,000 training labels.
-C.6. Final Model: Error Analysis
+C.4. Error Analysis
 The Conservative model incorrectly labels 27 papers.
@@
 These clear errors include publications that explicitly report the results of observational studies (one error), that re-analyze existing data (eight errors), and that report literature reviews (one error).
-This set also includes clinical trials that satisfy all criteria in 3.1, except the treatment being studied is not a drug (three errors).
-In 14 cases, however, errors are associated with records that are difficult for human labellers to categorize, either because the content of the publication does not fit neatly into the inclusion and exclusion criteria implied by 3.1 or because the publication is written in a way that makes it difficult to determine details of the study.
+This set also includes clinical trials that satisfy all criteria in 2.1, except the treatment being studied is not a drug (three errors).
+In 14 cases, however, errors are associated with records that are difficult for human labellers to categorize, either because the content of the publication does not fit neatly into the inclusion and exclusion criteria implied by 2.1 or because the publication is written in a way that makes it difficult to determine details of the study.
 PubMed record 27880726 is illustrative.
@@
 We constructed a list of phrases and keywords that consistently indicated, to a human labeler, that a record is a meta-analysis or review article.
-This list is reproduced in Footnote 51.
+This list is reproduced in Footnote 38.
 Using this list of search terms, we search the abstracts of all records in our citing sample.
-We flag any record with any of the keywords listed in Footnote 51 as a meta-analysis and flag any record that mentions at least two of the databases listed in Footnote 51 as well.
+We flag any record with any of the keywords listed in Footnote 38 as a meta-analysis and flag any record that mentions at least two of the databases listed in Footnote 38 as well.
 The objective of this exercise is to yield a conservative estimate of the change in the quantity of meta-analyses over our period of interest.

### v5

@@
 We develop a method for assigning high-quality labels to unstructured text.
-The method is based on fine-tuning an open-source language model with data extracted from a proprietary large language model.
+This method is based on fine-tuning an efficient, open-source language model with data extracted from a large, proprietary language model.
 We apply this method to construct a census of published clinical trials.
-We revisit a literature that contends that pharmaceutical research productivity is declining, based on measured increases in the quantity of clinical trials, which outpace trends in output.
-In our data, the quantity and composition of trials are stable since 2010.
-Previous measurements are an artifact of biases driven by shifts in the composition of other forms of research.
+With these data, we revisit a literature that contends that pharmaceutical sector productivity is declining.
+Central to this conclusion are measurements of substantial increases in the quantity of clinical trials over time, unmatched by trends in measures of output.
+In our data, the quantity, quality, and composition of clinical trials are stable since 2010.
+We show that previous measurements are an artifact of biases introduced by shifts in the composition of other forms of research.
 Keywords: Large Language Models, Medical Research Productivity, Clinical Trials
 JEL: C81, O32
-Perhaps the most serious task facing empirical work in the area of “technological change” and “invention and innovation” is the construction and interpretation of measures (indices) of advances in knowledge.
-— Pakes and Griliches,, 1980
 1. Introduction
-Considerable progress in the measurement of innovation has been made in the half-century since the publication of Pakes and Griliches, (1980).
-This progress has been enabled in large part by the proliferation of large-scale corpora of patents and scientific publications.
-Yet measurement remains an active constraint for empirical research (Bryan and Williams,, 2021).
-Patents and publications are attractive proxies for innovation, as they necessarily relate to instances of invention and research.
-Their utility is limited, however, by heterogeneity in their technical content and value (see e.g., Pakes,, 1986; Griliches,, 1990, 1994).
-Consequently, trends in counts of documents are often confounded by trends in their composition (Lerner and Seru,, 2022).
-Correction of this confounding can be viewed as a classification problem: how accurately can unstructured text be mapped to specific quantities of interest?
-[^1] The most prominent example of such a classification effort is the construction of the National Bureau of Economic Research (NBER) Patent Data, which, among other things, maps patent records to technology classes (Hall et al.,, 2001).
-Earlier, Scherer, (1984) classified, by hand, over 15,000 patents according to their technological category.
-In this paper, we demonstrate that large language models can facilitate classification tasks of this kind, achieving an accuracy and precision comparable to hand-made labels at a research-relevant scale and budget.
-We develop a procedure for constructing task-specific large language models, particularly suited to data construction problems that involve classification.
+For nearly thirty years, policy discussions about the pharmaceutical industry have been colored by the concern that sectoral productivity is in a state of decline (see e.g., Engelberg,, 1998; Harris,, 2011).
+A series of influential analyses, drawing on industry reports and U.S.
+Food and Drug Administration records, suggest that the resources required to bring a new drug to market have been increasing substantially year-on-year.
+Scannell et al., (2012), for example, estimate that the productivity of the pharmaceutical sector has halved roughly every decade since 1950; Ruffolo, (2006), Pammolli et al., (2011), Myers and Pauly, (2019), and Bloom et al., (2020) arrive at similar conclusions.
+The policy consequences of these findings have been substantial.
+Proposals that would restructure public subsidies administered by the National Institutes for Health (NIH) (e.g., Executive Office of the President,, 2019; Moss et al.,, 2020; The Economist,, 2022), relax the regulatory standards for approval of new medicines by the U.S.
+Food and Drug Administration (FDA) (e.g., Roy,, 2021; Chertman,, 2023; Graboyes and Reimschisel,, 2017), and prevent the Centers for Medicare and Medicaid Services (CMS) from negotiating drug prices (e.g., York,, 2023; Long,, 2023; Inflation Reduction Act,, 2022) are premised, in no small part, on the concern that existing policy is contributing to this pattern of decline.
+[^1] In recent years, discussions of declining medical and pharmaceutical research productivity have been especially extensive in the popular press.
+Collison and Nielsen, (2018), Broad, (2023), Thompson, (2021), and Piper, (2023), for example, interpret trends in pharmaceutical research productivity, alongside patterns collected from other fields, as a ‘warning sign” for the future of technological progress (Piper,, 2023).
+For the same reasons, an architect of the 1984 Hatch Waxman Act, which established the modern system for pharmaceutical regulation in the United States, called for an overhaul of that regulatory system just fifteen years after its inception (Engelberg,, 1998).
+These patterns also spurred the 2011 establishment of the National Center for Translational Sciences, a federal agency with a roughly one billion dollar annual budget and a mandate to speed up the commercialization of life-saving medicines—on the grounds that, left to private firms, this process had become too inefficient.
+[^2] The fiscal year 2025 budget request for NCATS is \$926.1 million.
+For details on its founding, see Harris, (2011).
+For details on its budget, see https://ncats.nih.gov/about/budget.
+Though this declining pharmaceutical sector productivity is often characterized as an empirical regularity on par with Moore’s Law in computing, its empirical support is quite thin.
+[^3] Moore’s Law is the observation that the number of transistors on an integrated circuit will double every two years with minimal rise in cost.
+See Moore, (1965).
+Scannell et al., (2012) coined “Eroom’s Law” (Moore, spelled backwards) to describe pharmaceutical sector productivity.
+On some measures, the quantity of resources devoted toward the commercialization of new medicines appears to be sharply increasing over time.
+Given that the number of innovative drugs approved by the FDA has been roughly constant, year-on-year, since 1990, this divergence between inputs and outputs generates the conclusion of declining productivity.
+This paper constructs new data on pharmaceutical sector inputs.
+We develop a method for assigning high-quality labels to unstructured text.
+This method is based on fine-tuning an efficient, open-source language model with data extracted from a large, proprietary language model.
+We apply this method to construct a new census of clinical trials disclosed in scientific publications.
+We document that the quantity, quality, and composition of published clinical trials have remained remarkably stable year-over-year.
+On the basis of these facts, we find that there is insufficient evidence to reject the null hypothesis that pharmaceutical sector productivity has remained stable since 2010.
+We show that reported increases in the quantity of published clinical trials, offered as recent evidence of declining productivity in Bloom et al., (2020), result from imprecise classification methods.
+These methods misattribute shifts in the composition of biomedical research to trends in clinical trial production.
+When our data are combined with evidence from Sertkaya et al., (2024)—that clinical trial costs, and broader drug development expenses, have remained roughly constant since 2000—our findings leave little in the way of evidence for the conventional wisdom of a persistent decline in pharmaceutical sector productivity.
+This paper makes two primary contributions.
+First, we revise a fact central to debates on biomedical productivity.
+Second, we demonstrate how advances in generative AI can address longstanding challenges in measuring innovation.
+The paper is, thus, organized in two parts.
+In the first part, we introduce a procedure for constructing task-specific large language models, particularly suited to data construction problems that involve classification.
+[^4] We include, as a supplemental appendix to this paper, a more detailed summary of this procedure, which may be useful for future work.
 The class of methods that we consider has two steps.
@@
 [^6] Here, we are referring to classes of open-source language models such as RoBERTa, based on Google’s BERT model, among others.
-This process is an instance of model distillation, which is a growing area of methodological research in computer science (Taori et al.,, 2023; Xu et al.,, 2024).
+This process is an instance of model distillation, a growing area of methodological research in computer science (Taori et al.,, 2023; Xu et al.,, 2024).
 This procedure allows researchers to capture the benefits of complex, frontier models, while retaining the resource efficiency and transparency of lighter-weight open-source models.
-The primary contribution of this paper is the application of this procedure in a context that is emblematic of the econometric challenges faced by studies of innovation: measurement of trends in clinical trial production.
+We apply this method in stages.
+Our objective is to identify all records in the National Library of Medicine’s PubMed/MEDLINE database, from 2010 forward, that disclose the results of a clinical trial that studied the effects of a medicine in human subjects.
+We aim to classify records based only on the contents of the publication’s abstract, with an accuracy close to that of a human labeler.
+We devise an interface to iteratively revise model prompts.
+With the highest performing prompts, we use proprietary language models, OpenAI’s GPT-3.5 and GPT-4, to classify a large number of records (Achiam et al.,, 2023; Bubeck et al.,, 2023).
+These noisy labels are used as training data to fine-tune a set of open-source models.
+An ensemble of the fine-tuned models matches the performance of the best proprietary model, with false positive and false negative rates below 5 percent, and produces a sample of approximately 150,000 publications that report the results of clinical trials from 2010 to 2022.
+The second part of this paper interprets trends in these data, in light of existing evidence.
+We begin by establishing that trends in clinical trials, as measured in publication data, are central to conclusions of declining biomedical and pharmaceutical productivity.
 We show that standard methods for data classification identify clinical trials in publication data imprecisely.
 These imprecise measurements indicate, spuriously, that the quantity of clinical trials has increased sharply since 2010.
-In our data, the quantity, quality, and composition of clinical trials have been stable year-on-year.
-When considered alongside recent evidence on the stability of drug development costs during this period (Sertkaya et al.,, 2024), these new measurements of clinical trial production call into question the conclusions of an influential literature that finds diminishing returns to pharmaceutical research (Scannell et al.,, 2012; DiMasi et al.,, 2016; Bloom et al.,, 2020).
-[^4] Goldin et al., (2024) survey evidence related to declining productivity across the economy and note that, in the pharmaceutical sector, evidence for declining productivity primarily relies on measures of total numbers of clinical trials and estimates of R&D expenditures.
-The conclusions of this literature are based on measures that suggest that the quantity of clinical trials is increasing substantially, in a way unmatched by trends in measures of outputs.
-In light of our new data, we find that there is little rigorous empirical support for falling pharmaceutical research productivity in the most recent period.
-This finding is directly relevant to a set of active policy debates, which were initiated, in no small part, in response to concerns that the production of innovative output in this sector is slowing down.
-These debates include proposals that would restructure public subsidies administered by the National Institutes for Health (NIH) (e.g., Executive Office of the President,, 2019; Moss et al.,, 2020; The Economist,, 2022), relax the regulatory standards for approval of new medicines by the U.S.
-Food and Drug Administration (FDA) (e.g., Roy,, 2021; Chertman,, 2023; Graboyes and Reimschisel,, 2017), and prevent the Centers for Medicare and Medicaid Services (CMS) from negotiating drug prices (e.g., York,, 2023; Long,, 2023; Inflation Reduction Act,, 2022).
-Clinical trials are an especially attractive measure of research in medicine (Bryan and Williams,, 2021).
-In the pharmaceutical industry, clinical trials are discrete, regulated experiments that generate useful statistical information about new and existing medicines (Chavez-MacGregor and Giordano,, 2016).
-Trials top the hierarchy of medical evidence employed by firms, regulators, physicians, and patients (Jones and Podolsky,, 2015).
-Since the early 2000s, clinical trials have been reported in a consistent, standardized format across scientific publications and, with less consistency, in administrative databases (DeVito et al.,, 2020; DeVito and Goldacre,, 2021).
-Trials have, thus, been widely used to measure relative and absolute differences in the composition of research investment (see Finkelstein, (2004) and Budish et al., (2015) for two influential examples).
-Studies of information production and disclosure have examined the substantive content of trials (see e.g., Alsan et al., (2024), Oostrom, (2024), and Kao and Oostrom, (2024)).
-Recently, and prominently, measures of clinical trial quantities have been used to support arguments that medical and pharmaceutical research productivity are declining over time (Scannell et al.,, 2012; Bloom et al.,, 2020).
-Nevertheless, it is remarkably difficult to generate accurate measures of trends in clinical trial production.
-Administrative databases—such as ClinicalTrials.gov—have experienced large shifts in compliance with reporting regulations over time, which make it difficult to distinguish genuine increases in research effort from greater adherence to reporting requirements (see e.g., DeVito et al.,, 2020).
-Proprietary databases (e.g., Cortellis,, 2024) take as an input the contents of publicly-maintained registries and, thus, are confounded by the same shifts in compliance.
-Given these limitations, scientific publication data have served as a natural alternative or complement in recent empirical studies (e.g., Lichtenberg,, 2018; Bloom et al.,, 2020; Kao,, 2024; Guzman et al.,, 2024).
-With unstructured publication text data, the challenge becomes, again, one of classification: of the millions of records in large-scale publication databases, which disclose the results of a clinical trial?
-Manually identifying clinical trial records at a scale sufficient to study aggregate trends is infeasible.
-[^5] A small number of papers curate select samples of clinical trial data, for specific research questions.
-For example, Oostrom, (2024) collects records of psychiatric clinical trials as identified in prominent meta-analyses for a study of the relationship between sources of clinical trial funding and reported drug efficacy.
-We show that standard methods—based on publication metadata and machine learning tools—perform poorly.
-In particular, we show that keyword-based searches, including methods employed by Bloom et al., (2020), are very imprecise and generate spurious measurements of increases in the quantity of clinical trials.
-[^6] Specifically, Bloom et al., (2020) categorize many records as clinical trials that discuss trials (e.g., meta-analyses and literature reviews).
-We find that changes in the scientific publishing landscape, discussed in Section 4, increase the size of this categorization error over time.
-Specialized classifiers developed for use in the health services literature, as in Thomas et al., (2021), similarly lack the precision necessary to recover meaningful trends.
-Instead, we leverage recent advances in generative AI, which enable data extraction from unstructured text at scale.
-Our objective is to identify all records in the National Library of Medicine’s PubMed / MEDLINE database, from 2010 forward, that disclose the results of a clinical trial that studied the effects of a medicine in human subjects.
-We aim to classify records based only on the contents of the publication’s abstract, with an accuracy close to that of a human labeler.
-[^7] Many records in PubMed / MEDLINE have full-texts available for download, via the PubMed Central database.
-However, these full-texts are subject to copyright restrictions.
-Roughly one quarter of records in PubMed—$8$ million of the $34$ million total records—have full-texts that are available under licenses that would permit this type of research use (the PubMed Central Open Access Subset, at https://www.ncbi.nlm.nih.gov/pmc/tools/openftlist/).
-Thus, we restrict our focus to abstracts, for which no such restrictions apply.
-We proceed in stages.
-Given the abstract of a scientific publication, we prompt two proprietary language models, OpenAI’s GPT-3.5 and GPT-4 (Ouyang et al.,, 2022; Achiam et al.,, 2023; Bubeck et al.,, 2023), to label whether the record satisfies our desired sample restrictions.
-We find considerable variation in performance with seemingly small differences in prompt text.
-We iteratively revise prompt text until small changes yield no improvement in performance.
-GPT-4 performs quite well off-the-shelf.
-The performance of GPT-3.5 is unsuitable for our application.
-However, use of GPT-4 at scale—to classify millions of records—is computationally and financially impractical.
-We classify a moderate number of abstracts with the best-performing prompts.
-The labels extracted from the proprietary language models are used as training data to fine-tune a set of open-source models.
-Our preferred sample is constructed using an ensemble of fine-tuned models, which matches the performance of the best proprietary model.
-False positive and false negative rates are below 5 percent.
-When this model is used to label the full sample of PubMed records, it yields a sample of approximately 150,000 publications that report the results of clinical trials from 2010 to 2022.
-Three technical features of this fine-tuning process are worth highlighting.
-First, we document a phase-transition in model quality in the number of training labels used.
-Put differently, we observe a threshold in the number of training labels used in fine-tuning, beyond which the performance of the fine-tuned model increases dramatically.
-Second, models fine-tuned with labels extracted from GPT 3.5 exceed the performance of GPT 3.5.
-Third, models fine-tuned with labels extracted from GPT-4 match the performance of GPT-4 itself.
-That is, we show that researchers can distill the full benefits of large proprietary models into light-weight, resource-efficient models that can be realistically applied to standard data construction tasks.
-The resulting data allow us to document two sets of facts about the quantity of medical research over time.
-First, we find that the quantity of clinical trials has been effectively constant, year-on-year, across our period of interest.
+With improved classification, we find that the quantity, quality, and composition of clinical trials have been remarkably stable over this period.
 We collect a set of ex-ante and ex-post quality measures, derived from scientific publication data.
@@
 Nearly 15 percent are never cited by any other scientific publication.
-However, we find that the distribution across clinical trials of measures of quality, importance, and geography is unchanging.
-Second, our data allow us to shed light on the source of the classification errors that drive the trends documented in Bloom et al., (2020).
-[^8] These data are relevant outside of this particular context, for studies of medical innovation and for studies in the health services literature that use clinical trial publications to construct prescribing guidelines for physician practice.
-Identifying trial records in publication data has been a challenge for this health services literature, where existing best practices have both high true and false positive rates.
-See Thomas et al., (2021).
-As similar trends in the total quantity of scientific publications have been documented elsewhere, including in Park et al., (2023) and Chu and Evans, (2021), these insights may be more generally applicable to discussions about how large increases in the quantity of publications should be interpreted.
-We contrast trends in clinical trial production with trends in the production of medical research that is similar—conceptually and in its textual presentation in scientific publications—to clinical trials.
-In our preferred version, this set of other medical research is defined as the set of papers that cite clinical trials.
-[^9] In Section 4, we inspect the contents of this set and document that nearly all records fall into one of three categories: observational studies, preclinical studies, and meta-analyses/literature reviews.
-While the quantity of clinical trials is constant over this period, the quantity of this other non-trial research has increased by a factor of two.
-Three facts about shifts in the composition and quality of these publications provide insight into the source of measured increases in research quantity.
-First, the large increase in the total number of non-trial papers in our sample is largely driven by researchers in China.
-Second, this rise in quantity coincides with a decrease (on available measures) in the average quality of publications.
-Third, there is a substantial increase in the quantity of meta-analyses and literature reviews.
-The main conclusion that emerges from this analysis is that errors in the classification of scientific publications can yield spurious findings about trends in research effort.
+However, we find that the distribution across clinical trials of measures of quality, importance, and geography is unchanging over time.
+Our data offer more suggestive evidence on a second piece of evidence cited in support of declining research productivity—that the quantity of scientific publications is increasing sharply over time.
+We document evidence of large compositional shifts in other biomedical publications, including a shift in the geographic composition of research and a large increase in the number of articles that summarize existing literature.
+In line with the arguments of a long-standing literature on similar shifts in patent data (e.g., Griliches,, 1990, 1994), these patterns cast doubt on whether simple counts of scientific publications are stable, informative measures of research effort.
+These insights may be more generally applicable to discussions about how large increases in the quantity of publications—as documented in Park et al., (2023) and Chu and Evans, (2021)—ought to be interpreted.
 1.1. Related Literature
 We draw on, and contribute to, a long literature on the measurement of invention.
-Decades of work have focused, in particular, on the development of proxies for the outputs of technological innovation.
-An especially prominent approach, due to Hall et al., (2000), uses citation-weighted counts of patents.
-Our focus, by contrast, is on the difficulties associated with measuring the inputs of innovation, which can similarly bias estimates of macroeconomic trends.
-The new methods developed in this paper yield measures that call into question a fact at the center of debates involving health, innovation, and fiscal policy.
-Discussions of declining medical and pharmaceutical research productivity have been especially extensive in the popular press.
-Collison and Nielsen, (2018), Broad, (2023), Thompson, (2021), and Piper, (2023), for example, interpret trends in pharmaceutical research productivity, alongside patterns collected from other fields, as a “warning sign” for the future of technological progress (Piper,, 2023).
+This paper considers, in effect, a binary classification task, of a kind that has long been a challenge for studies of innovation.
+The proliferation of large-scale corpora of patents and scientific publications has, seemingly, expanded opportunities to measure instances of research and invention (Bryan and Williams,, 2021).
+Patents and publications are attractive proxies for innovation, as they necessarily relate to instances of invention and research.
+Their utility is limited, however, by heterogeneity in their technical content and value (see e.g., Pakes,, 1986; Griliches,, 1990, 1994).
+Consequently, trends in counts of documents are often confounded by trends in their composition (Lerner and Seru,, 2022).
+Large-scale classification efforts have focused on mapping pieces of unstructured text to economic quantities of interest.
+[^7] The most prominent example of such a classification effort is the construction of the National Bureau of Economic Research (NBER) Patent Data, which, among other things, maps patent records to technology classes (Hall et al.,, 2001).
+Earlier, Scherer, (1984) classified, by hand, over 15,000 patents according to their technological category.
+In this paper, we demonstrate that large language models can facilitate classification tasks of this kind, achieving an accuracy and precision comparable to hand-made labels at a research-relevant scale and budget.
+More specifically, within studies of medical innovation, we contribute to a body of work that uses clinical trials as a measure of research investment.
+Trials—discrete, regulated experiments that generate structured statistical information on new and existing health technologies (Chavez-MacGregor and Giordano,, 2016)—are an especially attractive proxy for the level and composition of medical research.
+Since the early 2000s, clinical trials have been reported in a consistent, standardized format across scientific publications and, with less consistency, in administrative databases (DeVito et al.,, 2020; DeVito and Goldacre,, 2021).
+Trials have, thus, been widely used to measure relative and absolute differences in the composition of research investment (see e.g., Finkelstein,, 2004; Yin,, 2008; Budish et al.,, 2015).
+Recent studies of information production and disclosure have examined the substantive content of trials (see e.g., Alsan et al., (2024), Oostrom, (2024), and Kao and Oostrom, (2024)).
+Given limitations in publicly-available administrative trial registries (Zarin et al.,, 2017; DeVito et al.,, 2020; DeVito and Goldacre,, 2021), recent empirical studies in economics have used trial data extracted from scientific publications (e.g., Lichtenberg,, 2018; Bloom et al.,, 2020; Kao,, 2024; Guzman et al.,, 2024).
+The challenge, for each, has been again one of classification: of the millions of records in large-scale publication databases, which disclose the results of a clinical trial?
+Manually identifying clinical trial records at a scale sufficient to study aggregate trends is infeasible.
+[^8] A small number of papers curate select samples of clinical trial data, for specific research questions.
+For example, Oostrom, (2024) collects records of psychiatric clinical trials as identified in prominent meta-analyses for a study of the relationship between sources of clinical trial funding and reported drug efficacy.
+We show that standard methods for classification—based on publication data and machine learning tools—perform poorly.
+In particular, we show that keyword-based searches, including methods used in a productivity analysis in Bloom et al., (2020), are very imprecise and generate spurious measurements of increases in the quantity of clinical trials.
+[^9] Specifically, Bloom et al., (2020) categorize many records as clinical trials that discuss trials (e.g., meta-analyses and literature reviews).
+We show, in Section 4, that changes in the scientific publishing landscape increase the size of this categorization error over time.
+Specialized classifiers developed for use in the health services literature, as in Thomas et al., (2021), similarly lack the precision necessary to recover meaningful trends.
+We introduce both a new method and new data that address this problem.
 The idea that patterns like those considered in this paper—rising research quantities and stagnant outcomes—signal the “exhaustion of inventive opportunities” (Griliches,, 1994) is old and closely tied to the development of methods for the measurement of innovation.
@@
 We show that advances in data construction technology, if deployed carefully, can allow researchers to resolve some long-standing data construction challenges.
-A nascent literature investigates the use of large language models in economics (Dell,, 2024; Korinek,, 2023).
+A nascent literature investigates the use of large language models in economics (Dell,, 2024; Korinek,, 2023; Ludwig et al.,, 2025).
 We join a small set of papers—notably, Bartik et al., (2023) and Dell, (2024)—that use generative AI as a tool to construct high-quality economic data.
@@
 We highlight a set of steps—in validation, prompt design, and fine-tuning of open source models—that yield high-quality data at comparatively low cost.
-2. Identifying Clinical Trials
+2. Counting Clinical Trials
 We construct a new census of clinical trials disclosed in scientific publications.
@@
 Declining pharmaceutical research productivity has been a topic of policy concern since the late 1990s.
-In the past thirty years, discussions have largely centered on two sets of empirical patterns, which Cockburn, (2006) summarizes as follows: “More and more money is being invested in R&D, but the rate at which new drugs are introduced is failing to keep pace.” R&D, in this setting, refers to a particular type of biomedical research: investments in preclinical and, especially, clinical studies intended to translate discoveries made in basic science into commercialized medicines suitable for use in human patients.
-That inputs per new drug appear to be rising sharply over time is viewed, generally, as an established fact in the empirical productivity literature (Bloom et al.,, 2020; Goldin et al.,, 2024).
-Measures of productivity in this setting are somewhat crude.
-The measure of outputs in the most influential articulation of this phenomenon, by Scannell et al., (2012), is the number of “new molecular entities”—innovative drugs—approved by the U.S.
+In the past thirty years, discussions have largely centered on two sets of empirical patterns, which Cockburn, (2006) summarizes as follows: “More and more money is being invested in R&D, but the rate at which new drugs are introduced is failing to keep pace.” In these exercises, the empirical object of interest is a measure of investments in preclinical and, especially, clinical studies intended to translate discoveries made in basic science into commercialized medicines suitable for use in human patients.
+The “puzzle” as characterized by Cockburn, (2006) and others centers on the apparent decline in the efficiency of pharmaceutical firms tasked with translational research.
+Hypotheses for the decline are wide-ranging.
+Ruffolo, (2006) emphasizes shifts in managerial incentives.
+Cockburn, (2006) highlights “vertical disintegration” of the pharmaceutical industry.
+Scannell et al., (2012) points to rising regulatory thresholds.
+Myers and Pauly, (2019) and Bloom et al., (2020) suggest that the “low-hanging fruit”—high-efficacy, easy-to-test drug candidates—have already been brought to market.
+Though the candidate mechanisms in these discussions differ, declining productivity in the pharmaceutical sector is treated as a fact (Bloom et al.,, 2020; Goldin et al.,, 2024).
+[^11] Scannell et al., (2012) coin the term “Eroom’s Law”—Moore’s Law in computing, spelled backwards—to described this trend.
+Policy analyses and academic surveys (e.g., Goldin et al., (2024)) have characterized both Eroom’s Law and Moore’s Law as well-supported stylized facts.
+The measures of industry productivity central to these debates are somewhat crude.
+For the pharmaceutical industry, a standard measure of output is the number of “new molecular entities”—innovative drugs—approved each year by the U.S.
 Food and Drug Administration.
@@
 In practice, this means that every active element of the drug (e.g., acetaminophen and caffeine in the commonly used over-the-counter migraine medicine Excedrin) must be novel.
-As Bloom et al., (2020) note, that this count has been stagnant or declining is viewed as an established fact.
+This is the same output measure used in discussions of changes in federal research policy.
+See, for example, Harris, (2011).
+As Bloom et al., (2020) note, that this count has been stagnant or declining over time is viewed as an established fact.
 [^13] In data on new molecular entities approved by the FDA between 2005 and 2021, we observe a small increase in the number of approvals, largely driven by an uptick in approvals in 2020 and 2021.
@@
 Measures of inputs are, typically, some estimate of R&D expenditures.
-The simplest construction of a measure of R&D expenditure scales the (inflation-adjusted) cost of a clinical trial by the number of clinical trials.
-Recent evidence on these objects comes from Bloom et al., (2020) and Sertkaya et al., (2024).
-Bloom et al., (2020) find a sharp increase in the number of clinical trials, as measured in publication data over time.
+The most widely cited analysis of pharmaceutical research productivity, Scannell et al., (2012), uses a measure of annual, aggregated industry-wide R&D expenditures, published by the trade organization Pharmaceutical Research and Manufacturers of America (PhRMA).
+[^14] These PhRMA estimates are used more widely.
+See e.g., Myers and Pauly, (2019) for use in an economics paper on pharmaceutical productivity.
+See e.g., Harris, (2011) for a discussion in the popular press.
+There are few other sources of R&D expenditure data for the pharmaceutical industry.
+DiMasi et al., (2003) and DiMasi et al., (2016) provide, perhaps, the most commonly cited estimates of drug development costs, derived from internal records of a select set of firms.
+In an editorial, Frank, (2003) describes these estimates as ”a matter of heated debate since they were first made public.” [^15] Light and Warburton, (2005), for example, criticizes these estimates derived from industry data as ”nonrandom and small” and characterizes the findings as based on ”unverifiable industry data.”
+Given these substantial limitations of industry-wide aggregate estimates, a natural alternative is to decompose aggregate R&D into its component parts.
+The simplest such decomposition would develop estimates of the cost of pre-clinical and clinical studies and scale this quantity by the number of such clinical trials.
+In practice, nearly all work focuses on clinical trials and assumes that pre-clinical trials contribute a fixed proportion of total development costs (DiMasi et al.,, 2003, 2016; Sertkaya et al.,, 2024).
+[^16] Ideally, estimates of clinical trial costs would be inflation-adjusted.
+To date, there are no price indices that are appropriate for such an adjustment.
+See Berndt and Cockburn, (2013) for an effort to construct such a price index.
+See Sertkaya et al., (2016) for a discussion of efforts to construct cost estimates, including limitations associated with this type of adjustment.
+Two recent estimates of these objects come from Bloom et al., (2020) and Sertkaya et al., (2024).
+In an influential analysis, Bloom et al., (2020) document a sharp increase in the number of clinical trials over time, using data constructed using scientific publications.
+In Bloom et al., (2020) and in a recent survey article, Goldin et al., (2024), this fact is central to the argument that pharmaceutical productivity is declining.
 Using granular data on clinical trial site contracts and internal records of the U.S.
-Food and Drug Administration, Sertkaya et al., (2024) find that the cost of clinical trials and the cost of drug development was roughly constant between 2000 and 2018.
+Food and Drug Administration, Sertkaya et al., (2024) find both that the cost of clinical trials and the cost of drug development were roughly constant between 2000 and 2018.171717The data used in the Sertkaya et al., (2024) analysis have previously been used by a small set of other researchers—notably, Azoulay, (2004) and Berndt and Cockburn, (2013).
+As of 2019, the firm that collects data on clinical trial contracts has been acquired.
+As of 2022, the data are no longer available for research use.
+Correspondence on file with the authors.
 Arguments of declining productivity in the pharmaceutical industry, then, depend substantially on the finding that the quantity of clinical trials is rising.
@@
 In the United States, the Food and Drug Administration Amendments Act (FDAAA), as passed in 2007 and revised in a ”Final Rule” that took effect in 2017, imposes these requirements on sponsors of regulated trials.
-Nominally, effected sponsors are required to register their studies on “ClinicalTrials.gov.” For nearly all studies, sponsors must also report the results of a trial within a designated period after completion.
+Technically, affected sponsors are required to register their studies on “ClinicalTrials.gov.” For nearly all studies, sponsors must also report the results of a trial within a designated period after completion.
 Moreover, increasingly, medical journals require registration as a pre-condition of publication (Laine et al.,, 2007).
-In practice, there is widespread noncompliance with these regulations and requirements.
+In practice, there is widespread noncompliance with these regulations and requirements—by private firms, academic institutions, and the federal government itself.
+[^18] There have been many extremely careful, detailed analyses of ClinicalTrials.gov non-compliance.
+See Anderson et al., (2015) for one comprehensive analysis.
+The FDAAA Trials Tracker—https://fdaaa.trialstracker.net/—provides up-to-date, record-by-record details on non-compliance.
 Deborah Zarin, then-director of ClinicalTrials.gov, notes that “funders, sponsors, and institutional review boards continue to allow unregistered trials or trials with late registration to be conducted, and some journals continue to allow the results of such trials to potentially be published” (Zarin et al.,, 2017).
 Noncompliance with federal reporting requirements is similarly widespread.
-In a series of papers assessing compliance over time DeVito et al., (2020) and DeVito and Goldacre, (2021) find that roughly 60 percent of trials registered on ClinicalTrials.gov fail to report required data elements at the time of their reporting deadline.
+In a series of papers assessing compliance over time, DeVito et al., (2020) and DeVito and Goldacre, (2021) find that roughly 60 percent of trials registered on ClinicalTrials.gov fail to report required data elements at the time of their reporting deadline.
 [^19] Note that noncompliance in this industry persists despite the existence of high statutory penalties.
@@
 Table B.1 lists five changes between 2000 and 2020 that altered the set of studies that should appear in a complete version of the database.
+Oostrom, (2024) documents, in a case study of psychiatric drugs, large changes in reporting propensity over this time period.
+[^20] More details on the Oostrom, (2024) comparison of registry data and publication data appear in Section 4.
 Second, sponsors engage in considerable “back-filling” of past studies.
@@
 In Appendix B, we illustrate the difficulty of recovering accurate measures of trends in the quantity of clinical trials from this database.
-We show that small changes in data construction yield meaningfully different estimates of trial levels and that many records are not in fact clinical trials.
-Moreover, due to widespread data gaps resulting from noncompliance in reporting of results and study attributes, there is little reliable information about characteristics of registered studies.
+In particular, we show that small changes in data construction yield meaningfully different estimates of trial levels and that many records are not in fact clinical trials.
+Moreover, due to widespread data gaps resulting from noncompliance in reporting of results and study attributes, there is little complete information on characteristics of registered studies.
+[^22] See Alsan et al., (2024) for a discussion of gaps in the reporting of the racial composition of studies.
 Other databases, such as the PDQ Cancer Information database maintained by the U.S.
-National Cancer Institute and used in papers including Budish et al., (2015), are disease-specific and, thus, insufficiently general for our purposes.
-A small number of firms (notably, GlaxoSmithKline) have disclosed all clinical trials over a longer period of time.
-[^15] See GlaxoSmithKline’s registry here: https://www.gsk-studyregister.com/ Of course, records from a single firm are not informative about aggregate trends, especially because many investigational drugs change hands frequently during the course of development (Cunningham et al.,, 2021).
+National Cancer Institute and used in papers including Budish et al., (2015), are disease-specific and thus insufficiently general for our purposes.
+A small number of firms (for example, GlaxoSmithKline) have disclosed all clinical trials over a longer period of time.
+[^23] See GlaxoSmithKline’s registry here: https://www.gsk-studyregister.com/.
+Gibson, (2004) provides an overview of the lawsuit that led to the development of this registry.
+Of course, records from a single firm are not informative about aggregate trends, especially because many investigational drugs change hands frequently during the course of development (Cunningham et al.,, 2021).
 Researchers, instead, often draw on proprietary databases.
-Although these databases are sufficient for certain research questions—for example, related to changes in the ownership of investigational drugs (Cunningham et al.,, 2021, using Pharmaprojects data) or on the link between insurance policy and drug development (Agha et al.,, 2022, using Cortellis data)—they are unsuitable for our particular productivity measurement exercise.
+Although these databases are extremely useful for certain research questions—for example, regarding changes in the ownership of investigational drugs (Cunningham et al.,, 2021, using Pharmaprojects data) or on the link between insurance policy and drug development (Agha et al.,, 2022, using Cortellis data)—they are unsuitable for analyses of either research trends (as in this paper) or summaries of the evidence base relevant for clinical decision-making (as in the health services literature).
 Our aim is to construct a census that allows for comparison of quantities across points in time.
-First, we are unable to determine, with specificity, how records are identified and catalogued based on available documentation.
+First, we are unable to determine, with specificity, how records are identified and cataloged based on available documentation.
 Second, these databases take public registries as inputs and, thus, risk adopting the same shifts in reporting over time (see e.g., the list of registry inputs for one frequently-used database, Cortellis,, 2024).
@@
 [^24] In Section 4, we discuss the two natural concerns with measures constructed from publication data—publication bias and shifts in reporting norms over time.
-Neither is likely to affect the interpretation of trends in our data.
+Our data, a recent analysis by Oostrom, (2024), and the relevant institutional context suggest, together, that neither is likely to threaten our interpretation of trends in publication data.
 The database PubMed / MEDLINE indexes roughly $34$ million publications, a near universe of published biomedical research.
-[^17] See Section A.1 for further details on on our treatment of the PubMed data.
+[^25] See Section A.1 for further details on our treatment of PubMed data.
 The challenge is to classify the specific forms of research captured by this corpus.
@@
 The existing literature takes several approaches.
-In health sciences—where data on the universe of clinical trials are relevant for the construction of meta-analyses and the development of prescribing guidelines—researchers typically employ over-inclusive methods with high true positive rates.
+In health sciences—where data on the universe of clinical trials are relevant for the construction of meta-analyses and the development of prescribing guidelines—researchers typically employ overly-inclusive methods characterized by high true positive rates.
 To use one prominent example, Cochrane, a British organization that synthesizes the findings of medical research, developed the “Cochrane RCT Classifier,” a machine learning-based classifier for retrieving randomized controlled trials.
@@
 For every eight true positives, however, it returns 92 false positives.
-See Thomas et al., (2021) for details.
+Thomas et al., (2021) provides details on the Cochrane approach.
 Approaches used by researchers in social science, instead, largely draw on metadata contained in PubMed.
@@
 The results are displayed in Figure 2.2.
-Further details on the construction of this figure are given in Section C.2.
 Each algorithm is estimated on the hand-labels assigned to the publications in the training and validation sample splits.
@@
 Third, it draws attention to differences in the performance of the two models.
-Observe that, in the “Other” error type, we include a sub-type called “overly literal interpretation of inclusion criteria.” We primary record errors of this type for Prompt 3, which asked the model to return a completion that described why, or why not, a record was classified as being in our sample.
+Observe that, in the “Other” error type, we include a sub-type called “overly literal interpretation of inclusion criteria.” We primarily record errors of this type for Prompt 3, which asked the model to return a completion that described why, or why not, a record was classified as being in our sample.
 Here, GPT-3.5 makes 30 such errors—classifying records as FALSE by recapitulating the sample definition provided.
@@
 BigBird is useful because abstracts of clinical trials regularly exceed 512 tokens.
-Prior to pre-training, the model was warm-started from the RoBERTa checkpoint (Liu et al.,, 2019) and then pre-trained on the standard BERT corpus.242424 The standard BERT corpus consists of the Books (Zhu et al.,, 2015), CC-News (Guu et al.,, 2020), Stories (Trinh and Le,, 2019), and Wikipedia (Wikimedia Foundation,, 2023) datasets.
+Prior to pre-training, the model was warm-started from the RoBERTa checkpoint (Liu et al.,, 2019) and then pre-trained on the standard BERT corpus.323232 The standard BERT corpus consists of the Books (Zhu et al.,, 2015), CC-News (Guu et al.,, 2020), Stories (Trinh and Le,, 2019), and Wikipedia (Wikimedia Foundation,, 2023) datasets.
 We evaluate two model sizes: BigBird Base (125 million parameters) and BigBird Large (355 million parameters).
@@
 Review of the associated full text confirms that this study did, in fact, enroll only rats.
-We report results associated with the moderate and liberal thresholds in the appendix as tests of robustness.
+All results that follow are robust to the use of the moderate and liberal thresholds.
 Our final, conservative, sample consists of 152,027 publications classified as satisfying the sample restrictions.
@@
 3. Disentangling Trends in Publication Data
-Figure 2.1 indicates that, across measures, the total quantity of potentially-relevant records in PubMed has increased substantially over time, including since 2010.282828Here, we focus on patterns in the production of publications that are similar to those that disclose the results of a clinical trial.
+Figure 2.1 indicates that, across measures, the total quantity of potentially-relevant records in PubMed has increased substantially over time, including since 2010.363636Here, we focus on patterns in the production of publications that are similar to those that disclose the results of a clinical trial.
 We find similar patterns in PubMed as a whole, and in alternative cuts of records.
@@
 To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
-These trends are robust to alternative forms of data construction.
-In Figure D.1, we construct these plots using our liberal and moderate sample definitions and find, qualitatively, no change.
-In Figure D.2, we explore a potential concern with a citation-based measure.
-We document in Figure 2.1 that the quantity of scientific publications is rising over time.
-It is reasonable to expect that this rise, coupled with the ease of citing papers enabled by the internet, coincides with an increase in the total number of citations in the aggregate and per paper.
-We confirm, empirically, that both hypotheses are borne out by our data.
-To account for the fact that the “cost” of citing a clinical trial may have decreased over time, thus generating the trend documented in Figure 3.1, we construct a version of this plot in which each citing paper is weighted by the total number of papers that it cites.
-That is, a citing paper that cites a trial and one other paper receives a weight of 1/2, and a citing paper that cites a trial and nine other papers receives a weight of 1/10.
-In doing so, we “penalize” papers that are especially generous in their tendency to cite other papers.
-Figure D.2 suggests, again, that this trend is robust to this alternate construction.
 3.2. Clinical Trials as a Productivity Indicator
@@
 We view the first potential explanation as unlikely for two reasons.
-Somewhat suggestively, NIH RePORTER data allow us to tab total expenditures associated with grants acknowledged in these publications.
+Somewhat suggestively, NIH RePORTER data allow us to tabulate total expenditures associated with grants acknowledged in these publications.
 These grant-based expenditure measures, too, are stable over this period.
@@
 Figure 3.3 provides a more granular view of this heterogeneity.
-We plot quantiles of the distribution of the number of (three-year) citations to publications in our census from publications in leading journals, for each year between 2010 and 2019.323232Appendix D presents these results using alternative cuts of our data and five-year citations.
+We plot quantiles of the distribution of the number of (three-year) citations to publications in our census from publications in leading journals, for each year between 2010 and 2019.404040Appendix D presents these results using alternative cuts of our data and five-year citations.
 The patterns remain unchanged.
@@
 The sample of published clinical trials is constructed with the conservative model.
-Figure D.4 is constructed analogously, where instead we display quantiles of the distribution of the number of citations received from all publications indexed in PubMed.
-There, we find that the right-tail of this distribution is increasing over time.
-In other words, as the total number of publications is increasing, the total number of citations received by clinical trials is increasing.
-However, these citations are not originating in leading journals.
-We interpret this set of facts as evidence that trends in clinical trial production have been stable over this time period. There are four potential concerns with this interpretation.
-First, if the costs of either upstream, basic inputs into clinical trials or the costs of clinical trials themselves have increased substantially over this time, the total resources devoted toward research will have increased, even though Figure 3.1 suggests that counts of trials produced are stable.
+We interpret this set of facts as evidence that trends in clinical trial production have been stable over this time period.
+At the very least, we find that there is insufficient evidence to reject the null that pharmaceutical research productivity has been stable over this time period.
+There are four potential concerns with this interpretation.
+First, if the costs of clinical trials have increased substantially during this time, the total resources devoted to the commercialization of drugs will have increased, although Figure 3.1 suggests that the quantity of trials produced has been stable.
 In our view, this is unlikely to be the case.
-On the issue of clinical trial costs, the most up-to-date, comprehensive evidence on drug development costs, in our read, is given in Sertkaya et al., (2024), who use internal data from the U.S.
+The most up-to-date, comprehensive evidence on clinical trial and drug development costs is given in Sertkaya et al., (2024), which uses internal data from the U.S.
 Food and Drug Administration and contracts associated with clinical trial sites to compute estimates for the period from 2000 to 2018.
@@
 Thus, we do not consider the prospect of rising study prices a concern for the interpretation of this pattern.
-On the issue of basic science inputs into clinical trials, there is little empirical evidence that increases in cost or decreases in the productivity have occurred.
+A distinct and somewhat related concern is that the cost of basic scientific research “upstream” of clinical trials could have increased substantially, either because more studies are necessary to produce a candidate drug or because the cost of inputs has increased.
+To be clear, our interest in this paper is on the productivity of the pharmaceutical industry, not biomedical research as a whole.
+Debates about the productivity of the pharmaceutical industry hinge on the idea that efforts to commercialize medicines are slowing—because firms are too inefficiently managed to run new studies (Ruffolo,, 2006), because concerns about commercial viability encourage firms to shelve valuable drug candidates (Scannell et al.,, 2012), or because regulation has become too onerous (Scannell et al.,, 2012).
+Changes in the productivity of basic science are interesting and important, to be sure, but not the central object.
+Nonetheless, in our read, there is essentially no empirical evidence indicating either decreases in the productivity of basic scientific studies or increases in their cost.
 A small number of papers examine frictions in markets for basic research.
 See, for example, Hill and Stein, (2021) and Myers, (2020).
-To our knowledge, this literature has not provided evidence of substantial shifts in inputs or factor prices over time, which in turn would increase the cost or decrease the productivity of basic science.
+To our knowledge, this literature has not provided evidence of substantial shifts in the quantity of inputs or factor prices over time, which in turn would increase the cost or decrease the productivity of basic science.
 In fact, changes in the production of basic scientific inputs around the year 2000—enabled by breakthroughs such as the Human Genome Project and the advent of high-throughput screening for drug discovery (see Scannell et al.,, 2012)—suggest, if anything, that costs of basic science inputs have fallen.
-Moreover, it is reasonable that such trends would affect trends in trial quantity, quality, composition, or cost.
-Given the absence of such patterns, it is unlikely that changes in basic science threaten the interpretation of our findings.
-[^34] Constructing a comprehensive measure of medical research productivity, which grapples with these various input markets, remains an open question.
-To our knowledge, the work that is closest in spirit is Berndt and Cockburn, (2013), who aim to develop a price index for clinical trial research, with the eventual goal of developing price indices for biomedical research more generally.
+In fact, the pharmaceutical productivity literature often characterizes the productivity “puzzle” in terms of this decrease in costs (see e.g., Cockburn,, 2006).
+There is some evidence that the total quantity of scientific papers has increased over this time period, which has been interpreted in turn as evidence of increasing (basic) scientific research effort—notably in Bloom et al., (2020) and Park et al., (2023).
+In our view, moving from these trends in publication data to a conclusion of decreasing productivity is premature.
+In the next section, we document that measured increases in the quantity of scientific publications capture, in substantial part, a sharp increase in the quantity of low-quality papers over time.
+[^42] Park et al., (2023) document the same: in their data, the quantity of low-impact research has increased substantially, while the quantity of high-impact research has remained essentially stable over time.
+Their measure of impact is a measure of “disruptiveness”—how much a given piece of research deviates from existing work.
+If measured increases in the quantity of papers reflect, say, an increase in the likelihood that the marginal paper is published, perhaps because the cost of publication has decreased, publication trends alone are largely uninformative on the question of scientific productivity.
+Our data also provide a more suggestive piece of evidence on trends in upstream research.
+We find remarkable stability in trial quantity, quality, and composition.
+Sertkaya et al., (2024) finds analogous stability in cost.
+If there were large changes in upstream productivity or cost, we might expect to see some indication in these downstream markets.
+In the absence of such indicators, we conclude that large shifts in the productivity of upstream markets are unlikely to be relevant considerations for the interpretation of our findings.
+[^43] Constructing a comprehensive measure of biomedical research productivity, which grapples with these various input markets, remains an open, important challenge.
+Our work does not speak directly to this challenge, but does highlight limitations in existing measures of productivity and challenges with the use of publication data.
 Second, one might be concerned about how to interpret the x-axis in our plots.
 In particular, the relationship between clinical trial completion dates and clinical trial start dates varies systematically within and across disease classes (Budish et al.,, 2015).
-Here, we follow existing literatures that use dates affixed to patents and publications as measures of the timing of research.
-Practically speaking, as Footnote 14 suggests, records of the start dates of clinical trials are back-filled in administrative databases and, thus, are available only with a considerable lag.
-We focus on the date of publication, then, for ease of interpretation.
-Third, our findings are threatened by publication bias.
-The existence of publication bias means that counts of scientific publications only partially measure the quantity of research (Andrews and Kasy,, 2019).
-Our main finding is that the quantity of clinical trials is stable over time.
+Here, we follow existing papers that use completion dates affixed to patents and publications as measures of the timing of research.
+In the context of clinical trials, the use of completion dates is reasonable as a practical matter.
+As Footnote 21 suggests, records of the start dates of clinical trials are back-filled in administrative databases and, thus, available only with a considerable lag.
+The date of publication is more easily interpretable.
+Third, our findings may be threatened by publication bias.
+Specifically, the existence of publication bias means that counts of scientific publications only partially measure the quantity of clinical trials being produced in each time period (Andrews and Kasy,, 2019).
+To weigh concerns about publication bias, it is useful to recall that our objective is to produce an estimate of trends in clinical trial production, sufficient to test existing hypotheses about declining productivity in this industry.
+Our aim is not to produce a comprehensive census of all clinical trials.
+Given this objective, publication bias is a serious threat to the interpretation of our results if it is meaningfully changing over this time period.
+Recent empirical evidence in Oostrom, (2024) suggests that publication bias may have been stable over this time period.
+Oostrom collects a set of roughly 600 clinical trials associated with psychiatric medications, as listed in a comprehensive meta-analysis spanning five decades.
+For each record, the author collects any associated scientific publications and any ClinicalTrials.gov entries.
+Of interest to our analysis are two facts in this case study.
+First, Oostrom documents evidence of publication bias across years of her data, but finds that trends in this bias become stable by the early 2000s.
+[^44] See Oostrom, (2024), Figure5(b).
+That is, although there is publication bias, it is not meaningfully changing in her sample of trials across our period of interest.
+Second, she finds that trends in ClinicalTrials.gov reporting shift sharply over our period of interest.
+In 2010, roughly 40 percent of the trials in her sample were registered on ClinicalTrials.gov.
+By 2019, roughly 80 percent were registered.
+Taken together, her findings lend credence to our assumption that publication bias and publication reporting norms are relatively stable during this period, whereas database reporting is not.
+In our data, we find that the quantity, quality, and composition of published clinical trials have been stable over time.
 If a trend were in fact present, and our measure were confounded by publication bias, then it would need to be the case that the trend was perfectly offset by the bias.
-It is reasonable to view cancellation, in this way, as unlikely.
+Cancellation in this way—across all three margins we consider—is unlikely.
 Fourth, and finally, we must assume that the number of publications associated with each clinical trial is constant across our period of interest.
@@
 Under these rules, there is little scope for authors to, for example, split results across several publications.
-Second, when constructing our sample of clinical trials, we have excluded publications that re-analyze data generated by an existing study.
+Second, when constructing our sample of clinical trials, we excluded publications that re-analyze data generated by an existing study.
 Third, more suggestively, we find no evidence of a change in this relationship between the size of NIH grants and the quantity of resulting publications reporting the results of a clinical trial during our period of interest.
@@
 3.3. Why is the Quantity of Research Rising?
-The quantity of papers citing clinical trials has increased, per Figure 3.1, by nearly a factor of two over our period of interest.
+The quantity of papers citing clinical trials has increased, per Figure 3.1, by nearly a factor of two across our period of interest.
 In this section, we examine various measures of publication composition and quality to shed light on the source of this increase.
@@
 These patterns provide one indication that the central tendency of the quality of non-clinical-trial medical research may be decreasing over time.
-[^44] Figure D.5 gives an analogous figure, displaying the distribution in citations from all publications to the sample of publications that cite clinical trials.
-Here, the right tail of the distribution of citations is growing.
 Figure 3.5. Citation Distribution Across Time, Papers that Cite Clinical Trials
@@
 Sites may struggle to find patients who meet these criteria, especially in small disease markets (Kolata,, 2017), and the costs of accruing large samples for certain studies can be prohibitive (Alsan et al.,, 2024).
-Our findings focus on one aspect of the larger biomedical research ecosystem: investments in efforts to bring basic science discoveries to market, in the form of new medicines.
-Development of supplemental measures of medical research productivity, which capture the efficiency of basic scientific research, is a useful area for future research.
+Our findings focus on one aspect of the larger biomedical research ecosystem: investments in commercialization efforts in the pharmaceutical industry, which bring products to market in the form of new medicines.
+Development of new measures of upstream research productivity—centered on basic discoveries in biomedical research and translational efforts to map these discoveries to products safe for testing in human subjects—are a natural area for future research.
 Our findings, also, apply to roughly one decade.
@@
 A long-standing discussion in medicine and statistics, spurred by Altman, (1994), centers on the argument that “we need less research, better research, and research done for the right reasons.” Our data indicate that nearly half of all clinical trials published in our sample are never cited by leading medical journals, and roughly 15 percent are never cited at all.
-Supplemental Appendix to: Distilling Data from Large Language Models:
-An Application to Research Productivity Measurement [^1] Date: September 11, 2024
+Supplemental Appendix to: Counting Clinical Trials:
+New Evidence on Pharmaceutical Sector Productivity [^1] Date: January 24, 2025
 $\star$ Stanford Department of Economics and Stanford Law School, maya.durvasula@stanford.edu
@@
 Maya M. Durvasula Sabri Eyuboglu David M. Ritzwoller
-Contents  appendix.Asubsection.A.1subsection.A.2subsubsection.A.2.1subsubsection.A.2.2subsubsection.A.2.3subsubsection.A.2.4subsubsection.A.2.5subsection.A.3subsection.A.4appendix.Bappendix.Csubsection.C.1subsection.C.2subsection.C.3subsection.C.4appendix.Dsubsection.D.1subsection.D.2subsection.D.3subsubsection.D.3.1subsubsection.D.3.2subsubsection.D.3.3appendix.E
+Contents  appendix.Asubsection.A.1subsection.A.2subsubsection.A.2.1subsubsection.A.2.2subsubsection.A.2.3subsubsection.A.2.4subsubsection.A.2.5subsection.A.3subsection.A.4appendix.Bappendix.Csubsection.C.1subsection.C.2subsection.C.3appendix.Dsubsection.D.1subsection.D.2subsubsection.D.2.1subsubsection.D.2.2subsubsection.D.2.3appendix.E
 Appendix A Publication Data
-In this Appendix, we detail our treatment of the PubMed / MEDLINE database (hereafter, “PubMed”).
-We give a general overview of the PubMed data in Section A.1.
-In Section A.2, we describe the process of identifying our initial sample of abstracts, i.e., the “Potential Universe” displayed in Figure 3.1.
-We discuss the process of determining the set of leading medical journals in Section A.3.
 A.1. The PubMed / MEDLINE Database
@@
 In practice, this means that so-called predatory journals are excluded, as are pre-prints and non-peer reviewed articles.
-Informal conversations with staff at the NLM suggest that MEDLINE should contain the universe of peer-reviewed publications in legitimate journals.
 PubMed includes a broader set of records, including pre-prints and publications deposited through alternative processes.
@@
 Throughout the text, and from this point forward in the Appendix, we refer to this database as “PubMed.”
-A natural question for researchers who rely on these data as a census of research investments is whether PubMed records are complete and accurate.
-To our knowledge, there is no paper that reports such validation exercises for each field of PubMed.
-We focus on those data elements most relevant to our work.
-There are 2,335,653 records in PubMed that have no associated year of publication, which we drop from our baseline sample.
 There are 16 records, across all years of data, that are not indexed with tags describing their contents (NML tags).
@@
 We find that the two measures have a correlation of 0.9.
-Researchers interested in different cuts of the data may need to conduct additional validation exercises.
-Figure 1 highlights that the frequencies of certain NLM tags have changed over time, often sharply.
-Researchers should account for such changes in any use of these flags.
-For other assessments of the completeness of certain PubMed fields, see Durvasula et al., (2021) and Ouellette and Sampat, (2024).
 A.2. The Universe of Potential Clinical Trials
@@
 Many countries have adopted similar guidance.
-Several countries and international organizations now maintain registries of trials.
 Registry identifiers are distinctive strings of letters and numbers.
@@
 Table A.2 reports the frequency of each registry identifier flagged across all records in PubMed and subset to those published after 1 January 2010.
-Figure 1 plots the number of trials with any registry identifier over time.
+Figure 2.1 plots the number of trials with any registry identifier over time.
 Reassuringly, we observe registry identifiers in PubMed data beginning around 2010, as registration mandates were implemented, and find a steady, smooth increase over time.
@@
 [^59] We use a copy of the Web of Science licensed to Stanford University.
-Web of Science data allow us to examine more detailed information about each author in our sample.
 While PubMed includes some details on authors, including limited institutional affiliation and address information, the Web of Science provides standardized author addresses for roughly 75 percent of observations in our sample.
@@
 When researchers have names that are more frequent in the population, our spot-checking (against online publication records) suggests that there are more likely to be errors.
-Nonetheless, tracking researchers using these identifiers is standard in this literature, as there are no obviously superior alternatives.
-Especially given that researcher productivity over time is not a centerpiece of our analysis, we do not undertake additional validation exercises.
 Appendix B ClinicalTrials.gov
@@
 Figure B.1. Number of Entries in Clinical Trials.gov, by Sample Restrictions
-Notes: Figure D.1 displays the counts of the number of records in ClinicalTrials.gov that satisfy various sample restrictions.
+Notes: Figure B.1 displays the counts of the number of records in ClinicalTrials.gov that satisfy various sample restrictions.
 The $x$-axis of the top panel is the year in which the clinical trial was completed.
@@
 Appendix C Prompt Design, Fine-Tuning, and Performance Assessment
-In this appendix, we detail the construction of the sample of clinical trials studied in Section 3.
-We begin in Section C.1 by describing the process we used to hand-label a sample of publications according to whether they satisfy the restrictions enumerated in 2.1.
-In Section C.2, we document poor performance of several approaches for classifying publications according to whether they belong in our sample based on standard machine learning methods.
-Section C.3 overviews the process of constructing prompts that perform well when used to query proprietary language models.
 C.1. Hand-Labeling
@@
 We devise an interface for abstract labeling, which allows us to review records quickly.
-Figure C.1 provides an example of this interface.
 A labeller reads in a set of publication records and is shown one PubMed identifier (PMID) and one abstract at a time.
@@
 The buttons for exclusion mirror the inclusion and exclusion criteria in 2.1, using short-hand convenient for labellers.
-We exclude records that report, only, the protocol for a clinical trial without results (“protocol”), that report the results of an observational (“observational”) or retrospective (“retrospective”) study, that report the results of a study involving animals (“animal”), that do not include human subjects (“no human subjects”), [^55] In practice, “animal” and “no human subjects” catch different sets of records.
-We flag records that include any animal with “animal.” Many studies in our sample of candidate records occur in vitro—in test tubes or other laboratory-based settings.
-There, no living subjects of any kind are enrolled and the “no human subjects” button is used.
-that do not study a drug (“no drug”), that report the results of a specific clinical case or patient experience (“case report”), or that summarize the findings of existing work (“meta-analysis”).
-By our own estimates, the labelling interface increased the speed of labelling by a factor of seven.
-Each record was reviewed, and labelled, twice.
 The hand-labeled data are split into three subsets—validation, training, and testing—based on their eventual use.
@@
 We do not add extra records to the testing dataset.
-Figure C.1. Abstract Labeling Interface
-Notes: Figure C.1 displays a screenshot of the application we developed to hand-label abstracts.
-Each abstract was labelled according to whether it satisfied the sample criteria enumerated in 2.1.
-If a publication did not satisfy our sample restrictions, we stipulated a reason.
-The user interface was developed with the “Meerkat” Python package available at the link: https://github.com/HazyResearch/meerkat.
-C.2. Benchmark Comparison to Standard Machine Learning Methods
-We measure the performance of a variety of standard machine learning algorithms for classifying publications according to whether they meet our sample restrictions.
-Models are trained in the training and validation samples and tested in the testing sample.
-We compute two embeddings for each abstract in our hand-labeled sample.
-First, we compute Term Frequency, Inverse Document Frequency (TF-IDF) embeddings based on the corpus of abstracts in the hand-labeled sample.
-Second, we extract embeddings associated with the abstract of each publication in the hand-labeled sample with the SentenceTransformers language model (Reimers and Gurevych,, 2019).
-The SentenceTransformers language model is a standard source for embeddings, and is a modification of the BERT language model (Devlin et al.,, 2019).
-We consider six classes of prediction algorithms: logistic ridge regression, logistic lasso regression, Support Vector Machines, Random Forest Regression, Boosted Regression, and Convolutional neural networks.
-Each model is obtained from the “sklearn” Python package.
-The hyper-parameters of each model are chosen with 10-fold cross-validation implemented in the training and validation samples.
-A balanced loss is used to measure the performance of each hyper-parameter.
-Figure 2.2 displays receiver operating characteristics for the each class of models estimated with both types of embedding.
-Error rates are estimated in the testing data.
-The performance of the receiver operating characteristic for the “ensemble” model whose performance is displayed in Figure 2.4 is given in purple.
-The fine-tuned large language model developed in this paper substantially out-paces the performance of standard machine learning methods.
-C.3. Prompt Design and Error Analysis
+C.2. Prompt Design and Error Analysis
 We extract weak labels for 64,000 randomly selected abstracts with two proprietary large language models: OpenAI’s GPT-3.5 and GPT-4 (Nori et al.,, 2023; Bubeck et al.,, 2023).
@@
 We report counts of errors, by model and prompt, of each Type and Sub-type.
-C.4. Error Analysis
+C.3. Error Analysis
 The Conservative model incorrectly labels 27 papers.
@@
 Appendix D Additional Figures and Further Analyses
-This appendix collects additional figures and analyses.
-D.1. Robustness of Trends
-Figure D.1 displays series analogous to the series displayed in Figure 3.1, constructed with the Moderate and Liberal sample restrictions. The results are very similar.
-One concern with the comparison made in Figure 3.1 is that researchers may cite a growing number of publications over time.
-That is, more papers may cite clinical trials over time because citations are becoming, in some sense, “cheaper.” The relevant clinical literature, then, may not be any larger.
-To alleviate such concerns, Figure D.2 reports counts, where we scale each citation assigned to a paper by the total number of citations in the originating paper.
-That is, a paper that receives two citations, each from a paper that cites two papers, will have a weighted citation count of 1.
-Figure D.3 displays series analogous to Figure 3.2 with the Moderate and Liberal sample restrictions.
-Figure D.4 gives a displays analogous Figure 3.3 for the distribution of citations received by clinical trials in each calendar year originating from all publications.
-Figure D.5 gives a displays analogous to Figure 3.5 for the distribution of citations received by publications citing clinical trials (three year sample) originating from all publications.
-Figures D.6 and D.7 display heat maps analogous to Figure 3.3 and Figure D.4, with the Moderate and Liberal sample restrictions and both 3-Year and 5-Year citation counts.
-In each case, the results are very similar across specifications.
-Figure D.1. Growth in Medical Research, by Sample Stringency
-Notes: Figure D.1 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” Each series is reported in terms of the percent change relative to its initial value.
-To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
-Figure D.2. Growth in Cost-of-Citation Weighted Medical Research
-Notes: Figure D.2 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” Here, we report the number of scientific papers using a citation-weighted metric.
-In particular, we scale each citation assigned to a paper by the total number of citations in the originating paper.
-Each series is reported in terms of the percent change relative to its initial value.
-To address truncation, we report the number of publications that cite clinical trials published in the preceding $t$ years for each $t$ between 2 and 6.
-Figure D.3. Quality Measures, by Sample Stringency
-Notes: Figure D.3 illustrates the stability of the heterogeneity in the composition and quality of published clinical trials between 2010 and 2019 in all three cuts of our data—which we term “conservative,” “moderate,” and “liberal.” See Figure 2.4 for the true positive and false positive rates of these models.
-The figure displays three time series: the proportion of clinical trials that are publicly funded, the proportion of clinical trials that are ever cited by a leading journal, and the proportion of publicly funded clinical trials that are ever cited by a leading journal.
-Figure D.4. Citation Distribution Across Time, All Publications to Clinical Trials
-Notes: Figure D.4 displays a heat map measuring the distribution of citations received by clinical trials in each calendar year from all publications.
-The $y$-axis has been stretched to elongate the right-tail of the citation distribution.
-Colors are displayed in a log scale.
-The sample of published clinical trials is constructed with the conservative model.
-Figure D.5. Citation Distribution Across Time, All Publications to Citing Publications
-Notes: Figure D.4 displays a heat map measuring the distribution of citations received by papers that cite clinical trials, in each calendar year, from all publications.
-The $y$-axis has been stretched to elongate the right-tail of the citation distribution.
-Colors are displayed in a log scale.
-The sample of published clinical trials is constructed with the conservative model.
-The sample of papers that cite clinical trials is constructed using a three-year window.
-Figure D.6. Distribution of 3-Year Cites, Clinical Trials, by Sample Stringency
-Panel A: Citations From Publications in Leading Journals
-Panel B: Citations From All Publications
-Notes: Figure D.6 displays heat maps measuring the distribution of three-year citations received by clinical trials in each calendar year from leading journals and from all publications, respectively.
-The $y$-axis has been stretched to elongate the right-tail of the citation distribution.
-Colors are displayed in a log scale.
-We consider the sample of clinical trials constructed with the conservative, moderate, and liberal model.
-Figure D.7. Distribution of 5-Year Cites, Clinical Trials, by Sample Stringency
-Panel A: Citations From Publications in Leading Journals
-Panel B: Citations From All Publications
-Notes: Figure D.7 displays heat maps measuring the distribution of five-year citations received by clinical trials in each calendar year from leading journals and from all publications, respectively.
-The $y$-axis has been stretched to elongate the right-tail of the citation distribution.
-Colors are displayed in a log scale.
-We consider the sample of clinical trials constructed with the conservative, moderate, and liberal model.
-D.2. Growth in Citing Papers by Geography
+D.1. Growth in Citing Papers by Geography
 In Section 3.3.2, we document differential trends in the production of publications by type.
@@
 The fifth column displays the percent change in the number of ‘citing’ papers between 2013 and 2019.
-D.3. Quantifying Meta-Analyses
+D.2. Quantifying Meta-Analyses
 In Section 3.3.1, we present three methods that allow us to identify “review” publications—publications that summarize the findings of existing studies—in our set of citing papers.
@@
 Here, we provide additional details on these three approaches.
-D.3.1. Hand-Labeling Meta-Analyses
+D.2.1. Hand-Labeling Meta-Analyses
 From the set of publications that cite clinical trials (three-year sample), we drew 100 observations randomly from each publication year.
@@
 We report the number of publications, in each year, that fall into each category.
-D.3.2. Text Searches for Meta-Analyses
+D.2.2. Text Searches for Meta-Analyses
 Recall that, in the process of constructing the training data that were used to fine-tune our large language models, we identified many records in PubMed that were meta-analyses, literature reviews, etc., based on the contents of their text.
@@
 We constructed a list of phrases and keywords that consistently indicated, to a human labeler, that a record is a meta-analysis or review article.
-This list is reproduced in Footnote 38.
+This list is reproduced in Footnote 48.
 Using this list of search terms, we search the abstracts of all records in our citing sample.
-We flag any record with any of the keywords listed in Footnote 38 as a meta-analysis and flag any record that mentions at least two of the databases listed in Footnote 38 as well.
+We flag any record with any of the keywords listed in Footnote 48 as a meta-analysis and flag any record that mentions at least two of the databases listed in Footnote 48 as well.
 The objective of this exercise is to yield a conservative estimate of the change in the quantity of meta-analyses over our period of interest.
@@
 Panel B displays proportions of publications that satisfy both sets of criteria.
-D.3.3. NLM Tag Queries for Meta-Analyses
+D.2.3. NLM Tag Queries for Meta-Analyses
 Next, we observe that NLM tags, added when publication records are indexed in PubMed, also record information that may signal whether a record is a meta-analysis.